### v2

# Asymptotic Velocity Domination

Abstract
Asymptotic velocity domination (AVD) posits that when back-propagated to the Big Bang generic cosmological spacetimes solve a drastically simplified version of the Einstein field equations, where all dynamical spatial gradients are absent (similar as in the Belinski-Khalatnikov-Lifshitz scenario). Conversely, a solution can in principle be reconstructed from its behavior near the Big Bang. This property has been rigorously proven for the Gowdy class of cosmologies, both polarized and unpolarized. Here we establish for the polarized case a quantum version of the AVD property formulated in terms of two-point functions of (the integrands of) Dirac observables: these correlators approach their much simpler velocity dominated counterparts when the time support is back-propagated to the Big Bang. Conversely, the full correlators can be expressed as a uniformly convergent series in averaged spatial gradients of the velocity dominated ones.
1.  Introduction
The Big Bang singularity is one of the prime happenstances where quantum gravitational effects are expected to be important, both in imprinting observable signatures and as a quest for ultimate origins. So far theoretical efforts to develop a quantum framework specific to the situation have mostly been limited to mini-super space models with a few degrees of freedom. Gowdy cosmologies offer an interesting alternative: (i) they have field theoretical degrees of freedom; (ii) mirror the structure of full Einstein gravity; (iii) are amenable to standard quantization techniques; and (iv) they exhibit classically a remarkable behavior near the Big Bang known as Asymptotic Velocity Domination (AVD). AVD posits that, when generic cosmological solutions are evolved backward towards the Big Bang, they enter a velocity dominated regime in which temporal derivatives overwhelm spatial ones (similar as in the Belinski-Khalatnikov-Lifshitz scenario [1]) and in which they are governed by a much simpler system of equations (called “Velocity Dominated” VD) without dynamical spatial gradients. Conversely, the full solution can in principle be reconstructed from the limiting one and eventually from data specified on the Big Bang “boundary”; see [2] for an overview. For Gowdy cosmologies, AVD has been rigorously proven [3, 4], as reviewed in [5].
Gowdy cosmologies [6] are exact, spatially inhomogeneous, vacuum solutions of the Einstein equations with two commuting spacelike Killing vectors. They describe nonlinear colliding gravitational waves emerging from a Big Bang singularity. Prominent among them are soliton-like solutions reviewed in [7]. The system can be quantized and for generic polarizations leads to an effectively 1+11\!+\!1 dimensional quantum field theory with exponential self-interactions and a noncompact internal symmetry. It is found to be quasi-renormalizable to all orders of perturbation theory in a way compatible with the broader Asymptotic Safety Scenario [8, 9, 10]. These results do not address aspects related to AVD. Precisely those aspects, however, have the potential to generalize beyond the two-Killing vector subsector and thus should contain pointers towards a ‘quantum theory of the Big Bang’. With this motivation, we investigate here whether a quantum version of AVD holds in polarized Gowdy cosmologies. The polarized case shares most qualitative features with the general case, but is technically much simpler in that the action is quadratic in the main dynamical field $\phi$. In a reduced phase space formulation the quantum theory can therefore be explored using methods familiar from free quantum field theories on curved backgrounds; see [11, 12, 13, 14]. We shall do so here as well, but focus on aspects pertaining to AVD.
An important property of polarized Gowdy cosmologies is that they admit an infinite set of Dirac observables [15]. Dirac observables are thought to capture the intrinsic ‘gauge invariant’ content of a gravity theory, but are normally elusive, see e.g. [16]. In Appendix B we construct a new one-parameter family of the form
${\cal O}(\theta)=\int_{0}^{2\pi}\!\frac{dx^{1}}{2\pi}\,\jmath_{0}^{{\textsc{p}}}(x^{0},x^{1};\theta)={\cal O}_{0}(\theta)+O(\rho^{2}/\lambda_{{\textsc{n}}}^{2})\,,\quad{\rm Im}\theta\neq 0\,.$ (1.1)
Here $\rho>0$ is a temporal function whose $\rho=0$ level set can be identified with the Big Bang. Further, $(x^{0},x^{1})$ are the non-Killing coordinates of the Gowdy spacetime and $\lambda_{{\textsc{n}}}>0$ is the dimensionless reduced Newton constant. These quantities strongly Poisson commute with the constraints without using equations of motion. The leading term ${\cal O}_{0}(\theta)$ is in one-to-one correspondence to a family of Dirac observables in the VD system, while the subleading terms are organized alternatively according to powers of $\rho^{2}/\lambda_{{\textsc{n}}}^{2}$ or by the number of dynamical spatial derivatives. The integrand $\jmath_{0}^{{\textsc{p}}}$ is linear in the Gowdy scalar $\phi$ (defining $M={\rm diag}(e^{\phi},e^{-\phi})$ in (2.1) below) and its momentum $\pi^{\phi}$, and in a gauge fixed reduced phase space formulation only $(\phi,\pi^{\phi})$ is canonically quantized, while the field $\tilde{\sigma}$ in (2.1) is treated as a renormalized composite operator. Suitable gauge specializations single out coordinates $(\tau,\zeta)\in\mathbb{R}^{2}$, such that the Big Bang is located at $\tau\rightarrow$-$\infty$. Then ${\cal O}(\theta)$ becomes an on-shell conserved charge, $\partial_{\tau}{\cal O}^{\rm on}(\theta)=0$, whose integrand $q(\tau,\zeta;\theta)$ differs from $\jmath_{0}^{{\textsc{p}}}(e^{\tau},\zeta;\theta)$ by a total $\partial_{\zeta}$ derivative, and stays finite as $\tau\rightarrow$-$\infty$. The two-point function of the on-shell $q$ can be taken as the central object in the quantum theory, $\langle 0_{T}|q(\tau,\zeta;\theta)q(\tau^{\prime},\zeta^{\prime};\theta)|0_{T}\rangle$, where $|0_{T}\rangle$ is some (non-unique) Fock vacuum defining a ‘state’. Only its symmetric part is state dependent and on account of the linearity in $\phi$ its properties are coded by those of $\langle 0_{T}|\phi(\tau,\zeta)\phi(\tau^{\prime},\zeta^{\prime})+\phi(\tau^{\prime},\zeta^{\prime})\phi(\tau,\zeta)|0_{T}\rangle$ and its $\partial_{\tau},\partial_{\tau^{\prime}},\partial_{\tau}\partial_{\tau^{\prime}}$ derivatives. These are conveniently combined into a $2\times 2$ matrix two-point function ${\cal W}^{s}$. As described in Appendix A, one can view the velocity dominated Gowdy system as a Carroll-type gravity theory in its own right and follow analogous steps through VD Dirac observables and their integrand’s two-point function to find its properties coded by a matrix two-point function $\mathfrak{W}^{s}$ in the VD system analogous to ${\cal W}^{s}$. The question of quantum AVD can then be posed as follows: does ${\cal W}^{s}$ have a $\tau,\tau^{\prime}\rightarrow$-$\infty$ asymptotics that relates to its VD counterpart $\mathfrak{W}^{s}$? If so, can ${\cal W}^{s}$ be reconstructed from $\mathfrak{W}^{s}$? We find that the answers are affirmative for a large class of states we dub “time consistent”. Specifically, we obtain the
Result. For any time consistent state the matrix two-point function ${\cal W}^{s}$ admits upon averaging with $f,g\in C^{\infty}(S^{1})$ test functions a series expansion of the form
$\int\!\!\frac{d\zeta}{2\pi}\frac{d\zeta^{\prime}}{2\pi}f(\zeta)g(\zeta^{\prime}){\cal W}^{s}(\tau,\tau^{\prime},\zeta\!-\!\zeta^{\prime})=\int\!\!\frac{d\zeta}{2\pi}\frac{d\zeta^{\prime}}{2\pi}f(\zeta)g(\zeta^{\prime})\,\mathfrak{W}^{s}(\tau,\tau^{\prime},\zeta\!-\!\zeta^{\prime})$
$\qquad+\sum_{l\geq 1}\sum_{k=0}^{l}e^{2k\tau}e^{2(l-k)\tau^{\prime}}\int\!\!\frac{d\zeta}{2\pi}\frac{d\zeta^{\prime}}{2\pi}f(\zeta)g(\zeta^{\prime})(\partial_{\zeta}\partial_{\zeta^{\prime}})^{l}I_{k}\,\mathfrak{W}^{s}(\tau,\tau^{\prime},\zeta\!-\!\zeta^{\prime})\,I_{l-k}^{T}\,.$ (1.2)
Here $I_{k}$ are numerical $2\times 2$ matrices with a known generating functional. The expansion is fully determined by the spatial gradients of the two-point function in the VD system, $\mathfrak{W}^{s}(\tau,\tau^{\prime},\zeta-\zeta^{\prime})$, which is linear in $\tau,\tau^{\prime}$. The series (1) is uniformly convergent for all $\tau,\tau^{\prime}<-\delta$, $\delta>0$ and $\tau-\tau^{\prime}$ bounded.
The paper is organized as follows. After introducing the polarized Gowdy system in relation to its AVD property in Subsections 2.1 and 2.2, we outline the reduced phase space quantization and define the $qq$ two-point functions as well as ${\cal W}^{s}$. In Section 3 the properties of the various two-point functions are explored, leading to the notion of time dependent states and the above spatial gradient expansion. The construction of the $\sigma$ field solving the quantum constraints is done in Section 4. The Hamiltonian actions and their symmetries are detailed in Appendix A, which are prerequisites for the construction of Dirac observables (1.1) in Appendix B.
2.  Polarized Gowdy cosmologies and AVD
Gowdy cosmologies are exact solutions of the 4-dimensional Einstein equations with two commuting spacelike Killing vectors $K_{1},K_{2}$. In adapted coordinates $X=(x^{0},x^{1},y^{1},y^{2})$ the line element can be brought into the form
$g^{\rm 2K}_{IJ}(X)dX^{I}dX^{J}=\gamma_{\mu\nu}(x)dx^{\mu}dx^{\nu}+\rho(x)M_{ab}(x)dy^{a}dy^{b}\,.$ (2.1)
Here, all fields only depend on $(x^{0},x^{1})$ while $(y^{1},y^{2})$ are Killing coordinates, i.e. are such that locally $K_{a}=K_{a}^{I}\partial_{I}=\partial/\partial y^{a}$, $a=1,2$. Further, $\rho(x)>0$, and $M(x)$ is a real-valued $2\times 2$ matrix with unit determinant. A selfcontained derivation of (2.1) (for the case of one timelike and one spacelike Killing vector) can be found in [17], Appendix C. For the Lorentzian metric $\gamma$ we adopt the following lapse-shift type parameterization
$\gamma_{\mu\nu}(x)dx^{\mu}dx^{\nu}=e^{\tilde{\sigma}}[-n^{2}(dx^{0})^{2}+(dx^{1}+s\,dx^{0})^{2}]\,$, (2.2)
i.e. $\gamma_{00}=e^{\tilde{\sigma}}(-n^{2}+s^{2})$, $\gamma_{01}=e^{\tilde{\sigma}}s=\gamma_{10}$, $\gamma_{11}=e^{\tilde{\sigma}}$. By a slight abuse of terminology we shall refer to $n,s$ as lapse, shift, respectively. [^1] For a two-dimensional Lorentzian metric the lapse $N$, shift $N^{1}$, spatial metric $\gamma_{11}$ proper would be $N=e^{\tilde{\sigma}/2}n$, $N^{1}=s$, $\gamma_{11}=e^{\tilde{\sigma}}$. Note that $e^{\tilde{\sigma}},n$ are spatial densities of weight 2,12,-1, respectively, and that $s$, being one-dimensional spatial vector, can also be viewed as a spatial density of weight 1-1. Cosmological spacetimes arise in this setting if $\rho$ is assumed to have everywhere timelike gradient, $\gamma^{\mu\nu}(x)\partial_{\mu}\rho\partial_{\nu}\rho<0$. This evaluates to $e_{0}(\rho)^{2}>n^{2}(\partial_{1}\rho)^{2}$, where $e_{0}(\rho)=(\partial_{0}-s\partial_{1})\rho$. The field equations are presented in (LABEL:evolpt) below. The solutions generically have a Big Bang singularity, which in the coordinates singled out by $s=0,n=1,\rho=t$ occurs at the $t\rightarrow 0^{+}$ hypersurface. The evolution equations can be coded by a Lax pair and some of the integrable systems methodology can be adapted to this situation. In particular, there is a rich set of gravitational soliton solutions [7].
The Killing coordinates $y^{1},y^{2}$ are assumed to be periodic, corresponding to orbits with the topology of a 2-torus $T^{2}$. The $x^{1}$ spatial sections are normally taken to be circular as well, in which case all fields in (2.1) are periodic functions of $x^{1}\in\mathbb{R}$. The overall topology of the manifold then is $\mathbb{R}\times T^{3}$ and the solutions are referred to as $T^{3}$ Gowdy cosmologies. Technically, it is often simpler to allow the fields to be be nonperiodic, with suitable fall-off conditions. The topology then is $\mathbb{R}\times\mathbb{R}\times T^{2}$. Other possible topologies are $\mathbb{R}\times S^{3}$ or $\mathbb{R}\times S^{1}\times S^{2}$ [6].
Here we shall focus on the $\mathbb{R}\times T^{3}$ case, arguably the most natural one in that all spatial dimensions are treated on the same footing. Moreover, AVD has been rigorously proven for this class [5]. Specifically, we consider the polarized $T^{3}$ Gowdy cosmologies, where the matrix $M$ is diagonal, $M(x)={\rm diag}(e^{\phi(x)},e^{-\phi(x)})$. AVD for this case has been established much earlier in [3]; moreover, the key field equations are linear, and on the reduced phase space the system is amenable to a straightforward canonical quantization [11, 12, 13, 18]. For the construction of Dirac observables on the full phase space a Hamiltonian action principle and its gauge symmetries is the appropriate starting point. These aspects are prepared in Appendices A and B. Below we outline how the action leads to the reduced phase space convenient for quantization.
2.1 Action, proper time gauge, and reduced phase space
Inserting (2.1) with $M(x)={\rm diag}(e^{\phi(x)},e^{-\phi(x)})$ into the Gibbons-Hawking action results in a valid action principle for the polarized Gowdy cosmologies:
$S^{\textsc{l}}=\frac{1}{\lambda_{{\textsc{n}}}}\int_{t_{i}}^{t_{f}}\!\!dx^{0}\!\!\int_{0}^{2\pi}\!\!dx^{1}\Big\{\!-\frac{1}{n}e_{0}(\rho)e_{0}(\sigma)+n\big(\partial_{1}\rho\partial_{1}\sigma\!-\!2\partial_{1}^{2}\rho\big)+\frac{\rho}{2n}e_{0}(\phi)^{2}-\frac{\rho n}{2}(\partial_{1}\phi)^{2}\Big\}.$ (2.3)
This action can be inferred from [19, 13, 15] or directly verified using computer algebra. Here, $\lambda_{{\textsc{n}}}>0$ is the dimensionless reduced Newton constant, $e_{0}=\partial_{0}-{\cal L}_{s}$ (with ${\cal L}_{s}$ the one-dimensional Lie-derivative acting on spatial tensor densities) is the derivative transversal to the leaves of the $x^{0}={\rm const}.$ foliation, and $\sigma=\tilde{\sigma}+\frac{1}{2}\ln\rho$. We use $x^{1}\in\mathbb{R}$ but all fields in (2.3) are assumed to be spatially periodic with period $2\pi$ (and thus having a well-defined projection onto $S^{1}=\mathbb{R}/2\pi\mathbb{Z}$.) This form of the action is tailored towards holding the fields $\rho,\sigma,\phi$ fixed at the boundaries $x^{0}=t_{i},t_{f}$. Its Hamiltonian version (A.1) allows one to identify the canonically induced gauge symmetries (A.4). The subset (A.4,a,b) with the velocity-momentum relations inserted on the right hand side of (A.4b) yields the gauge variations $\delta_{\epsilon}^{\textsc{l}}$ of (2.3). Variation of $S^{\textsc{l}}$ with respect to $n,s$ gives the Lagrangian constraints
$\frac{\delta S^{\textsc{l}}}{\delta n}=-{\cal H}_{0}^{\textsc{l}}\,,\quad\frac{\delta S^{\textsc{l}}}{\delta s}=-{\cal H}_{1}^{\textsc{l}}\,.$ (2.4)
Here ${\cal H}_{0}^{\textsc{l}},{\cal H}_{1}^{\textsc{l}}$ coincide with ${\cal H}_{0},{\cal H}_{1}$ from (A.1) upon insertion of the velocity-momentum relations. Importantly, the vanishing conditions ${\cal H}_{0}^{\textsc{l}}=0={\cal H}_{1}^{\textsc{l}}$ can be solved algebraically for $e_{0}(\sigma)=F_{0}(\rho,\phi)$ and $n\partial_{1}\sigma=F_{1}(\rho,\phi)$, where
$F_{0}(\rho,\phi)$ $\!\!=\!\!$ $\frac{e_{0}(\rho)}{e_{0}(\rho)^{2}\!-\!(n\partial_{1}\rho)^{2}}\Big(2n^{2}\partial_{1}^{2}\rho+\frac{\rho}{2}[e_{0}(\phi)^{2}+(n\partial_{1}\phi)^{2}]\Big)$
- $\frac{n\partial_{1}\rho}{e_{0}(\rho)^{2}\!-\!(n\partial_{1}\rho)^{2}}\big(2n^{2}\partial_{1}(e_{0}(\rho)/n)+\rho e_{0}(\phi)n\partial_{1}\phi\big)\,$,
$F_{1}(\rho,\phi)$ $\!\!=\!\!$ $\frac{e_{0}(\rho)}{e_{0}(\rho)^{2}\!-\!(n\partial_{1}\rho)^{2}}\big(2n^{2}\partial_{1}(e_{0}(\rho)/n)+\rho e_{0}(\phi)n\partial_{1}\phi\big)$ (2.5)
- $\frac{n\partial_{1}\rho}{e_{0}(\rho)^{2}\!-\!(n\partial_{1}\rho)^{2}}\Big(2n^{2}\partial_{1}^{2}\rho+\frac{\rho}{2}[e_{0}(\phi)^{2}+(n\partial_{1}\phi)^{2}]\Big)\,.$
The fact that the constraints in 2-Killing vector reductions can be solved for one of the fields is well-known [7, 17]; we present (2.1) mostly to highlight that this can be done without gauge fixing or coordinate choices. By assumption, $e_{0}(\rho)^{2}\!-\!(n\partial_{1}\rho)^{2}>0$ for Gowdy cosmologies. Since $\partial_{1}e_{0}(\sigma+2\ln n)=e_{0}\partial_{1}(\sigma+2\ln n)$ the integrability condition $\partial_{1}[F_{0}+2e_{0}(\ln n)]=e_{0}[F_{1}/n+2\partial_{1}\ln n]$ must be satisfied. This is indeed the case subject to the evolution equations for $\rho,\phi$, and re-expresses the preservation (A.6) of the constraints under time evolution. The two first order relations can therefore locally be integrated to yield $\sigma$ as a functional of the on-shell $\rho$ and $\phi$, i.e. $\sigma=\sigma[\rho^{\rm on},\phi^{\rm on}]$. Further, $e_{0}(n^{-1}e_{0}(\sigma))-\partial_{1}(n\partial_{1}\sigma)=e_{0}(n^{-1}F_{0})-\partial_{1}F_{1}=2\partial_{1}^{2}n-e_{0}(\phi)^{2}/(2n)+n(\partial_{1}\phi)^{2}/2$ reproduces the equation of motion for $\sigma$, i.e. $\delta S^{\textsc{l}}/\delta\rho=0$. That is, the functional $\sigma=\sigma[\rho^{\rm on},\phi^{\rm on}]$ is automatically on-shell as well. Moreover, if $\rho^{\rm on},\phi^{\rm on}$ are parameterized by their boundary values at $x^{0}=t_{i},t_{f}$, so will be $\sigma[\rho^{\rm on},\phi^{\rm on}]$, rendering $\sigma|_{t_{i}},\sigma|_{t_{f}}$ a functional of $\rho|_{t_{i}},\rho|_{t_{f}},\phi|_{t_{i}},\phi|_{t_{f}}$. Together, this allows one to take the fields $\rho,\phi$ as basic and to treat $\sigma$ as a composite field in them.
Proper time gauge. So far, no gauge fixing entered. Both, for setting up a functional integral and for ensuring a unique classical time evolution one needs to fix a gauge. The action principle (2.3) with fixed boundary fields would enter the functional integral for the propagation kernel, in which case the proper time gauge is the appropriate choice [20]. The constraints are then imposed on the Schrödinger picture wave functionals and the propagation kernel is evaluated in the gauge
$s=0=\partial_{0}n\,.$ (2.6)
In this gauge the proper time interval $\tau(x^{1})=n(x^{1})\int_{t_{i}}^{t_{f}}dx^{0}(e^{\tilde{\sigma}/2})(x^{0},x^{1})$, between points on the initial and final spacelike hypersurfaces is fixed. Since $\tau$ is a spatial scalar, the elapsed proper time does not depend on the choice of spatial coordinates used (the same though for all level surfaces). The only gauge transformations $\delta^{\textsc{l}}_{\epsilon}$ (with $\epsilon|_{t_{i}}=0=\epsilon|_{t_{f}}$) preserving these conditions have parameters of the form $\epsilon=0,\epsilon^{1}=\epsilon^{1}(x^{1})$. Writing $\delta_{\epsilon^{1}}$ for the restricted gauge variations one has $\delta_{\epsilon^{1}}\sigma=\epsilon^{1}\partial_{1}\sigma+2\partial_{1}\epsilon^{1}$, $\delta_{\epsilon^{1}}n=\epsilon^{1}\partial_{1}n-n\partial_{1}\epsilon^{1}$, $\delta_{\epsilon^{1}}\rho=\epsilon^{1}\partial_{1}\rho$, $\delta_{\epsilon^{1}}\phi=\epsilon^{1}\partial_{1}\phi$. The residual purely spatial gauge invariance is useful for bookkeeping purposes.
As seen above, the solution of the constraints determines the $\sigma$ field in terms of on-shell $\rho,\phi$ satisfying (2.10). Consistent therewith, the periodic $\sigma$ field can in the functional integral for the propagation kernel in proper time gauge be integrated out to yield a $\delta$ distribution insertion with argument $\partial_{0}^{2}\rho-(n\partial_{1})^{2}\rho$. Here we allude to the fact (not elaborated here) that the Faddeev-Popov factors for the gauge (2.6) do not contain $\sigma$ in differentiated form and can naturally be interpreted as $\sigma$ independent. The $\delta$ function insertion produced implements, of course, the evolution equation for $\rho$, so the Schrödinger picture wave functionals should be projected onto the joint solution set of ${\cal H}_{0}^{\textsc{l}}=0,{\cal H}_{1}^{\textsc{l}}=0$, $\partial_{0}^{2}\rho-(n\partial_{1})^{2}\rho=0$. This leads to the reduced phase space quantization: The basic quantum fields are $\phi,\rho$ constrained by (2.10) below and $\partial_{0}^{2}\rho-(n\partial_{1})^{2}\rho=0$. The dynamics of the $\phi$ field is governed by the propagation kernel computed from the second part of the action $S^{\textsc{l}}$ evaluated in proper time gauge and with prescribed boundary values $\phi|_{t_{i}},\phi|_{t_{f}}$. The $\rho$ field at this point is nondynamical and just sets a background geometry of sorts. The propagation kernel could be evaluated explicitly in terms of the Hamilton-Jacobi principal function, but will not be needed. It acts on spatially reparameterization invariant Schrödinger picture wave functionals $\Psi[\phi|_{t_{i}}]$ by convolution. The non-unique vacuum wave functionals will be Gaussians, analogous to the situation for free quantum field theories on spatially homogeneous time dependent backgrounds, see e.g. [14, 21]. By construction, the $\phi$ two-point functions in the so-defined Schrödinger picture will coincide with the ones evaluated in the Heisenberg picture with an underlying (non-unique) Fock vacuum, see Section 3 for the latter. The only remnants of the gravitational origin of the system are: (i) the constraint (2.10) needs to be taken into account, and (ii) the $\sigma$ field needs to be constructed as a renormalized composite operator in $\rho,\phi$, see Section 4.
So far, we maintained spatial reparameterization invariance by carrying $n=n(x^{1})$ along. This residual gauge invariance could be gauge fixed as well. Technically, it is simpler to absorb $n$ into a pseudo-scalar
$\zeta(x^{1})=\int_{y^{1}}^{x^{1}}\!\frac{dx}{n(x)}\,.$ (2.7)
Note that with a spatially periodic $n$ the quantity (2.7) is quasi-periodic, $\zeta(x^{1}+2\pi)=\zeta(x^{1})+\zeta_{0}$, $\zeta_{0}=\int_{0}^{2\pi}dx/n(x)$. By suitable constant rescalings we may assume that $\zeta_{0}=2\pi$, so that the periods in $x^{1}$ and $\zeta$ are the same. Since $\partial_{1}\zeta=1/n>0$ the map $\zeta:\mathbb{R}\rightarrow\mathbb{R}$ is invertible and all quantities can be regarded as functions of $\zeta$ rather than $x^{1}$. For simplicity we retain the original function symbols and write $\phi(x^{0},\zeta)$, $n(\zeta)$, etc., viewed as $2\pi$ periodic functions of $\zeta$. The transcription of the action and field equations is straightforward as mostly $n\partial_{1}=\partial_{\zeta}$ enters, while the exceptional $n\partial_{1}^{2}n$ term transcribes into $\partial_{\zeta}^{2}\ln n$. In particular, the constraints are solved by the transcription of (2.1) and yield $\partial_{0}\sigma$ and $\partial_{\zeta}(\sigma+2\ln n)$ as a function of the $\partial_{0}$ and $\partial_{\zeta}$ derivatives of $\phi,\rho$, while $n$ no longer occurs. For later reference we note the independent field equations in this version of the proper time gauge
$\partial_{0}(\rho\partial_{0}\phi)=\partial_{\zeta}(\rho\partial_{\zeta}\phi)\,,\quad\partial_{0}^{2}\rho=\partial_{\zeta}^{2}\rho\,$$\quad$
$\partial_{0}\sigma=F_{0}^{\rm pt}(\rho,\phi)\,,\quad\partial_{\zeta}(\sigma+2\ln n)=F_{1}^{\rm pt}(\rho,\phi)\,$, (2.8)
where $F_{0}^{\rm pt}$, $F_{1}^{\rm pt}$ are obtained from $F_{0},F_{1}$ in (2.1) by the specialization to (2.6), (2.7). Explicitly,
$F_{0}^{\rm pt}(\rho,\sigma)$ $\!\!=\!\!$ $\frac{\partial_{0}\rho}{(\partial_{0}\rho)^{2}\!-\!(\partial_{\zeta}\rho)^{2}}\Big(2\partial_{\zeta}^{2}\rho+\frac{\rho}{2}[(\partial_{0}\phi)^{2}+(\partial_{\zeta}\phi)^{2}]\Big)$
- $\frac{\partial_{\zeta}\rho}{(\partial_{0}\rho)^{2}\!-\!(\partial_{\zeta}\rho)^{2}}\big(2\partial_{\zeta}\partial_{0}\rho+\rho\partial_{0}\phi\partial_{\zeta}\phi\big)$
$F_{1}^{\rm pt}(\rho,\sigma)$ $\!\!=\!\!$ $\frac{\partial_{0}\rho}{(\partial_{0}\rho)^{2}\!-\!(\partial_{\zeta}\rho)^{2}}\big(2\partial_{\zeta}\partial_{0}\rho+\rho\partial_{0}\phi\partial_{\zeta}\phi\big)$ (2.9)
- $\frac{\partial_{\zeta}\rho}{(\partial_{0}\rho)^{2}\!-\!(\partial_{\zeta}\rho)^{2}}\Big(2\partial_{\zeta}^{2}\rho+\frac{\rho}{2}[(\partial_{0}\phi)^{2}+(\partial_{\zeta}\phi)^{2}]\Big)\,.$
The integrability condition mentioned after Eq.(2.5) specializes to $\partial_{\zeta}F_{0}^{\rm pt}=\partial_{0}F_{1}^{\rm pt}$, and can be checked to hold on account of the first two equations in (2.8). Note that the $\phi$ field equation is through $\rho$ explicitly $(x_{0},\zeta)$-dependent; in particular, it is not conformally invariant. The other vestige of the gravitational origin of the system then is that in order for $\sigma$ to be spatially periodic, one must have
$\int_{0}^{2\pi}\frac{d\zeta}{2\pi}F_{1}^{\rm pt}(\rho,\phi)\stackrel{{\scriptstyle\displaystyle{!}}}{{=}}0\,.$ (2.10)
VD Gowdy system. The velocity dominated (VD) Gowdy system is normally introduced only on the level of the gauge fixed field equations [3, 5]. It can, however, be viewed as a (electric) Carroll-type gravity theory [22, 23] in its own right with Lagrangian action
$S^{{\textsc{lvd}}}=\frac{1}{\lambda_{{\textsc{n}}}}\int_{t_{i}}^{t_{f}}\!\!dx^{0}\!\!\int_{0}^{2\pi}\!\!dx^{1}\Big\{\!-\frac{1}{n}e_{0}(\varrho)e_{0}(\varsigma)+\frac{\rho}{2n}e_{0}(\varphi)^{2}\Big\},$ (2.11)
where $\varrho,\varsigma,\varphi$ are the counterparts $\rho,\sigma,\phi$ in the VD system. Strictly speaking different symbols should be used also for lapse and shift, for readability’s sake we retain the original notation. Observe that in (2.11) spatial points are only coupled through the Lie derivative term in $e_{0}$. Correspondingly, there is a Diffeomorphism constraint $\delta S^{{\textsc{lvd}}}/\delta s=-{\cal H}_{1}^{{\textsc{lvd}}}$ of the same form as in the full Gowdy system, while the Hamiltonian constraint $\delta S^{{\textsc{lvd}}}/\delta n=-{\cal H}_{0}^{{\textsc{lvd}}}$ simplifies. The constraints can be solved for $e_{0}(\varsigma)$ and $\partial_{1}\varsigma$ and determine $\varsigma$ as a functional of the on-shell $\varrho,\varphi$, i.e. $\varsigma=\varsigma[\varrho^{\rm on},\varphi^{\rm on}]$. A more detailed exposition of the VD Hamiltonian action and its gauge symmetries is relegated to Appendix A. The transition to a reduced phase space formulation proceeds in parallel to the previous discussion.
For comparison’s sake, we note the independent field equations in the proper time gauge (2.6), (2.7):
$\partial_{0}(\varrho\partial_{0}\varphi)=0\,,\quad\partial_{0}^{2}\varrho=0\,$, (2.12)
$\partial_{0}\varsigma=\frac{\varrho}{2\partial_{0}\varrho}(\partial_{0}\varphi)^{2}\,,\quad\partial_{\zeta}(\varsigma+2\ln n)=\frac{1}{\partial_{0}\varrho}\big(\partial_{0}\partial_{\zeta}\varrho+\varrho\partial_{0}\varphi\partial_{\zeta}\varphi\big)-\frac{\varrho\partial_{\zeta}\varrho}{2(\partial_{0}\varrho)^{2}}(\partial_{0}\varphi)^{2}\,.$
Here the counterpart of (2.10) needs to be imposed in order to ensure that $\varsigma$ is spatially periodic.
Matched foliations. So far we kept $\rho$ as a field, which in the reduced phase space formulation sets a background geometry of sorts for the dynamical field $\phi$ and the induced $\sigma$. It is worth noting that $\rho$ meets the criteria for a temporal function in Lorentzian geometry: it is a strictly positive scalar function which has an everywhere timelike gradient $d\rho=\partial_{\mu}\rho dx^{\mu}$ such that the associated vector field $\gamma^{\mu\nu}\partial_{\nu}\rho$ is past pointing. This means, the level surfaces $\rho={\rm const}.$ define a foliation of the two-dimensional Lorentzian manifold $(\mathbb{R}\times S^{1},\gamma)$. Specifically, in proper time gauge one can use the equation of motion $(\partial_{0}-\partial_{\zeta})(\partial_{0}+\partial_{\zeta})\rho=0$ and the spatial periodicity requirement to conclude that $\rho$ must be quasi-periodic in $x^{0}$, i.e. $\rho(x^{0}+n2\pi,\zeta)=\rho(x^{0},\zeta)+n\rho_{0}$, with $\rho_{0}=\int_{0}^{2\pi}dx(\partial_{0}\rho)$. Positivity then requires either $x^{0}>0,\rho_{0}>0$ or $x^{0}<0,\rho_{0}<0$. Choosing the former and taking into account $(\partial_{0}\rho)^{2}>(\partial_{\zeta}\rho)^{2}>0$ one sees that $\rho(x^{0},\zeta)$ is strictly increasing in $x^{0}$, pointwise in $\zeta$. Its level surfaces $\rho(x^{0},\zeta)={\rm const}.$ define a foliation of the half cylinder $\mathbb{R}_{+}\times S^{1}$.
On the other hand, the defining relation for the lapse in terms of an ADM temporal function $T$ specializes to $\sqrt{\gamma}\gamma^{\mu\nu}\partial_{\mu}T\partial_{\nu}T=-1/n$ for the metric (2.2). This is satisfied for the choice $T=x^{0}$, and the level surfaces $x^{0}={\rm const}$ set the standard ADM foliation to which the coordinatization of the fields in terms of $(x^{0},\zeta)$ refers. In most of the Gowdy literature the $\rho$-foliation and the standard ADM foliation are identified by stipulating [^2] As a metric component $\rho$ is dimensionless. For simplicity, we suppress a dimensionful conversion factor in this and subsequent identifications.
$\rho=x^{0}=:t>0\,.$ (2.13)
This is especially natural in the reduced phase space formulation, where $\rho$ merely sets a background for the $\phi,\sigma$ evolution. Technically, (2.13) should not be regarded as a gauge choice complementing the proper time gauge. Indeed, for an off shell $\rho$ the condition $n\partial_{1}\rho=0$ is compatible with the residual gauge symmetry of the proper time gauge: $\delta_{\epsilon^{1}}\partial_{1}\rho|_{\partial_{1}\rho=0}=\partial_{1}[\delta_{\epsilon^{1}}\rho]|_{\partial_{1}\rho=0}=0$. One can still take $\rho=\rho(x^{0})$ to be an arbitrary (positive, strictly increasing) function of $x^{0}$. The identification (2.13) picks one of them. We stress this because in the discussion of Dirac observables, the full phase space must be used, including the $(\rho,\pi^{\rho})$ canonical pair as off-shell dynamical degree of freedom, see Appendix B. Once Dirac observables ${\cal O}(\theta)$ have successfully constructed the choice (2.13) eliminates the $(\rho,\pi^{\rho})$ canonical pair and reduces them to (semilocal) charges which are conserved on-shell, $\partial_{t}{\cal O}(\theta)=0$.
A similar discussion applies to the VD Gowdy system. To avoid a detour into Carroll geometry, we consider directly the properties of $\varrho$ in proper time gauge. The conditions $\partial_{0}^{2}\varrho=0$, $\varrho>0$, and the spatial periodicity fix $\varrho(x^{0},\zeta)=\varrho_{1}(\zeta)x^{0}+\varrho_{0}(\zeta)$. Taking $x^{0}>0$ one needs $\varrho_{0},\varrho_{1}$ to be periodic and positive. Again, $\varrho$ plays the role of a local time function, now linear in $x^{0}$. The counterpart of (2.13) is the choice $\varrho=x^{0}>0$, which also here it is not a gauge choice on top of the proper time gauge. Together, the matched foliations condition identifies three temporal functions, $\rho$, $\varrho$, and the ADM one:
$\rho=x^{0}=t=\varrho>0\,.$ (2.14)
This sets a shared coordinate time $t$ in which the solutions of the full and the VD Gowdy system evolve.
2.2 The AVD property
In the shared coordinate time $\rho=\varrho=t>0$ both sets of field equations (LABEL:evolpt) and (LABEL:VDevolpt) can essentially be solved in closed form. For (LABEL:VDevolpt) the general solution is
$\varphi(t,\zeta)=v(\zeta)\ln(t/t_{0})+\varphi_{0}(\zeta)\,,\quad\varsigma(t,\zeta)=\ln(t/t_{0})\frac{1}{2}v(\zeta)^{2}+\varsigma_{0}(\zeta)\,$,
$\varsigma_{0}(\zeta)=-2\ln(\frac{n(\zeta)}{n(\zeta_{0})})+\int_{\zeta_{0}}^{\zeta}\!d\zeta^{\prime}v(\zeta^{\prime})\partial_{\zeta^{\prime}}\varphi_{0}(\zeta^{\prime})\,.$ (2.15)
Note that this carries a largely arbitrary $\zeta$ dependence merely through the choice of initial conditions at $t=t_{0}$. Further, $\varsigma$ is determined by $\varphi$ up to a constant. The solution of (LABEL:evolpt) in $\rho=t$ time proceeds via spatial Fourier decomposition of $\partial_{t}(t\partial_{t}\phi)=t\partial_{\zeta}^{2}\phi$ and leads to Bessel functions with index zero. Upon Fourier synthesis, one can insert the solution into the constraint equations and integrate them to find $\sigma$. Needless to say, this is only possible in the polarized case. In the non-polarized Gowdy system the evolution equations are nonlinear partial differential equations and despite the existence of a Lax pair they cannot be integrated in closed form. On the other hand, the VD Gowdy system admits a simple explicit solution even in the non-polarized case. This pattern underlines why the AVD property is interesting and nontrivial.
Foreshadowing much of the subsequent developments the AVD property for polarized Gowdy cosmologies has been proven by Isenberg and Moncrief [3], see Thm. III.5 for the $T^{3}$ topology considered here. [^3] The notational correspondence to [3] is: $\phi\mapsto W,\,\tilde{\sigma}+2\ln n=\sigma+\ln(n^{2}\sqrt{t})\mapsto 2a,\,\ln(t/t_{0})=\tau\mapsto-(\tau\!-\!\tau_{0})$. Schematically, the result asserts the following: Given initial data $\phi_{0}(\zeta),\pi_{0}(\zeta),\sigma_{0}(\zeta)$ at some $t_{0}>0$ there exists a solution (LABEL:VDsol) of the VD system and correction terms $\Phi(t,\zeta),\Sigma(t,\zeta)$ such that
$\phi(t,\zeta)=\varphi(t,\zeta)+\Phi(t,\zeta),\quad\sigma(t,\zeta)=\varsigma(t,\zeta)+\Sigma(t,\zeta)\,$,
$|\partial_{\zeta}^{k}\Phi(t,\zeta)|,|t\partial_{t}\partial_{\zeta}^{k}\Phi(t,\zeta)|,|\partial_{\zeta}^{k}\Sigma(t,\zeta)|\leq c(1\!+\!\ln^{2}(t/t_{0}))t^{2}\,$, (2.16)
(with $c>0$ and some finite differentiation order $k$) is a solution of the full Gowdy field equations (LABEL:evolpt) with initial data $\phi(t_{0},\zeta)=\phi_{0}(\zeta)$, $(t\partial_{t}\phi)(t_{0},\zeta)=\pi_{0}(\zeta)$, $\sigma(t,\zeta)=\sigma_{0}(\zeta)$. The same holds if instead of initial data a non-exceptional solution (LABEL:VDsol) for the VD system is prescribed.
In other words, all solutions of the full Gowdy system approach a VD solution near the Big Bang. Conversely, almost every solution of the VD system can be lifted to a solution of the full Gowdy system via (LABEL:AVD).
The origin of the exceptional points can be understood from the blow-up behavior of the curvature scalars. Considering ${\cal R}_{2}:=R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}$, the result of [3], Thm. IV.1, is that the rate of blow up for $t\rightarrow 0^{+}$ is determined solely by the coefficient $v(\zeta)$ in the velocity dominated solution (LABEL:VDsol). Specifically,
${\cal R}_{2}(t,\zeta_{0})=\left\{\begin{array}[]{cc}O(t^{-3-v(\zeta_{0})^{2}})&\mbox{if}\;\;v(\zeta_{0})^{2}\neq 1\,,\\[5.69054pt] O(t^{-2})&\mbox{if}\;\;v(\zeta_{0})^{2}=1\,,\;(\partial_{\zeta}v)(\zeta_{0})\neq 0\,,\\[5.69054pt] O(\ln^{2}t/t_{0})&\mbox{if}\;\;v(\zeta_{0})^{2}=1\,,\;(\partial_{\zeta}v)(\zeta_{0})=0\,,\;\;(\partial_{\zeta}^{2}v)(\zeta_{0})\neq 0\,.\end{array}\right$. (2.20)
This means, a form of “cosmic censorship” holds: except for exceptional cases with ‘fine tuned’ data, the polarized Gowdy cosmologies have a Big Bang singularity. Moreover, they do not admit an analytic extension beyond the singularity to include non-globally hyperbolic (‘acausal’) regions.
2.3 Fourier decomposition and canonical quantization
Returning to (LABEL:AVD), since $\sigma$ and $\varsigma$ are determined by $\phi$ and $\varphi$, respectively, the key aspect is the one-to-one correspondence between $\phi(t,\zeta)$ and $\varphi(t,\zeta)$. As their defining wave equations are linear, the problem can naturally be recast in terms of the spatial Fourier transforms. Our conventions for the Fourier transform are
$f(\tau,\zeta)$ $\!\!=\!\!$ $\sum_{n\in\mathbb{Z}}f_{n}(\tau)e^{in\zeta}\,,\quad\;\;f_{n}(\tau)=\int_{0}^{2\pi}\!\frac{d\zeta}{2\pi}\,e^{-in\zeta}f(\tau,\zeta)\,.$ (2.21)
For the scalar field, we use the mode expansion
$\phi(\tau,\zeta)=T_{0}(\tau)a_{0}+T_{0}(\tau)^{*}a_{0}^{*}+\sum_{n\neq 0}\big\{T_{n}(\tau)e^{in\zeta}a_{n}+T_{n}(\tau)^{*}e^{-in\zeta}a_{n}^{*}\big\}\,$, (2.22)
where the time variable is taken to be $\tau=\ln t/t_{0}$, with $t\in\mathbb{R}_{+}$ transcribing to $\tau\in\mathbb{R}$, and the Big Bang located at $\tau=$-$\infty$. Classically, the $a_{n},a_{n}^{*}$ are just complex parameters; later, they will be promoted to Fock space operators, see [11, 12, 13, 14]. Applying a spatial Fourier transform to the wave equation for $\phi$ in Eq. (LABEL:evolpt) for $\rho=t$ yields
$\partial_{\tau}^{2}+e^{2\tau}(t_{0}n)^{2}]T_{n}(\tau)=0\,,\quad(\partial_{\tau}T_{n}){T_{n}}^{*}-(\partial_{\tau}T_{n})^{*}T_{n}=-i\,.$ (2.23)
When using the dimensionless $\tau$ time variable only the dimensionless combination $t_{0}n$ occurs. For notational simplicity we will just write $n$ for this, with $t_{0}$ tacit whenever demanded for dimensionality reasons. In the second equation we imposed a Wronskian normalization condition based on the fact that the quantity on the left hand side is constant in $\tau$. The general solution can be written in terms of Hankel functions in the form
$T_{n}(\tau)$ $\!\!=\!\!$ $\frac{\sqrt{\pi}}{2}\lambda_{n}H_{0}^{(2)}(|n|e^{\tau})+\frac{\sqrt{\pi}}{2}\mu_{n}H_{0}^{(1)}(|n|e^{\tau})\,,\quad\,|\lambda_{n}|^{2}-|\mu_{n}|^{2}=1\,,\quad n\neq 0\,$,
$T_{0}(\tau)$ $\!\!=\!\!$ $-\frac{i}{\sqrt{\pi}}(\tilde{\lambda}_{0}-\tilde{\mu}_{0})\tau+\frac{\sqrt{\pi}}{2}(\tilde{\lambda}_{0}+\tilde{\mu}_{0})\,,\quad|\tilde{\lambda}_{0}|^{2}-|\tilde{\mu}_{0}|^{2}=1\,.$ (2.24)
We assume throughout that $\lambda_{n}=\lambda_{-n},\mu_{n}=\mu_{-n}$ are bounded in $|n|$ and take this as part of the definition of a ‘state’ via the two-point function (LABEL:Wsdef) below. An alternative parameterization is in terms of initial data
$z_{n}(\tau_{0}):=T_{n}(\tau_{0}),\quad w_{n}(\tau_{0}):=(\partial_{\tau}T_{n})(\tau_{0}).$ (2.25)
Displaying also a potential $\tau_{0}$-dependence of the Bogoliubov parameters $\lambda_{n}=\lambda_{n}(\tau_{0})$ and $\mu_{n}=\mu_{n}(\tau_{0})$, both parameterizations are for $n\neq 0$ related by
$\begin{pmatrix}z_{n}(\tau_{0})\\ w_{n}(\tau_{0})\end{pmatrix}=H(|n|e^{\tau_{0}})\begin{pmatrix}\lambda_{n}(\tau_{0})\\ \mu_{n}(\tau_{0})\end{pmatrix}\,,\quad H(x):=\frac{\sqrt{\pi}}{2}\begin{pmatrix}H_{0}^{(2)}(x)&H_{0}^{(1)}(x)\\[4.2679pt] -xH_{1}^{(2)}(x)&-xH_{1}^{(1)}(x)\end{pmatrix}$, (2.26)
with $\det H(x)=i$.
Upon canonical quantization the $a_{n}$ and $a_{n}^{*}$ are promoted to annihilation and creation operators normalized according to $a_{n},a_{m}^{*}]=\delta_{n,m}$. Each choice of Wronskian normalized solution $T_{n}(\tau)$, corresponds to a different decomposition (2.22), i.e. to different Fock space operators $a_{n}=a_{n}^{T}$, $a_{n}^{*}=(a_{n}^{T})^{*}$, and we associate a Fock vacuum to some such choice via
$a_{n}|0_{T}\rangle=0\,,\quad n\in\mathbb{Z}.$ (2.27)
For the velocity dominated system, we adapt (2.22) to
$\varphi(\tau,\zeta)=T_{0}(\tau)a_{0}+T_{0}(\tau)^{*}a_{0}^{*}+\sum_{n\neq 0}\big\{\mathfrak{t}_{n}(\tau)e^{in\zeta}a_{n}+\mathfrak{t}_{n}(\tau)^{*}e^{-in\zeta}a_{n}^{*}\big\}\,$, (2.28)
where the defining relation for $\mathfrak{t}_{n}(\tau)$ is simply $\partial_{\tau}^{2}\mathfrak{t}_{n}=0$. We parameterize the general Wronskian normalized solution as
$\mathfrak{t}_{n}(\tau)=-\frac{i}{\sqrt{\pi}}(\tilde{\lambda}_{n}-\tilde{\mu}_{n})\tau+\frac{\sqrt{\pi}}{2}(\tilde{\lambda}_{n}+\tilde{\mu}_{n})=\tilde{w}_{n}(\tau_{0})(\tau-\tau_{0})+\tilde{z}_{n}(\tau_{0})\,.$ (2.29)
Guided by (LABEL:AVD) one will describe the quantized velocity dominated scalar $\varphi(\tau,\zeta)$ by the same set of creation/annihilation operators in Fourier space as $\phi(\tau,\zeta)$, but use as the classical coefficient functions those $\mathfrak{t}_{n}$ of the velocity dominated system. We highlight the concomitant identification of the vacua
$|0_{T}\rangle=|0_{\mathfrak{t}}\rangle\,.$ (2.30)
For $n=0$, we identify $T_{0}(\tau)=\mathfrak{t}_{0}(\tau)$. The Wronskian normalization condition amounts respectively to $|\tilde{\lambda}_{n}|^{2}-|\tilde{\mu}_{n}|^{2}=1$ and $(\tilde{w}_{n}\tilde{z}_{n}^{*}-\tilde{w}_{n}^{*}\tilde{z}_{n})(\tau_{0})=-i$. For later use, we express this in parallel to (2.26) as
$\begin{pmatrix}\tilde{z}_{n}(\tau_{0})\\ \tilde{w}_{n}(\tau_{0})\end{pmatrix}=\tilde{H}(\tau_{0})\begin{pmatrix}\tilde{\lambda}_{n}(\tau_{0})\\ \tilde{\mu}_{n}(\tau_{0})\end{pmatrix}\,,\quad\tilde{H}(\tau_{0}):=\begin{pmatrix}h(\tau_{0})&h(\tau_{0})^{*}\\[5.69054pt] -i/\sqrt{\pi}&i/\sqrt{\pi}\end{pmatrix}\,$, (2.31)
where $h(\tau):=\sqrt{\pi}/2-i\tau/\sqrt{\pi}$.
2.4 Two-point functions of gauge-fixed Dirac observables
In Appendix B, a one-parameter family of off-shell Dirac observables for the polarized $T^{3}$ Gowdy cosmologies is constructed, which strongly Poisson commutes with the constraints. In proper time gauge, with $\lambda_{{\textsc{n}}}=1$, and $\rho=t$, $\tilde{\rho}=\zeta$ coordinates the on-shell Dirac observables (B) can after integration-by-parts be written as
${\cal O}(\theta)\!$ $\!\!=\!\!$ $\!\int_{0}^{2\pi}\frac{d\zeta}{2\pi}q(\tau,\zeta;\theta)\,,\quad{\rm Im}\theta\neq 0\,$, (2.32)
$q(\tau,\zeta;\theta)\!$ $\!\!:=\!\!$ $\!c_{0}(\zeta\!+\!\theta,e^{\tau})\partial_{\tau}\phi+e^{2\tau}\partial_{\zeta}c_{1}(\zeta\!+\!\theta,e^{\tau})\phi=\jmath_{0}^{{\textsc{p}}}(e^{\tau},\zeta;\theta)+\partial_{\zeta}[e^{2\tau}c_{1}(\zeta\!+\!\theta,e^{\tau})\phi]\,$,
where we use $\tau=\ln t$ as time variable and omit the superscript “on”. By construction, the $c_{0},c_{1}$ are $2\pi$-periodic functions in $\zeta$ and are such that $\partial_{\tau}{\cal O}(\theta)=0$ holds, subject to the evolution equations in the same specialization. This conservation is a consequence of the stronger gauge invariance of the ${\cal O}(\theta)$ in (B.21), but in the fully gauge-fixed setting convenient for (canonical) quantization it is the main indicator for the observable property. In relativistic quantum field theories conserved charges normally have to be regularized in order to give rise to well-defined (unbounded) operators. Remarkably, this is not necessary here. Inserting the mode expansion (2.22) for $\phi$ a well-defined unbounded operator on Fock space arises as long as ${\rm Im}\theta\neq 0$. It is given by (B) with the $a_{n}^{*},a_{n}$ read as creation, annihilation operators. The two-point function in the vacuum (2.27) is simply
$\langle 0_{T}|{\cal O}(\theta){\cal O}(\theta^{\prime})|0_{T}\rangle=\frac{1}{\pi}\sum_{n\geq 1}|\lambda_{n}\!-\!\mu_{n}|^{2}\,n^{2}e^{\pm in(\theta-{\theta^{\prime}}^{*})}\,,\quad\pm{\rm Im}\theta>0,\,\pm{\rm Im}\theta^{\prime}>0\,$, (2.33)
where $\lambda_{n},\mu_{n}$ are the Bogoliubov parameters entering via (2.3). Clearly, the sum in (2.33) is rapidly convergent for all $\lambda_{n},\mu_{n}$ bounded in $|n|$. In Section 4 we shall argue that physically viable states have in fact parameters obeying $\lambda_{n}\rightarrow 1$, $\mu_{n}\rightarrow 0$, as $n\rightarrow\infty$. The two-point function (2.33) is manifestly $\tau$ independent and turns to coincide exactly with its counterpart computed in the velocity dominated system, see (B). While conceptually satisfying this renders (2.33) unsuited to study AVD.
Better suited for this purpose is the two-point functions of the integrands $q(\tau,\zeta;\theta)$ in the Fock vacuum (2.27). As often, it is useful to consider the symmetric and anti-symmetric parts separately. These can be expressed in matrix form as follows
$\langle 0_{T}|q(\tau,\zeta;\theta)q(\tau^{\prime},\zeta^{\prime};\theta^{\prime})+q(\tau^{\prime},\zeta^{\prime};\theta^{\prime})q(\tau,\zeta;\theta)|0_{T}\rangle$ (2.34)
$\quad=\big(e^{2\tau}\partial_{\zeta}c_{1}(\zeta\!+\!\theta,e^{\tau}),c_{0}(\zeta\!+\!\theta,e^{\tau})\big){\cal W}^{s}(\tau,\tau^{\prime},\zeta-\zeta^{\prime})\begin{pmatrix}e^{2\tau^{\prime}}\partial_{\zeta^{\prime}}c_{1}(\zeta^{\prime}\!+\!\theta^{\prime},e^{\tau^{\prime}})\\[5.69054pt] c_{0}(\zeta^{\prime}\!+\!\theta^{\prime},e^{\tau^{\prime}})\end{pmatrix}$,
where the following symmetric matrix two-point function enters
${\cal W}^{s}(\tau,\tau^{\prime},\zeta-\zeta^{\prime})$ $:=$ $\begin{pmatrix}1&\partial_{\tau^{\prime}}\\[5.69054pt] \partial_{\tau}&\partial_{\tau}\partial_{\tau^{\prime}}\end{pmatrix}W^{s}(\tau,\tau^{\prime},\zeta-\zeta^{\prime})\,$,
$W^{s}(\tau,\tau^{\prime},\zeta-\zeta^{\prime})$ $:=$ $\langle 0_{T}|\phi(\tau,\zeta)\phi(\tau^{\prime},\zeta^{\prime})+\phi(\tau^{\prime},\zeta^{\prime})\phi(\tau,\zeta)|0_{T}\rangle\,.$ (2.35)
In the parameterization (2.3) this gives
$W^{s}(\tau,\tau^{\prime},\zeta-\zeta^{\prime})=W_{0}(\tau,\tau^{\prime})+2\sum_{n\geq 1}\cos n(\zeta\!-\!\zeta^{\prime})W_{n}(\tau,\tau^{\prime}),$ (2.36)
$W_{0}(\tau,\tau^{\prime})=\frac{2}{\pi}|\tilde{\lambda}_{0}\!-\!\tilde{\mu}_{0}|^{2}\tau\tau^{\prime}-i(\tilde{\lambda}_{0}\tilde{\mu}_{0}^{*}\!-\!\tilde{\lambda}_{0}^{*}\tilde{\mu}_{0})(\tau\!+\!\tau^{\prime})+\frac{\pi}{2}|\tilde{\lambda}_{0}+\tilde{\mu}_{0}|^{2}\,$,
$W_{n}(\tau,\tau^{\prime})=\frac{\pi}{2}\big(|\lambda_{n}|^{2}+|\mu_{n}|^{2}\big)\big(J_{0}(|n|e^{\tau})J_{0}(|n|e^{\tau^{\prime}})+Y_{0}(|n|e^{\tau^{\prime}})Y_{0}(|n|e^{\tau})\big)$
$\quad+\!\frac{\pi}{2}(\lambda_{n}\mu_{n}^{*}+\lambda_{n}^{*}\mu_{n})\big(J_{0}(|n|e^{\tau})J_{0}(|n|e^{\tau^{\prime}})-Y_{0}(|n|e^{\tau^{\prime}})Y_{0}(|n|e^{\tau})\big)$
$\quad-\!\frac{\pi}{2}i(\lambda_{n}\mu_{n}^{*}-\lambda_{n}^{*}\mu_{n})\big(J_{0}(|n|e^{\tau})Y_{0}(|n|e^{\tau^{\prime}})+Y_{0}(|n|e^{\tau})J_{0}(|n|e^{\tau^{\prime}})\big)\,,\quad n\neq 0\,.$
The anti-symmetric part of $q$’s two-point function is independent of the choice of vacuum state $|0_{T}\rangle$ and is just given by the commutator, i.e. $\langle 0_{T}|q(\tau,\zeta;\theta)q(\tau^{\prime},\zeta^{\prime};\theta)-q(\tau^{\prime},\zeta^{\prime};\theta)q(\tau,\zeta;\theta)|0_{T}\rangle=[q(\tau,\zeta;\theta),q(\tau^{\prime},\zeta^{\prime};\theta)]$. In matrix form
$i[q(\tau,\zeta;\theta),q(\tau^{\prime},\zeta^{\prime};\theta^{\prime})]$ (2.37)
$\quad=\big(e^{2\tau}\partial_{\zeta}c_{1}(\zeta\!+\!\theta,e^{\tau}),c_{0}(\zeta\!+\!\theta,e^{\tau})\big)D(\tau,\tau^{\prime},\zeta-\zeta^{\prime})\begin{pmatrix}0&-1\\[5.69054pt] 1&0\end{pmatrix}\begin{pmatrix}e^{2\tau^{\prime}}\partial_{\zeta^{\prime}}c_{1}(\zeta^{\prime}\!+\!\theta^{\prime},e^{\tau^{\prime}})\\[5.69054pt] c_{0}(\zeta^{\prime}\!+\!\theta^{\prime},e^{\tau^{\prime}})\end{pmatrix}$.
The matrix $D$ is defined in parallel to (2.4) but with a ‘symplectic unit’ matrix taken out for later convenience. Explicitly
$\!\!\!\!\!\!\!\!\!\!D(\tau,\tau^{\prime},\zeta-\zeta^{\prime})\begin{pmatrix}0&-1\\[5.69054pt] 1&0\end{pmatrix}$ $:=$ $\begin{pmatrix}1&\partial_{\tau^{\prime}}\\[5.69054pt] \partial_{\tau}&\partial_{\tau}\partial_{\tau^{\prime}}\end{pmatrix}\Delta(\tau,\tau^{\prime},\zeta-\zeta^{\prime})\,$,
$\!\!\!\!\!\!\!\!\!\!\Delta(\tau,\tau^{\prime},\zeta-\zeta^{\prime})$ $:=$ $i\langle 0_{T}|\phi(\tau,\zeta)\phi(\tau^{\prime},\zeta^{\prime})-\phi(\tau^{\prime},\zeta^{\prime})\phi(\tau,\zeta)|0_{T}\rangle\,.$ (2.38)
We recall from Appendix B that $c_{0}(\zeta\!+\!\theta,e^{\tau})$, $c_{1}(\zeta\!+\!\theta,e^{\tau})$ are explicitly known and given by convergent power series in $e^{2\tau}$, see (B). The computation of the two-point functions (LABEL:diractwopt2), (LABEL:diractwopt3) therefore amounts to the evaluation of the matrix two-point functions ${\cal W}^{s}$ and $D$. Our primary interest will lie in their dependence on $\tau,\tau^{\prime}$.
3.  Time evolution of two-point functions
The time evolution of the Heisenberg picture field operators and Fock space operators can be described by convolution with a symplectic matrix kernel that is defined in terms of the commutator function; see (LABEL:opevol). This in turn gives rise to an initial value parameterization for the symmetric part of the two-point functions; see (LABEL:twopmat2), (LABEL:twopmat3). The bonus property that the time evolution of (slightly modified) Heisenberg picture field operators can be unitarily implemented on Fock space [13] will not be needed. For simplicity, we set $\lambda_{{\textsc{n}}}=1$ throughout this section.
3.1 Time evolution of field operators
The solution of the initial value problem of Heisenberg picture field operators can concisely be expressed in terms of the commutator function $\Delta(\tau,\tau^{\prime},\zeta-\zeta^{\prime})$ defined in (2.4). Its Fourier components $\Delta_{n}(\tau,\tau^{\prime})$ are real valued and can be characterized by the following relations: $\partial_{\tau}^{2}+n^{2}e^{2\tau}]\Delta_{n}(\tau,\tau^{\prime})=0$, $\Delta_{n}(\tau,\tau^{\prime})=-\Delta_{n}(\tau^{\prime},\tau)$, $\partial_{\tau}\Delta_{n}(\tau,\tau^{\prime})|_{\tau^{\prime}=\tau}=~1$. Merely on account of these properties one has
$\begin{pmatrix}\phi_{n}(\tau)\\[4.2679pt] \pi_{n}(\tau)\end{pmatrix}=D_{n}(\tau,\tau_{0})\begin{pmatrix}\phi_{n}(\tau_{0})\\[4.2679pt] \pi_{n}(\tau_{0})\end{pmatrix},\quad D_{n}(\tau,\tau_{0}):=\begin{pmatrix}-\partial_{\tau_{0}}\Delta_{n}(\tau,\tau_{0})&\Delta_{n}(\tau,\tau_{0})\\[4.2679pt] -\partial_{\tau}\partial_{\tau_{0}}\Delta_{n}(\tau,\tau_{0})&\partial_{\tau}\Delta_{n}(\tau,\tau_{0})\end{pmatrix}$, (3.1)
for all $n\in\mathbb{Z}$, where $\pi_{n}(\tau):=\partial_{\tau}\phi_{n}(\tau)$. We take $\Delta_{0}(\tau,\tau_{0}):=\tau-\tau_{0}$. If the data are propagated in two time steps, from $\tau_{0}$ to $\tau_{1}$ and then from $\tau_{1}$ to $\tau$, uniqueness of the solution implies consistency conditions on the evolution matrix in (3.1). The four conditions arising consist of a basic one (the 1,2 matrix component) and derivatives thereof. The basic identity reads
$\partial_{\tau_{1}}\Delta_{n}(\tau,\tau_{1})\Delta_{n}(\tau_{1},\tau_{0})+\Delta_{n}(\tau,\tau_{1})\partial_{\tau_{1}}\Delta_{n}(\tau_{1},\tau_{0})=\Delta_{n}(\tau,\tau_{0})\,$, (3.2)
and again holds only on account of the above properties of $\Delta_{n}$. The matrix $D_{n}$ obeys $D_{n}(\tau,\tau_{1})D_{n}(\tau_{1},\tau_{0})=D_{n}(\tau,\tau_{0})$ by (3.2). In particular, $D_{n}(\tau,\tau_{0})^{-1}=D_{n}(\tau_{0},\tau)$ and by interpreting the adjoint in terms of the inverse it is also symplectic.
Upon Fourier synthesis the matrix kernel $D(\tau,\zeta;\tau^{\prime},\zeta^{\prime})$ governs the time evolution of the Heisenberg picture field operators, see (LABEL:opevol) below. It is manifestly built from the time derivatives of the basic commutator function $\Delta(\tau,\zeta;\tau^{\prime},\zeta^{\prime})$. However, the inverse Fourier transform of $\Delta_{n}$ can not be evaluated in closed form and even certain qualitative properties are masked. Explicitly, we seek to evaluate
$\Delta(t,t^{\prime},\zeta-\zeta^{\prime})$ $\!\!=\!\!$ $\ln(t/t^{\prime})+2\sum_{n\geq 1}\cos n(\zeta-\zeta^{\prime})\Delta_{n}(t,t^{\prime})\,$,
$\Delta_{n}(t,t^{\prime})$ $\!\!=\!\!$ $\frac{\pi}{2}\big\{Y_{0}(|n|t)J_{0}(|n|t^{\prime})-J_{0}(|n|t)Y_{0}(|n|t^{\prime})\big\}\,$, (3.3)
where for notational convenience we revert to $t,t^{\prime}>0$ as time variable. When the spatial sections are diffeomorphic to $\mathbb{R}$ the commutator function $\Delta_{\mathbb{R}}(t,t^{\prime},\zeta-\zeta^{\prime})$ can be shown analytically to vanish for spacelike distances, $|\zeta-\zeta^{\prime}|>|t-t^{\prime}|$ (despite the time dependent background). We state without derivation the following link to the commutator function (3.1) on the cylinder needed here:
$2\pi\sum_{n\in\mathbb{Z}}\Delta_{\mathbb{R}}(t,t^{\prime},\zeta-\zeta^{\prime}+2\pi n)=\Delta(t,t^{\prime},\zeta-\zeta^{\prime})\,$, (3.4)
where the sum is finite for each fixed set of arguments. This relation implies that within each periodicity interval $0<|\zeta-\zeta^{\prime}|<2\pi$ also the cylinder commutator function vanishes exactly for $|\zeta-\zeta^{\prime}|>|t-t^{\prime}|$.
A brute force implementation of the sum (3.1) turns out to be numerically unstable. We therefore perform the sums that govern its qualitative behavior analytically and treat only the rest numerically. By the exact anti-symmetry in $t,t^{\prime}$ we may restrict attention to $t\geq t^{\prime}>0$. For large $n>1$ the terms behave like
$\frac{\pi}{2}\big(Y_{0}(nt)J_{0}(nt^{\prime})-Y_{0}(nt^{\prime})J_{0}(nt)\big)$
$\quad=:\frac{1}{\sqrt{tt^{\prime}}}\frac{\sin n(t\!-\!t^{\prime})}{n}+\frac{t\!-\!t^{\prime}}{8(tt^{\prime})^{3/2}}\frac{\cos n(t\!-\!t^{\prime})}{n^{2}}+\Delta^{\rm sub2}_{n}(t,t^{\prime})\,.$ (3.5)
Here, the ‘subtracted’ summands $\Delta_{n}^{\rm sub2}(t,t^{\prime})$ are pointwise $O(1/n^{3})$. Then (3.1) can be rewritten as
$\Delta(t,t^{\prime},\zeta-\zeta^{\prime})=\ln(t/t^{\prime})+2\sum_{n\geq 1}\cos n(\zeta\!-\!\zeta^{\prime})\Delta_{n}^{\rm sub2}(t,t^{\prime})$
$\quad+2\sum_{n\geq 1}\cos n(\zeta\!-\!\zeta^{\prime})\Big[\frac{1}{\sqrt{tt^{\prime}}}\frac{\sin n(t\!-\!t^{\prime})}{n}+\frac{t\!-\!t^{\prime}}{8(tt^{\prime})^{3/2}}\frac{\cos n(t\!-\!t^{\prime})}{n^{2}}\Big]\,.$ (3.6)
The infinite sums in the second line can be performed exactly:
$2\sum_{n=1}^{\infty}\cos nx\,\frac{\sin nT}{n}$ $\!\!=\!\!$ $-T+\frac{\pi}{2}\sum_{n\in\mathbb{Z}}\big\{{\rm sign}(T\!+\!x\!+\!2\pi n)+{\rm sign}(T\!-\!x\!-\!2\pi n)\big\}\,$,
${\rm C}_{2}(T,x):=\sum_{n=1}^{\infty}\cos nx\,\frac{\cos nT}{n^{2}}$
$\quad=\frac{1}{4}\Big\{{\rm Li}_{2}\big(e^{i(T+x)}\big)+{\rm Li}_{2}\big(e^{-i(T+x)}\big)+{\rm Li}_{2}\big(e^{i(T-x)}\big)+{\rm Li}_{2}\big(e^{-i(T-x)}\big)\Big\}\,.$ (3.7)
Here, ${\rm Li}_{\nu}(z)=\sum_{n\geq 1}z^{n}n^{-\nu}$, $|z|<1$, is the polylogarithm function, analytically continued to $|z|=1$ via an integral representation, and ${\rm C}_{2}$ is a continuous periodic function of bounded variation of $O(1)$. Using (3.1) in (LABEL:PPtwop4) one obtains
$\Delta(t,t^{\prime},\zeta-\zeta^{\prime})=\ln(t/t^{\prime})+2\sum_{n\geq 1}\cos n(\zeta\!-\!\zeta^{\prime})\Delta_{n}^{\rm sub2}(t,t^{\prime})$
$\quad-\frac{t\!-\!t^{\prime}}{\sqrt{tt^{\prime}}}+\frac{\pi}{2\sqrt{tt^{\prime}}}\sum_{n\in\mathbb{Z}}\big\{{\rm sign}(t\!-\!t^{\prime}\!+\!(\zeta\!-\!\zeta^{\prime})\!+\!2\pi n)+{\rm sign}(t\!-\!t^{\prime}\!-\!(\zeta\!-\!\zeta^{\prime})\!-\!2\pi n)\big\}$
$\quad+\frac{t\!-\!t^{\prime}}{4\pi(tt^{\prime})^{3/2}}{\rm C}_{2}(t\!-\!t^{\prime},\zeta\!-\!\zeta^{\prime})\,$, (3.8)
The key aspect is that the remaining oscillatory sum is rapidly converging due to the pointwise $O(1/n^{3})$ decay of $\Delta_{n}^{\rm sub2}$. In fact, truncating the $\sum_{n\geq 1}$ sums to $\sum_{n=1}^{15}$, say, only introduces an error of order $O(10^{-5})$. The so-truncated expression (LABEL:PPtwop7) is then readily programmed and a plot is shown in Fig. 1. Usually there are two discontinuities per periodicity interval, located at $|\zeta-\zeta^{\prime}|=t-t^{\prime}\!\mod 2\pi$. For $t-t^{\prime}=\pi\!\mod 2\pi$, these merge to one. The vanishing outside the lightcone in the zero-centered periodicity interval $\zeta-\zeta^{\prime}\in[-\pi,\pi]$ is clearly visible in the figure.
Figure 1: Commutator function $\Delta(t,1,\zeta-\zeta^{\prime})$, $t\geq 1$, $\zeta-\zeta^{\prime}\in[-\pi,\pi]$.
Returning to (3.1) and $\tau=\ln t$ as time coordinate we obtain its position space version as
$\begin{pmatrix}\phi(\tau,\zeta)\\[4.2679pt] \pi(\tau,\zeta)\end{pmatrix}=\frac{1}{2\pi}\int_{0}^{2\pi}\!d\zeta_{0}\,D(\tau,\zeta;\tau_{0},\zeta_{0})\begin{pmatrix}\phi(\tau_{0},\zeta_{0})\\[4.2679pt] \pi(\tau_{0},\zeta_{0})\end{pmatrix}$
$D(\tau,\zeta;\tau_{0},\zeta_{0})=\begin{pmatrix}-\partial_{\tau_{0}}\Delta(\tau,\tau_{0},\zeta-\zeta_{0})&\Delta(\tau,\tau_{0},\zeta-\zeta_{0})\\[4.2679pt] -\partial_{\tau}\partial_{\tau_{0}}\Delta(\tau,\tau_{0},\zeta-\zeta_{0})&\partial_{\tau}\Delta(\tau,\tau_{0},\zeta-\zeta_{0})\end{pmatrix}\,$, (3.9)
where $\pi(\tau,\zeta):=\partial_{\tau}\phi(\tau,\zeta)$.
The main simplification in the velocity dominated system is that the Hamiltonian is time independent and is determined by the also time independent momentum operator. For the Fourier modes a precise counterpart of (3.1) holds
$\begin{pmatrix}\varphi_{n}(\tau)\\ \wp_{n}(\tau)\end{pmatrix}=\mathfrak{D}_{n}(\tau,\tau_{0})\begin{pmatrix}\varphi_{n}(\tau_{0})\\ \wp_{n}(\tau_{0})\end{pmatrix}$,
$\mathfrak{D}_{n}(\tau,\tau_{0})=\begin{pmatrix}-\partial_{\tau_{0}}\mathfrak{d}_{n}(\tau,\tau_{0})&\mathfrak{d}_{n}(\tau,\tau_{0})\\ -\partial_{\tau}\partial_{\tau_{0}}\mathfrak{d}_{n}(\tau,\tau_{0})&\partial_{\tau}\mathfrak{d}_{n}(\tau,\tau_{0})\end{pmatrix}=\begin{pmatrix}1&\tau-\tau_{0}\\ 0&1\end{pmatrix}$, (3.10)
where $\wp_{n}(\tau):=\partial_{\tau}\varphi_{n}(\tau)$, and $\mathfrak{d}_{n}$ are the ($n$-independent) Fourier modes of the velocity dominated commutator function. [^4] Note that $\mathfrak{D}_{n}(\tau,\tau_{0})=D_{0}(\tau,\tau_{0})$, $n\in\mathbb{Z}$, with $D_{0}$ the zero mode of the time evolution matrix (3.1) of the full system. In parallel to (LABEL:opevol), the position space version of (LABEL:vdopevol) is
$\begin{pmatrix}\varphi(\tau,\zeta)\\[4.2679pt] \wp(\tau,\zeta)\end{pmatrix}=\frac{1}{2\pi}\int_{0}^{2\pi}\!d\zeta_{0}\,\mathfrak{D}(\tau,\zeta;\tau_{0},\zeta_{0})\begin{pmatrix}\varphi(\tau_{0},\zeta_{0})\\[4.2679pt] \wp(\tau_{0},\zeta_{0})\end{pmatrix}$
$\mathfrak{D}(\tau,\zeta;\tau_{0},\zeta_{0})=\begin{pmatrix}1&\tau-\tau_{0}\\ 0&1\end{pmatrix}\delta_{2\pi}(\zeta-\zeta_{0})\,$, (3.11)
where $\wp(\tau,\zeta):=\partial_{\tau}\varphi(\tau,\zeta)$.
3.2 Initial value parameterization and time consistent states
We return to the symmetric matrix two-point function ${\cal W}^{s}$ in (2.4). Using (LABEL:opevol) in (2.4) one finds
${\cal W}^{s}(\tau,\tau^{\prime},\zeta-\zeta^{\prime})=\int_{0}^{2\pi}\!\frac{d\zeta_{0}}{2\pi}\int_{0}^{2\pi}\!\frac{d\zeta_{0}^{\prime}}{2\pi}$ (3.12)
$\quad D(\tau,\zeta;\tau_{0},\zeta_{0})\,{\cal W}^{s}(\tau_{0},\tau_{0},\zeta_{0}\!-\!\zeta^{\prime}_{0})\,D(\tau^{\prime},\zeta^{\prime};\tau_{0},\zeta^{\prime}_{0})^{T}\,$,
where $D^{T}$ is the pointwise matrix transpose of the matrix $D$. The matrix two-point function is therefore parameterized by its initial value kernel ${\cal W}^{s}(\tau_{0},\tau_{0},\zeta_{0}\!-\!\zeta^{\prime}_{0})$ depending only on the single reference time $\tau_{0}$. Below, we shall study the dependence on this reference time. For now, we only note its Fourier representation in terms of a positive definite matrix $Z_{n}(\tau_{0})$,
${\cal W}^{s}(\tau_{0},\tau_{0},\zeta\!-\!\zeta^{\prime})=\sum_{n\in\mathbb{Z}}e^{in(\zeta-\zeta^{\prime})}Z_{n}(\tau_{0})\,$,
$Z_{n}(\tau_{0})\!\!:=\!\!\begin{pmatrix}2|z_{n}|^{2}&z_{n}w_{n}^{*}+z_{n}^{*}w_{n}\\[4.2679pt] z_{n}w_{n}^{*}+z_{n}^{*}w_{n}&2|w_{n}|^{2}\end{pmatrix}(\tau_{0})>0\,.$ (3.13)
where $z_{n}(\tau_{0})$, $w_{n}(\tau_{0})$ are the initial data from (2.25). The positive definiteness of $Z_{n}(\tau_{0})$ is linked to the fact that the underlying Fock vacuum (2.27) is a state in the algebraic quantum field theory sense. It could be rendered manifest by factorizing $Z_{n}(\tau_{0})$ into a matrix and its adjoint. For the basic two-point function $W^{s}$ an initial value parameterization of this form is due to Lüders and Roberts [24]. The matrix version (LABEL:twopmat2) is convenient in that the initial values and the time evolved quantity have the same structure.
Mostly to set the notation we run through the same steps in the velocity dominated system. Defining in parallel to (2.4)
$\mathfrak{W}^{s}(\tau,\tau^{\prime},\zeta-\zeta^{\prime}):=\begin{pmatrix}1&\partial_{\tau^{\prime}}\\ \partial_{\tau}&\partial_{\tau}\partial_{\tau^{\prime}}\end{pmatrix}\mathfrak{w}^{s}(\tau,\tau^{\prime},\zeta-\zeta^{\prime})\,$, (3.14)
the counterpart of (LABEL:twopmat2) is spatially local
$\mathfrak{W}^{s}(\tau,\tau^{\prime},\zeta-\zeta^{\prime})=\begin{pmatrix}1&\tau\!-\!\tau_{0}\\ 0&1\end{pmatrix}\mathfrak{W}^{s}(\tau_{0},\tau_{0},\zeta\!-\!\zeta^{\prime})\,\begin{pmatrix}1&0\\ \tau^{\prime}\!-\!\tau_{0}&1\end{pmatrix}\,$. (3.15)
Again, the initial value kernel $\mathfrak{W}^{s}(\tau_{0},\tau_{0},\zeta\!-\!\zeta^{\prime})$ is parameterized by a positive definite matrix in a Fourier space $\tilde{Z}_{n}(\tau_{0})$.
$\mathfrak{W}^{s}(\tau_{0},\tau_{0},\zeta\!-\!\zeta^{\prime})=\sum_{n\in\mathbb{Z}}e^{in(\zeta-\zeta^{\prime})}\tilde{Z}_{n}(\tau_{0})\,$,
$\tilde{Z}_{n}(\tau_{0})\!\!:=\!\!\begin{pmatrix}2|\tilde{z}_{n}|^{2}&\tilde{z}_{n}w_{n}^{*}+\tilde{z}_{n}^{*}\tilde{w}_{n}\\[4.2679pt] \tilde{z}_{n}\tilde{w}_{n}^{*}+\tilde{z}_{n}^{*}\tilde{w}_{n}&2|\tilde{w}_{n}|^{2}\end{pmatrix}(\tau_{0})>0\,$, (3.16)
where $\tilde{z}_{n}(\tau_{0})$, $\tilde{w}_{n}(\tau_{0})$ are the initial data from (2.29).
Definition. A state in a linear quantum field theoretical system on a spatially homogeneous background is called time consistent if in the realization of the (matrix-) two-point function in terms of the commutator function and an initial data kernel at time $\tau_{0}$ the two-point function itself is independent of the choice of $\tau_{0}$.
The definition draws on the fact that a homogeneous pure state in a linear quantum field theoretical system can be characterized by the spatial Fourier transform of an inital data matrix [24]. The requirement that the two-point function associated with such a state is independent of the time $\tau_{0}$ where the initial data are specified is a very natural requirement, as the standard definition of a two-point function in terms of the expectation value of a product of smeared field operators does not refer to $\tau_{0}$. To the best of our knowledge it has never been formulated as an explicit requirement, since in earlier uses of the initial value formulation $\tau_{0}$ was always kept fixed [24, 31, 21]. In the present context often the initial data of intermediary or derived quantities ought to be backpropagated to the Big Bang as well, so $\tau_{0}$ cannot be kept fixed. In the following we characterize the initial data that give rise to a reconstructed two-point function independent of $\tau_{0}$. For definiteness we do so directly for the Gowdy scalar at hand, i.e. (LABEL:twopmat3) versus (LABEL:twopmat2), (2.4), but the ensued Result 3.1 would generalize to other linear quantum field theories on spatially homogeneous backgrounds. In Section 4.2 we shall see that these initial data are typically compatible with the Hadamard property.
We begin by spelling out the consistency condition imposed on $Z_{n}(\tau_{0})$ by time consistency. In view of (LABEL:twopmat2) this means that two initial value kernels ${\cal W}^{s}(\tau_{0},\tau_{0},\,\cdot\,)$ and ${\cal W}^{s}(\tau_{1},\tau_{1},\,\cdot\,)$ must be related by the $\tau=\tau^{\prime}=\tau_{1}$ specialization of (LABEL:twopmat2). In the Fourier representation (LABEL:twopmat3) this amounts to
$\begin{pmatrix}z_{n}\\ w_{n}\end{pmatrix}\!(\tau_{0})=D_{n}(\tau_{0},\tau_{1})\begin{pmatrix}z_{n}\\ w_{n}\end{pmatrix}\!(\tau_{1})\,,\quad Z_{n}(\tau_{0})=D_{n}(\tau_{0},\tau_{1})Z_{n}(\tau_{1})D_{n}(\tau_{0},\tau_{1})^{T}.$ (3.17)
The same applies to the velocity dominated system where two initial value kernels $\mathfrak{W}^{s}(\tau_{0},\tau_{0},\,\cdot\,)$ and $\mathfrak{W}^{s}(\tau_{1},\tau_{1},\,\cdot\,)$ must be related by the $\tau=\tau^{\prime}=\tau_{1}$ specialization of (3.15). In the Fourier representation (LABEL:twopmat8)
$\begin{pmatrix}\tilde{z}_{n}\\ \tilde{w}_{n}\end{pmatrix}\!(\tau_{0})=\begin{pmatrix}1&\tau_{0}\!-\!\tau_{1}\\ 0&1\end{pmatrix}\!\begin{pmatrix}\tilde{z}_{n}\\ \tilde{w}_{n}\end{pmatrix}\!(\tau_{1})\,,\quad\tilde{Z}_{n}(\tau_{0})=\begin{pmatrix}1\!&\!\tau_{0}\!-\!\tau_{1}\\ 0\!&\!1\end{pmatrix}\tilde{Z}_{n}(\tau_{1})\begin{pmatrix}1\!&\!0\\ \!\tau_{0}\!-\!\tau_{1}\!&\!1\end{pmatrix}$. (3.18)
We shall refer to initial data $\big(z_{n}(\tau_{0}),w_{n}(\tau_{0})\big)$ satisfying (3.17) as time consistent initial data; similarly, for the velocity dominated system and (3.18). We now claim that such time consistent initial data can be constructed by transitioning to the corresponding Bogoliubov parameters.
Result 3.1.
The transformations (2.26), (2.31) trivialize the consistency conditions (3.17), (3.18) for the initial data under a change of reference time: the latter hold iff the Bogoliubov parameters $\lambda_{n}(\tau_{0}),\mu_{n}(\tau_{0})$ and $\tilde{\lambda}_{n}(\tau_{0}),\tilde{\mu}_{n}(\tau_{0})$, $n\in\mathbb{Z}$, are independent of $\tau_{0}$. Further, both sets of Bogoliubov parameters are consistently related by the time independent relation (3.21) below. Both sets of time consistent initial data are in turn related by $I^{\rm grad}(|n|e^{\tau_{0}})$ from (3.2). Overall, the four maps give rise to the commutative diagram in Figure 2.
[…]
Figure 2: Commutative diagram. All four maps are invertible.
The derivation of this result is based on the fact that the solutions of the full Gowdy system are related to those of the velocity-dominated system through the following gradient map
$\begin{pmatrix}T_{n}(\tau)\\[4.2679pt] \partial_{\tau}T_{n}(\tau)\end{pmatrix}$ $\!\!=\!\!$ $I^{\rm grad}(|n|e^{\tau})\begin{pmatrix}\mathfrak{t}_{n}(\tau)\\[4.2679pt] \partial_{\tau}\mathfrak{t}_{n}(\tau)\end{pmatrix}\,,\quad n\in\mathbb{Z}\,$,
$I^{\rm grad}(x)$ $:=$ $\begin{pmatrix}J_{0}(x)&U_{0}(x)\\[4.2679pt] -xJ_{1}(x)&-xU_{1}(x)\end{pmatrix}=:\mbox{1\hskip-3.44444pt1}+\sum_{k\geq 1}I_{k}\,x^{2k}\,$, (3.19)
with $U_{k}(x):=(\pi/2)Y_{k}(x)-\ln(|x|e^{\gamma_{E}}/2)J_{k}(x)$, $k=0,1$. The matrix is sympletic, $\det I^{\rm grad}(x)=1$, which preserves the Wronskian normalizations. We refer to it as the gradient map, because all matrix elements have absolutely convergent Taylor expansions in powers of $x^{2}$, as indicated. For $x^{2}=n^{2}e^{2\tau}$ the $n^{2}$ powers in turn correspond to spatial gradients in position space.
To see how (3.2) comes about, one starts with
$T_{n}(\tau)$ $\!\!=\!\!$ $J_{0}(|n|e^{\tau})\mathfrak{t}_{n}(\tau)+U_{0}(|n|e^{\tau})\partial_{\tau}\mathfrak{t}_{n}(\tau)$ (3.20)
$\!\!=\!\!$ $J_{0}(|n|e^{\tau})\big[\mathfrak{t}_{n}(\tau)-\partial_{\tau}\mathfrak{t}_{n}(\tau)\ln(e^{\tau_{0}}e^{\gamma_{E}}|n|/2)\big]+\frac{\pi}{2}Y_{0}(|n|e^{\tau})\partial_{\tau}\mathfrak{t}_{n}(\tau)\,.$
The coefficients of $J_{0}(|n|e^{\tau})$ and $Y_{0}(|n|e^{\tau})$ are constants, and by rewriting the expression in terms of $H_{0}^{(1)}(|n|e^{\tau})$ and $H_{0}^{(2)}(|n|e^{\tau})$, one can match the result to (2.3). Repeating these steps for $\partial_{\tau}T_{n}(\tau)$, one obtains another set of matching conditions. Both combine to the following time independent identification of Bogoliubov parameters.
$\begin{pmatrix}\lambda_{n}(\tau_{0})\\ \mu_{n}(\tau_{0})\end{pmatrix}$ $\!\!=\!\!$ $\begin{pmatrix}1+\frac{i}{\pi}\ln\gamma_{n}(0)&-\frac{i}{\pi}\ln\gamma_{n}(0)\\[4.2679pt] \frac{i}{\pi}\ln\gamma_{n}(0)&1-\frac{i}{\pi}\ln\gamma_{n}(0)\end{pmatrix}\begin{pmatrix}\tilde{\lambda}_{n}(\tau_{0})\\[4.2679pt] \tilde{\mu}_{n}(\tau_{0})\end{pmatrix},\quad n\neq 0\,$, (3.21)
where $\gamma_{n}(0)=e^{\gamma_{E}}|n|/2$. Conversely, subject to the identification (3.21) the relation (3.2) holds.
Specializing (3.2) to $\tau=\tau_{0}$ one has
$\begin{pmatrix}z_{n}(\tau_{0})\\[4.2679pt] w_{n}(\tau_{0})\end{pmatrix}=I^{\rm grad}(|n|e^{\tau_{0}})\begin{pmatrix}\tilde{z}_{n}(\tau_{0})\\[4.2679pt] \tilde{w}_{n}(\tau_{0})\end{pmatrix}\,,\quad n\in\mathbb{Z}\,.$ (3.22)
Since $I^{\rm grad}(0)$ is the unit matrix, the $n=0$ instance of (LABEL:init1) simply identifies $z_{0}(\tau_{0})=\tilde{z}_{0}(\tau_{0}),\,w_{0}(\tau_{0})=\tilde{w}_{0}(\tau_{0})$. The interrelation (LABEL:init1) must be compatible with the respective consistency conditions (3.17), (3.18) on both sides. This amounts to the relation
$D_{n}(\tau_{1},\tau_{0})I^{\rm grad}(|n|e^{\tau_{0}})\begin{pmatrix}1&\tau_{0}\!-\!\tau_{1}\\ 0&1\end{pmatrix}=I^{\rm grad}(|n|e^{\tau_{1}})\,,\quad n\in\mathbb{Z}\,$, (3.23)
which is indeed an identity. Inserting (2.26), (2.31) into (LABEL:init1) one finds
$\begin{pmatrix}\lambda_{n}(\tau_{0})\\ \mu_{n}(\tau_{0})\end{pmatrix}$ $\!\!=\!\!$ $H(|n|e^{\tau_{0}})^{-1}I^{\rm grad}(|n|e^{\tau_{0}})\tilde{H}(\tau_{0})\begin{pmatrix}\tilde{\lambda}_{n}(\tau_{0})\\[4.2679pt] \tilde{\mu}_{n}(\tau_{0})\end{pmatrix}\,$. (3.24)
Consistency requires that the matrices in (3.24) and (3.21) are the same. This can be checked to be the case and yields the commutativity of the above diagram.
So far, we held the reference time $\tau_{0}$ fixed. Next we consider the transformation of the Bogoliubov parameters $\lambda_{n}(\tau_{0}),\mu_{n}(\tau_{0})$ and $\tilde{\lambda}_{n}(\tau_{0}),\tilde{\mu}_{n}(\tau_{0})$, $n\in\mathbb{Z}$, under a change of initial time. Note that the transformation law is uniquely determined by (3.17), (2.26) and (3.18), (2.31), respectively. In the first case one finds for $n\neq 0$
$\begin{pmatrix}\lambda_{n}(\tau_{0})\\ \mu_{n}(\tau_{0})\end{pmatrix}=H(|n|e^{\tau_{0}})^{-1}D_{n}(\tau_{0},\tau_{1})H(|n|e^{\tau_{1}})\begin{pmatrix}\lambda_{n}(\tau_{1}\\ \mu_{n}(\tau_{1})\end{pmatrix}=\begin{pmatrix}\lambda_{n}(\tau_{1})\\ \mu_{n}(\tau_{1})\end{pmatrix}\,$, (3.25)
as the matrix reduces to the unit matrix. In the velocity dominated system one has
$\begin{pmatrix}\tilde{\lambda}_{n}(\tau_{0})\\ \tilde{\mu}_{n}(\tau_{0})\end{pmatrix}=\tilde{H}(\tau_{0})^{-1}\begin{pmatrix}1&\tau_{0}\!-\!\tau_{1}\\ 0&1\end{pmatrix}\tilde{H}(\tau_{1})\begin{pmatrix}\tilde{\lambda}_{n}(\tau_{1})\\ \tilde{\mu}_{n}(\tau_{1})\end{pmatrix}=\begin{pmatrix}\tilde{\lambda}_{n}(\tau_{1})\\ \tilde{\mu}_{n}(\tau_{1})\end{pmatrix}\,$, (3.26)
where the matrix again reduces to the unit matrix. This shows the main assertion in the above result: the Bogoliubov parameters do not depend on the choice of reference time. The relation (3.26) applies to all $n\in\mathbb{Z}$, while for (3.25) the zero modes need to be augmented. We do so by identifying
$\lambda_{0}(\tau_{0})=\tilde{\lambda}_{0}(\tau_{0})\,,\quad\mu_{0}(\tau_{0})=\tilde{\mu}_{0}(\tau_{0})\,$, (3.27)
consistent with our initial identification $T_{0}(\tau)=\mathfrak{t}_{0}(\tau)$.
Conversely, re-expressing the first equality in (3.25) in terms of initial data using (2.26), one obtains (3.17). That is, time independent Bogoliubov parameters give via (2.26) automatically rise to time consistent initial data. Similarly, re-expressing the first equality in (3.26) in terms of initial data using (2.31), gives rise to (3.18). This completes the derivation of the above result.
3.3 Spatial gradient expansion of two-point functions
As seen before, in Fourier space the basic mode functions $(T_{n},\partial_{\tau}T_{n})$ and their velocity dominated counterparts $(\mathfrak{t}_{n},\partial_{\tau}\mathfrak{t}_{n})$ are related by the gradient map (3.2). Since $I^{\rm grad}(x)$ is real this extends to their complex conjugates and one has
$\begin{pmatrix}T_{n}(\tau)&T_{n}(\tau)^{*}\\[5.69054pt] \dot{T}_{n}(\tau)&\dot{T}_{n}(\tau)^{*}\end{pmatrix}=I^{\rm grad}(|n|e^{\tau})\begin{pmatrix}\mathfrak{t}_{n}(\tau)&\mathfrak{t}_{n}(\tau)^{*}\\[5.69054pt] \dot{\mathfrak{t}}_{n}(\tau)&\dot{\mathfrak{t}}_{n}(\tau)^{*}\end{pmatrix}\,$, (3.28)
where we write $\dot{f}(\tau)=(\partial_{\tau}f)(\tau)$. The Fourier modes of the basic two-point functions can be written as
$W_{n}^{s}(\tau,\tau^{\prime})$ $:=$ $T_{n}(\tau)T_{n}(\tau^{\prime})^{*}+T_{n}(\tau^{\prime})T_{n}(\tau)^{*}\,$,
$\Delta_{n}(\tau,\tau^{\prime})$ $:=$ $i\big(T_{n}(\tau)T_{n}(\tau^{\prime})^{*}-T_{n}(\tau^{\prime})T_{n}(\tau)^{*}\big)\,.$ (3.29)
We seek to express their matrix versions in terms of the left hand side in (3.28). One has
${\cal W}_{n}^{s}(\tau,\tau^{\prime})=\begin{pmatrix}1&\partial_{\tau^{\prime}}\\[5.69054pt] \partial_{\tau}&\partial_{\tau}\partial_{\tau^{\prime}}\end{pmatrix}W_{n}^{s}(\tau,\tau^{\prime})=\begin{pmatrix}T_{n}(\tau)&T_{n}(\tau)^{*}\\[5.69054pt] \dot{T}_{n}(\tau)&\dot{T}_{n}(\tau)^{*}\end{pmatrix}\begin{pmatrix}T_{n}(\tau^{\prime})&T_{n}(\tau^{\prime})^{*}\\[5.69054pt] \dot{T}_{n}(\tau^{\prime})&\dot{T}_{n}(\tau^{\prime})^{*}\end{pmatrix}^{\dagger}\!$. (3.30)
The matrix commutator function enters naturally in the form (3.1) via the solution of the initial value problem. This gives
$D_{n}(\tau,\tau^{\prime})\begin{pmatrix}0&-1\\[5.69054pt] 1&0\end{pmatrix}=\begin{pmatrix}1&\partial_{\tau^{\prime}}\\[5.69054pt] \partial_{\tau}&\partial_{\tau}\partial_{\tau^{\prime}}\end{pmatrix}\Delta_{n}(\tau,\tau^{\prime})$
$\quad=i\begin{pmatrix}T_{n}(\tau)&T_{n}(\tau)^{*}\\[5.69054pt] \dot{T}_{n}(\tau)&\dot{T}_{n}(\tau)^{*}\end{pmatrix}\begin{pmatrix}1&0\\[5.69054pt] 0&-1\end{pmatrix}\begin{pmatrix}T_{n}(\tau^{\prime})&T_{n}(\tau^{\prime})^{*}\\[5.69054pt] \dot{T}_{n}(\tau^{\prime})&\dot{T}_{n}(\tau^{\prime})^{*}\end{pmatrix}^{\dagger}\,$. (3.31)
The same relations apply to the velocity dominated matrix two-point functions $\mathfrak{W}_{n}^{s}(\tau,\tau^{\prime})$ and $\mathfrak{D}_{n}(\tau,\tau^{\prime})$, respectively, with the $T_{n}$’s replaced by $\mathfrak{t}_{n}$’s. Inserting (3.28) into (3.30) one obtains
${\cal W}_{n}^{s}(\tau,\tau^{\prime})=I^{\rm grad}(|n|e^{\tau})\mathfrak{W}_{n}^{s}(\tau,\tau^{\prime})I^{\rm grad}(|n|e^{\tau^{\prime}})^{T}\,,\quad n\neq 0.$ (3.32)
For the zero mode, we identified $T_{0}(\tau)$ with $\mathfrak{t}_{0}(\tau)$, so that ${\cal W}_{0}^{s}(\tau,\tau^{\prime})=\mathfrak{W}_{0}^{s}(\tau,\tau^{\prime})$ holds by definition. For the commutator function one needs in addition
$\begin{pmatrix}0&1\\[5.69054pt] -1&0\end{pmatrix}^{\!\!-1}I^{\rm grad}(x)^{T}\begin{pmatrix}0&1\\[5.69054pt] -1&0\end{pmatrix}=I^{\rm grad}(x)^{-1}\,.$ (3.33)
This results in
$D_{n}(\tau,\tau^{\prime})=I^{\rm grad}(|n|e^{\tau})\mathfrak{D}_{n}(\tau,\tau^{\prime})I^{\rm grad}(|n|e^{\tau^{\prime}})^{-1}\,,\quad n\neq 0.$ (3.34)
For $n=0$ one has again trivially $D_{0}(\tau,\tau^{\prime})=\mathfrak{D}_{0}(\tau,\tau^{\prime})$. Note that the inverse on the far right of (3.34) ensures compatibility with the composition law mentioned after (3.1). An alternative, slightly quicker, route to (3.34) is by rearranging (3.23) using $\mathfrak{D}_{n}(\tau^{\prime},\tau)=\mathfrak{D}_{n}(\tau,\tau^{\prime})^{-1}$.
Since $I^{\rm grad}(x)$ admits a convergent series expansion in powers of $x^{2}$ the relations (3.32), (3.34) will give rise to convergent expansions of the matrix two-point functions in Fourier space around their velocity dominated counterparts. Preparing
$I^{\rm grad}(x)$ $\!\!=\!\!$ $\mbox{1\hskip-3.44444pt1}+x^{2}I_{1}+O(x^{4})\,,\quad I_{1}=\begin{pmatrix}-1/4&1/4\\ -1/2&1/4\end{pmatrix}\,$, (3.35)
one finds to first order
${\cal W}^{s}_{n}(\tau,\tau^{\prime})=\mathfrak{W}_{n}^{s}(\tau,\tau^{\prime})+n^{2}[e^{2\tau}I_{1}\mathfrak{W}_{n}^{s}(\tau,\tau^{\prime})+e^{2\tau^{\prime}}\mathfrak{W}_{n}^{s}(\tau,\tau^{\prime})I_{1}^{T}]+O(n^{4})\,$,
$D_{n}(\tau,\tau^{\prime})=\mathfrak{D}_{n}(\tau,\tau^{\prime})+n^{2}[e^{2\tau}I_{1}\mathfrak{D}_{n}(\tau,\tau^{\prime})-e^{2\tau^{\prime}}\mathfrak{D}_{n}(\tau,\tau^{\prime})I_{1}]+O(n^{4})\,.$ (3.36)
This illustrates a systematic way to reconstruct the full two-point functions from the velocity dominated ones. The powers of $n^{2}$ encountered amount to a spatial gradient expansion, so the reconstruction does not rely on the time reversal in a unitary time evolution operator (supposed to undo the $\tau,\tau^{\prime}\rightarrow$-$\infty$ asymptotics, as in the classical AVD results.)
In position space, the velocity dominated two-point functions will in general contain distributional terms in $\zeta-\zeta^{\prime}$ and will not be conventionally differentiable to all orders. However once spatially averaged with real, smooth test functions $f,g\in C^{\infty}(S^{1})$, the differentiations can be defined distributionally. That is, for any element $\mathfrak{F}$ in the dual [^5] That is, $\mathfrak{F}(\zeta)=\sum_{m}\mathfrak{F}_{m}e^{im\zeta}$, where there exists $c^{\prime}>0,M\in\mathbb{N}$ such that $|\mathfrak{F}_{m}|\leq c^{\prime}(1+|m|)^{M}$, for all $m\in\mathbb{Z}$. This is is dual to the characterization of $f\in C^{\infty}(S^{1})$, i.e. $f(\zeta)=\sum_{n}f_{n}e^{in\zeta}$, with $|f_{n}|\leq c(1+|n|)^{-N}$, for all $N\in\mathbb{N}$. of $C^{\infty}(S^{1})$ and $l\in\mathbb{N}$ we set
$\int\!\frac{d\zeta}{2\pi}f(\zeta)\int\!\frac{d\zeta^{\prime}}{2\pi}g(\zeta^{\prime})(\partial_{\zeta}\partial_{\zeta^{\prime}})^{l}\mathfrak{F}(\zeta\!-\!\zeta^{\prime})$ $:=$ $\int\!\frac{d\zeta}{2\pi}\partial_{\zeta}^{l}f(\zeta)\int\!\frac{d\zeta^{\prime}}{2\pi}\partial_{\zeta^{\prime}}^{l}g(\zeta^{\prime})\mathfrak{F}(\zeta\!-\!\zeta^{\prime})$
$=$ $\sum_{n\in\mathbb{Z}}n^{2l}f_{-n}g_{n}\,\mathfrak{F}_{n}\,.$ (3.37)
Convergence of the sum is ensured by the rapid decay of $f_{n},g_{-n}$ in $n$. This can be used to show the
Result 3.2.
The matrix two-point function ${\cal W}^{s}(\tau,\tau^{\prime},\zeta-\zeta^{\prime})$ admits upon averaging with $f,g\in C^{\infty}(S^{1})$ test functions a series expansion of the form
$\int\!\!\frac{d\zeta}{2\pi}\frac{d\zeta^{\prime}}{2\pi}f(\zeta)g(\zeta^{\prime})\,{\cal W}^{s}(\tau,\tau^{\prime},\zeta\!-\!\zeta^{\prime})=\int\!\!\frac{d\zeta}{2\pi}\frac{d\zeta^{\prime}}{2\pi}f(\zeta)g(\zeta^{\prime})\mathfrak{W}^{s}(\tau,\tau^{\prime},\zeta\!-\!\zeta^{\prime})$
$+\sum_{l\geq 1}\sum_{k=0}^{l}e^{2k\tau}e^{2(l-k)\tau^{\prime}}\int\!\!\frac{d\zeta}{2\pi}\frac{d\zeta^{\prime}}{2\pi}f(\zeta)g(\zeta^{\prime})(\partial_{\zeta}\partial_{\zeta^{\prime}})^{l}I_{k}\,\mathfrak{W}^{s}(\tau,\tau^{\prime},\zeta\!-\!\zeta^{\prime})I^{T}_{l-k}\,$, (3.38)
Here, $I_{k}$ are the numerical $2\times 2$ matrices in the expansion of $I^{\rm grad}(x)$ in powers of $x^{2}$. All terms in the expansion are spatial gradients of the two-point function $\mathfrak{W}^{s}(\tau,\tau^{\prime},\zeta-\zeta^{\prime})$ of the velocity dominated system, which is linear in $\tau,\tau^{\prime}$. The series (LABEL:grad20) is uniformly convergent for all $\tau,\tau^{\prime}<-\delta$, $\delta>0$ and $|\tau-\tau^{\prime}|$ bounded.
Only the convergence part needs to be shown. We start from (3.32) and insert the Taylor series $I^{\rm grad}(x)=\sum_{k=0}^{K}I_{k}x^{2k}+C_{K}x^{2K}$ with Peano remainder, i.e. $C_{K}$ goes to zero as $K\rightarrow\infty$ in one (and hence any) $2\times 2$ matrix norm. A brief computation gives
${\cal W}^{s}_{n}(\tau,\tau^{\prime})-\mathfrak{W}_{n}^{s}(\tau,\tau^{\prime})-\sum_{l=1}^{2K}n^{2l}\Big(\sum_{k=0}^{l}I_{k}\mathfrak{W}^{s}_{n}(\tau,\tau^{\prime})I_{l-k}^{T}e^{2k\tau}e^{2(l-k)\tau^{\prime}}\Big)$
$\quad=\sum_{k=0}^{K}n^{2(k+K)}\Big(e^{2k\tau+2K\tau^{\prime}}I_{k}\mathfrak{W}^{s}_{n}(\tau,\tau^{\prime})C_{K}^{T}+e^{2k\tau^{\prime}+2K\tau}C_{K}\mathfrak{W}^{s}_{n}(\tau,\tau^{\prime})I_{k}^{T}\Big)$
$\quad+n^{4K}e^{2K(\tau+\tau^{\prime})}C_{K}\mathfrak{W}^{s}_{n}(\tau,\tau^{\prime})C_{K}^{T}\,.$ (3.39)
In view of (3.3) we multiply this by $f_{n}g_{-n}+f_{-n}g_{n}$ and sum over $n\in\mathbb{N}$. The resulting right hand side is denoted by RHS and we seek to bound it in some $2\times 2$ matrix norm. For definiteness we use the $\|\cdot\|_{\infty}$ norm from (3.40) below. We only sketch the rest of the argument, deferring some of the details to Section 3.4: For large enough $K$ we may assume $\|C_{K}\|_{\infty}\leq 1$. By inspection of (3.2) one sees that $\sum_{k=0}^{K}\|I_{k}\|_{\infty}\leq 2$, for all $K$. This readily gives $\|{\rm RHS}\|_{\infty}\leq 2(2e^{2K\tau}+2e^{2K\tau^{\prime}}+e^{2K(\tau+\tau^{\prime}})\sum_{n\geq 1}|f_{n}||g_{n}|n^{4K}\|\mathfrak{W}^{s}_{n}(\tau,\tau^{\prime})\|_{\infty}$. A bound for $\|\mathfrak{W}^{s}_{n}(\tau,\tau^{\prime})\|_{\infty}$ will be provided in (3.45), (3.47). Since $|\tilde{\lambda}_{n}|,|\tilde{\mu}_{n}|$ grow like $\log n$ for the relevant cases, see Section 4.2, we may safely bound the sum by $\sum_{n\geq 1}|f_{n}||g_{n}|n^{4K+2\nu}$, for a small power $0<\nu<1$. From footnote 5 we know $|f_{n}|,|g_{n}|\leq c(1+|n|)^{-N}$, for all $N\geq 1$. Taking $N=2\nu+5K$, the previous sum can be bounded by $c\sum_{n\geq 1}n^{-K}=c\,\zeta(K)$. For large $K$ the $\zeta$ function approaches $1$, and for real $K\geq 2$ it is bounded by $2$. It remains to take into account the $\tau,\tau^{\prime}$ powers from (3.45). They lead to terms of the form $e^{2K\tau},e^{2K\tau^{\prime}}$, $\tau e^{2K\tau},\tau^{\prime}e^{2K\tau^{\prime}}$, and $\tau^{\prime}e^{2K\tau},\tau e^{2K\tau^{\prime}}$, $K\geq 2$. All of them are readily seen to be uniformly bounded for $|\tau\!-\!\tau^{\prime}|<d$ and $\tau,\tau^{\prime}<-\delta$, for some (small) $\delta>0$ and (large) $d>0$. This establishes Result 3.2.
3.4 Asymptotic velocity domination for two-point functions
Here we derive an explicit bound on the second line of (LABEL:grad20) that pins down the leading rate of decay. This could be done for each of the $2\times 2$ matrix components separately, but it is technically convenient to use a matrix norm. A matrix norm on the space of $n\times n$ matrices obeys: $\|A\|\geq 0$; $\|A\|=0$ iff $A=0$; $\|\alpha A\|=|\alpha|\|A\|$, $\alpha\in\mathbb{C}$; $\|A+B\|\leq\|A\|+\|B\|$; $\|AB\|\leq\|A\|\|B\|$. For finite dimensional matrices any two of such norms are equivalent and we use the sup-norm for convenience
$\|A\|_{\infty}=\max_{1\leq j\leq n}\sum_{i=1}^{n}|a_{ij}|\,,\quad(\mbox{row sums})$ (3.40)
where $a_{ij}$, $1\leq i,j\leq n$, are the matrix elements of $A$. In this norm we consider
$\bigg\|\int\!\!\frac{d\zeta}{2\pi}\frac{d\zeta^{\prime}}{2\pi}f(\zeta)g(\zeta^{\prime})\Big\{{\cal W}^{s}(\tau,\tau^{\prime},\zeta\!-\!\zeta^{\prime})-\mathfrak{W}^{s}(\tau,\tau^{\prime},\zeta\!-\!\zeta^{\prime})\,\Big\}\bigg\|_{\infty}$
$\leq\sum_{n\neq 0}\Big\|f_{n}g_{n}{\scriptsize\Big\{{\cal W}_{n}^{s}(\tau,\tau^{\prime})-\mathfrak{W}_{n}^{s}(\tau,\tau^{\prime})\Big\}\Big\|_{\infty}}\,.$ (3.41)
The $n=0$ term drops out on account of identification of zero modes. Next, we insert (3.32) and split off the unit matrix part from $I^{\rm{grad}}$, defining $\hat{I}(x)=I^{\rm{grad}}(x)-\mbox{1\hskip-3.44444pt1}$. Termwise this gives
$\Big\|f_{n}g_{n}\Big\{{\cal W}_{n}^{s}(\tau,\tau^{\prime})-\mathfrak{W}_{n}^{s}(\tau,\tau^{\prime})\Big\}\Big\|_{\infty}$
$=\Big\|f_{n}g_{n}\Big\{\hat{I}(|n|e^{\tau})\mathfrak{W}_{n}^{s}(\tau,\tau^{\prime})+\mathfrak{W}_{n}^{s}(\tau,\tau^{\prime})\hat{I}(|n|e^{\tau^{\prime}})^{T}+\hat{I}(|n|e^{\tau})\mathfrak{W}_{n}^{s}(\tau,\tau^{\prime})\hat{I}(|n|e^{\tau^{\prime}})^{T}\Big\}\Big\|_{\infty}$
$\leq|f_{n}||g_{n}|\Big\{\|\hat{I}(|n|e^{\tau})\|_{\infty}\|\mathfrak{W}_{n}^{s}(\tau,\tau^{\prime})\|_{\infty}+\|\mathfrak{W}_{n}^{s}(\tau,\tau^{\prime})\|_{\infty}\|\hat{I}(|n|e^{\tau^{\prime}})^{T}\|_{\infty}$
$\qquad+\|\hat{I}(|n|e^{\tau})\|_{\infty}\|\mathfrak{W}_{n}^{s}(\tau,\tau^{\prime})\|_{\infty}\|\hat{I}(|n|e^{\tau^{\prime}})^{T}\|_{\infty}\Big\}.$ (3.42)
Clearly only bounds on $\hat{I}$ and $\mathfrak{W}^{s}_{n}$ are needed. To obtain a bound on $\hat{I}(x)$, we recall the definition of $I^{\rm grad}(x)$ from (3.2) and the redefinitions $J_{0}(x)=1-(x^{2}/4)\hat{\jmath}_{0}(x)$, $J_{1}(x)=(x/2)\hat{\jmath}_{1}(x)$, $U_{0}(x)=(x^{2}/4)\hat{u}_{0}(x)$, $U^{\prime}_{0}(x)=(x/2)\hat{u}_{1}(x)$. For the lower right entry of $I^{\rm grad}(x)$ we need $xU_{1}(x)$, where $U_{1}(x):=(\pi/2)Y_{1}(x)-\ln(|x|e^{\gamma_{E}}/2)J_{1}(x)$. Differentiating $U_{0}(x):=(\pi/2)Y_{0}(x)-\ln(|x|e^{\gamma_{E}}/2)J_{0}(x)$ one finds $U^{\prime}_{0}(x)=-U_{1}(x)-J_{0}(x)/x$, so that $xU_{1}(x)=1+(x^{2}/4)(2\hat{u}_{1}(x)-\hat{\jmath}_{0}(x))$. Inserted into (3.2) one gets
$\hat{I}(x)=I^{\rm grad}(x)-\begin{pmatrix}1&0\\ 0&1\end{pmatrix}=\frac{x^{2}}{4}\begin{pmatrix}-\hat{\jmath}_{0}(x)&\hat{u}_{0}(x)\\ -2\hat{\jmath}_{1}(x)&2\hat{u}_{1}(x)-\hat{\jmath}_{0}(x)\end{pmatrix}$. (3.43)
Since the modulus of the ‘hatted’ functions is globally bounded by $1$, this gives straightforwardly
$\|\hat{I}(x)\|_{\infty}\leq\frac{5}{4}x^{2}\,,\quad\|\hat{I}(x)^{T}\|_{\infty}\leq x^{2}\,.$ (3.44)
Next, for a bound on $\mathfrak{W}^{s}_{n}(\tau,\tau^{\prime})$ we combine equations (3.15) and (LABEL:twopmat8) to obtain
$\|\mathfrak{W}^{s}_{n}(\tau,\tau^{\prime})\|_{\infty}\leq(1+|\tau-\tau_{0}|)(1+|\tau^{\prime}-\tau_{0}|)\|\tilde{Z}_{n}(\tau_{0})\|_{\infty}\,.$ (3.45)
For the $\tilde{Z}_{n}$ norm we use
$\|\tilde{Z}_{n}(\tau_{0})\|_{\infty}$ $\!\!=\!\!$ $\max\big\{2|\tilde{z}_{n}|^{2}\!+\!|\tilde{w}_{n}\tilde{z}_{n}^{*}\!+\!\tilde{w}_{n}^{*}\tilde{z}_{n}|,2|\tilde{w}_{n}|^{2}\!+\!|\tilde{w}_{n}\tilde{z}_{n}^{*}\!+\!\tilde{w}_{n}^{*}\tilde{z}_{n}|\big\}(\tau_{0})$ (3.46)
$\leq$ $2\big(|\tilde{z}_{n}(\tau_{0})|\!+\!|\tilde{w}_{n}(\tau_{0})|\big)^{2}\,.$
This is still $\tau_{0}$ dependent, but for time consistent initial data we know that $\tilde{z}_{n}(\tau_{0}),\tilde{w}_{n}(\tau_{0})$ arise from $\tau_{0}$-independent Bogoliubov parameters $\tilde{\lambda}_{n},\tilde{\mu}_{n}$ via (2.31). This gives
$\|\tilde{Z}_{n}(\tau_{0})\|_{\infty}\leq\mathfrak{z}_{n}(\tau_{0})\,,\quad\mathfrak{z}_{n}(\tau_{0})=\frac{2}{\pi}\big(|\tilde{\lambda}_{n}|+|\tilde{\mu}_{n}|\big)^{2}\big(\tau_{0}+1+\frac{\pi}{2}\big)^{2}\,.$ (3.47)
With these bounds in place we return to (LABEL:Ibound0) to estimate
$\Big\|f_{n}g_{n}{\scriptsize\Big\{{\cal W}_{n}^{s}(\tau,\tau^{\prime})-\mathfrak{W}_{n}^{s}(\tau,\tau^{\prime})\Big\}\Big\|_{\infty}}$ (3.48)
$\leq|f_{n}||g_{n}|(1+|\tau-\tau_{0}|)(1+|\tau^{\prime}-\tau_{0}|)\mathfrak{z}_{n}(\tau_{0})\Big\{\frac{5}{4}n^{2}e^{2\tau}+n^{2}e^{2\tau^{\prime}}+\frac{5}{4}n^{4}e^{2(\tau+\tau^{\prime})}\Big\}$
Defining
$m_{2l}(\tau_{0})=\frac{5}{4}\sum_{n\neq 0}|f_{n}||g_{n}|\,\mathfrak{z}_{n}(\tau_{0})\,n^{2l}\,,\quad l\in\mathbb{N}\,$, (3.49)
we can return to the mode sum (LABEL:avdpr1) and arrive at the
Result 3.3.
The difference between the spatially averaged matrix two-point function ${\cal W}^{s}(\tau,\tau^{\prime},\zeta\!-\!\zeta^{\prime})$ in the full Gowdy system and its counterpart $\mathfrak{W}^{s}(\tau,\tau^{\prime},\zeta\!-\!\zeta^{\prime})$ in the velocity dominated system can for $f,g\in C^{\infty}(S^{1})$ be bounded as follows:
$\bigg\|{\scriptsize\int\!\!\frac{d\zeta}{2\pi}\frac{d\zeta^{\prime}}{2\pi}f(\zeta)g(\zeta^{\prime})\Big\{{\cal W}^{s}(\tau,\tau^{\prime},\zeta\!-\!\zeta^{\prime})-\mathfrak{W}^{s}(\tau,\tau^{\prime},\zeta\!-\!\zeta^{\prime})\,\Big\}}\bigg\|_{\infty}$
$\leq(1+|\tau-\tau_{0}|)(1+|\tau^{\prime}-\tau_{0}|)\Big\{e^{2\tau+2\tau^{\prime}}m_{4}(\tau_{0})+(e^{2\tau}+e^{2\tau^{\prime}})m_{2}(\tau_{0})\Big\}\,.$ (3.50)
Here, both $W^{s}$ and $\mathfrak{W}^{s}$ are taken to derive from time consistent initial data, so that the left hand side (though not the bound) is independent of $\tau_{0}$.
Remarks.
(i) This bound entails the desired result: if $\tau,\tau^{\prime},\tau_{0}$ are back-propagated towards the Big Bang via $\tau=\tilde{\tau}+\eta$, $\tau^{\prime}=\tilde{\tau}^{\prime}+\eta$, $\tau_{0}=\tilde{\tau}_{0}+\eta$, $\eta\rightarrow$-$\infty$, the right hand side vanishes like $O(\eta^{2}e^{2\eta})$. In particular $\tau-\tau^{\prime}=\tilde{\tau}-\tilde{\tau}^{\prime}$ stays fixed in the limit, mirroring the condition for uniform convergence in Result 3.2. In the latter case the time consistency of the underlying state does not directly enter, as it is tacit on both sides of (3.32).
(ii) The assumption $f,g\in C^{\infty}(S^{1})$ is clearly not necessary for the proof to go through. Even with the crude Bessel function bounds leading to (3.44) only the moments up to $m_{4}(\tau_{0})$ need to be finite. With $\tilde{\lambda}_{n},\tilde{\mu}_{n}$ to be at most of $\ln|n|$ type growth this requires the sequence $|f_{n}||g_{n}|n^{4}\ln|n|$ to be summable. This means one could replace $C^{\infty}(S^{1})$ by the space of twice differentiable functions $C^{2}(S^{1})$, equipped with the usual norm $\|f\|_{C^{2}(S^{1})}=\sum_{j=0}^{2}\sup_{\zeta\in S^{1}}|\partial_{\zeta}^{j}f|$.
(iii) A similar bound can be obtained for the difference of the commutator functions. Starting from (3.34) and proceeding as above, one finds
$\bigg\|{\scriptsize\int\!\!\frac{d\zeta}{2\pi}\frac{d\zeta^{\prime}}{2\pi}f(\zeta)g(\zeta^{\prime})\Big\{D(\tau,\tau^{\prime},\zeta\!-\!\zeta^{\prime})-\mathfrak{D}(\tau,\tau^{\prime},\zeta\!-\!\zeta^{\prime})\,\Big\}}\bigg\|_{\infty}$
$\leq(1+|\tau-\tau^{\prime}|)\big\{e^{2\tau+2\tau^{\prime}}k_{4}+(e^{2\tau}+e^{2\tau^{\prime}})k_{2}\big\}\,$, (3.51)
where we defined
$k_{2l}=\frac{5}{4}\sum_{n\neq 0}|f_{n}||g_{n}|\,\,n^{2l}\,,\quad l\in\mathbb{N}\,.$ (3.52)
For illustration, we show in Fig. 3 the actual behavior of the difference for the two-point functions in the Bunch-Davies vacuum, as detailed in Section 4.
Figure 3: Decay of the difference $\|\mathcal{W}^{s}-\mathfrak{W}^{s}\|_{\infty}$ in the Bunch–Davies vacuum state, plotted as a function of the two time variables $\tau$ and $\tau^{\prime}$.
4.  Hadamard states and composite operators
The concept of Hadamard states normally only applies to (free) quantum field theories on a curved, non-dynamical background. In the present context, the metric components of the Gowdy cosmologies themselves are treated as quantum fields, so the notion of a Hadamard state is not directly applicable. Nevertheless, the basic wave equation for the Gowdy scalar $\phi$ in (LABEL:evolpt) "looks as if" it lived on a $1+1$ dimensional time-dependent background set by $\rho$. In this sense, the notion of a Hadamard state is applicable to the quantized $\phi$ fields. In the following we argue that the Fock vacuum entering in (2.27) should in fact be associated with a Hadamard state.
4.1 Hadamard condition in Fourier space
Hadamard states on a globally hyperbolic background admit a characterization in terms of a local parametrix defined in terms of the Synge function, see e.g. [25, 26], but constructing them is difficult. On spatially homogeneous backgrounds (with spatial sections diffeomorphic to $\mathbb{R}^{d}$), one can alternatively characterize and construct Hadamard states in Fourier space.
Result[27]. Let $T_{p}$ be a Wronskian‑normalized solution of $\bigl[\partial_{\tau}^{2}+\nu(\tau)+|p|^{2}w(\tau)\bigr]T_{p}(\tau)=0$, with $\nu,w\in C^{\infty}(\mathbb{R})$, $w>0$, and $p=(p_{1},\ldots,p_{d})$ the spatial momentum. Define the two‑point function
$\varpi(\tau,x;\tau^{\prime},x^{\prime})=\int\!\frac{d^{d}p}{(2\pi)^{d}}\,T_{p}(\tau)\,T_{p}(\tau^{\prime})^{*}\,e^{ip\cdot(x-x^{\prime})}.$ (4.1)
Then $\varpi$ is associated with a Hadamard state if and only if $|T_{p}(\tau)|^{2}$ admits an asymptotic expansion of the form
$|T_{p}(\tau)|^{2}\asymp\frac{1}{2\sqrt{w}|p|}\Big\{1+\sum_{n\geq 1}(-)^{n}G_{n}(\tau)|p|^{-2n}\Big\}\,$, (4.2)
where the $G_{n}$ are known functionals of $\nu,w$. In particular,
$G_{1}$ $\!=\!$ $\frac{\nu}{2w}+\frac{5}{32}\frac{{w^{\prime}}^{2}}{w^{3}}-\frac{1}{8}\frac{w^{\prime\prime}}{w^{2}}\,$,
$G_{2}$ $\!=\!$ $\frac{3}{8w^{2}}\Big(\nu^{2}+\frac{1}{3}\nu^{\prime\prime}\Big)-\frac{5}{16w^{3}}\Big(\nu w^{\prime\prime}+\nu^{\prime}w^{\prime}-\nu\frac{7{w^{\prime}}^{2}}{4w}\Big)$ (4.3)
$+$ $\frac{1}{32w^{3}}\Big(\!-w^{(4)}+\frac{21{w^{\prime\prime}}^{2}}{4w}+\frac{7w^{(3)}w^{\prime}}{w}-\frac{231{w^{\prime}}^{2}w^{\prime\prime}}{8w^{2}}+\frac{1155{w^{\prime}}^{4}}{64w^{3}}\Big)\,.$
The asymptotics (4.2) is consistent with the one induced by the adiabatic iteration but much streamlined, see [28, 29] and Appendix A of [27]. Heuristically, Hadamard states are adiabatic vacua of infinite order. Mathematically, this is expressed by Lemma III.2 of [27] which we assume to extend to generic $d\geq 1$. Importantly, the Fourier space Hadamard condition (4.2) only prescribes the large $|p|$ asymptotics of $|T_{p}(\tau)|^{2}$. Different Hadamard states will differ by their non-universal small $|p|$ behavior. It is commonly assumed that a state is also infrared finite, which requires $|T_{p}(\tau)|^{2}=o(|p|^{-\delta})$, $\delta<d$, for the integral to be infrared finite. For example, the States of Low Energy are Hadamard states with $\delta=1$ for all $d\geq 1$, so they are infrared finite for $d\geq 2$ [27].
Two prominent Hadamard states for the Gowdy scalar. For $d=1$, $\nu=0$, $w=e^{2\tau}$ the above result entails that $T_{n}(\tau)$ is associated with a Hadamard state if and only if $|T_{n}(\tau)|^{2}$ has the a large $|n|$ asymptotic expansion of the form
$|T_{n}(\tau)|^{2}\asymp\frac{1}{2e^{\tau}|n|}\Big\{1-\frac{1}{8}\frac{e^{-2\tau}}{n^{2}}+\frac{27}{128}\frac{e^{-4\tau}}{n^{4}}+O(n^{-6})\Big\}\,$, (4.4)
where the coefficients turn out to be those of $(\pi/4)|H_{0}^{(2)}(|n|e^{\tau})|^{2}$. Here we transitioned to discrete momenta and circular spatial sections. This is expected to be legitimate on general grounds, but in the situation at hand it follows more directly from an identity of the form
$\sum_{n=-N}^{N}2\pi W^{s}_{\mathbb{R}}(\tau,\tau^{\prime},\zeta-\zeta^{\prime}+2\pi n)-[W^{s}(\tau,\tau^{\prime},\zeta-\zeta^{\prime})-W_{0}(\tau,\tau^{\prime})]=c_{N}(\tau,\tau^{\prime})\,.$ (4.5)
Here $W^{s}_{\mathbb{R}}$ is the symmetric two-point function for noncompact spatial sections and the term in square brackets is the previous two-point function $W^{s}$ from (LABEL:Wsdef) defined on $S^{1}$, with the zero mode term $W_{0}(\tau,\tau^{\prime})$ removed. The function $c_{N}(\tau,\tau^{\prime})$ is explicitly computable and is regular at $\tau=\tau^{\prime}$. It diverges for $N\rightarrow\infty$ in a way that depends on the infrared properties of $W^{s}_{\mathbb{R}}$’s Fourier transform. In $d=1$ one needs this Fourier transform to scale as $p^{-\delta},0\leq\delta<1$, for the integral to be infrared finite. In this case the leading divergence is $O(N^{\delta}\ln^{2}N)$. For $\delta=1$, relevant for States of Low Energy, the Fourier integral needs to be cut off at some small $0<\mu/(2N+1)\ll 1$. In this case the leading divergence is $O(N\ln^{2}\mu)$ or $O(N\ln^{2}N\ln\mu)$. Note that the periodic two-point function is always infrared finite as $W_{0}(\tau,\tau^{\prime})$ of the form in (LABEL:Wsdef) can be chosen at will. Based on (4.5) one can recover $W^{s}(\tau,\tau^{\prime},\zeta\!-\!\zeta^{\prime})-W_{0}(\tau,\tau^{\prime})$ as a subtracted limit of $2\pi W_{\mathbb{R}}^{s}$’s periodic extension. The notion of a Hadamard state thereby carries over from the noncompact to the compact spatial sections without having to construct the parametrix in the latter case. We now discuss two important instances of so-understood Hadamard states for the $T^{3}$-Gowdy system.
(Gowdy) Bunch–Davies vacuum. The choice $\lambda_{n}=1,\mu_{n}=0$, $n\neq 0$, in (2.3) gives
$T_{n}^{\rm BD}(\tau)=\frac{\sqrt{\pi}}{2}H_{0}^{(2)}(|n|e^{\tau})\,,\quad n\neq 0\,$, (4.6)
which satisfies (4.4). Augmented by the zero-mode choice $\tilde{\lambda}_{0}=1,\tilde{\mu}_{0}=0$ it defines a spatially periodic two-point function via (LABEL:Wsdef). The Fock vacuum associated with the decomposition (2.22) for this choice will be referred to as the ($T^{3}$-Gowdy) Bunch-Davies vacuum. The rationale for this terminology is as follows. Starting from the basic wave equation for $\phi$ in (LABEL:evolpt) with $\rho=t$ and redefining $\phi$ according to $\Phi(t,\zeta):=\sqrt{t}\phi(t,\zeta)$ leads to [13]
$\Big[\partial_{t}^{2}-\partial_{\zeta}^{2}+\frac{1}{4t^{2}}\Big]\Phi=0\,,\quad\Phi(t,\zeta):=\sqrt{t}\phi(t,\zeta)\,.$ (4.7)
This can be recognized as the wave equation of a free massive field on $1+1$ dimensional de Sitter space in the Poincaré patch and conformal time $t$, see e.g. [30]. Except for the late time limit $t\rightarrow\infty$ it is not conformally invariant. The mass parameter here is $m=1/2$, which is at the unitarity threshold. For noncompact spatial sections the two-point function $W_{\mathbb{R}}(t,t^{\prime},\zeta\!-\!\zeta^{\prime})$ can be computed explicitly and indeed coincides with the known $W_{{\rm dS}_{2}}(t,t^{\prime},\zeta\!-\!\zeta^{\prime})/\sqrt{tt^{\prime}}$, expressible in terms of an $_{2}F_{1}$ hypergeometric function for $m=1/2$ and with an appropriate $i\epsilon$-prescription. The symmetric part can be expressed as follows
$W^{s}_{\mathbb{R}}(t,t^{\prime},\zeta-\zeta^{\prime})=\frac{1}{\pi\sqrt{tt^{\prime}}}\Big\{\theta(-\xi)\frac{1}{\sqrt{1-\xi}}{\bf K}\Big(\frac{1}{1-\xi}\Big)+\theta(\xi){\bf K}(1-\xi)\Big\}\,.$ (4.8)
Here ${\bf K}(z)=\int_{0}^{\pi/2}\!ds(1-z\sin s)^{-1/2}$ is the complete elliptic integral and $\xi=[(\zeta-\zeta^{\prime})^{2}-(t-t^{\prime})^{2}]/(4tt^{\prime})$ is the de Sitter embedding distance. Near the lightcone $|t-t^{\prime}|=|\zeta-\zeta^{\prime}|$ the function $W^{s}_{\mathbb{R}}$ has a logarithmic singularity, at $t+t^{\prime}=|\zeta-\zeta^{\prime}|$ it is regular, and for $|t-t^{\prime}|\rightarrow\infty$ it behaves like $1/|t-t^{\prime}|$. The relation (4.5) can be directly verified in this case and used to (re-)construct the periodic two-point function (LABEL:Wsdef) with the above Bogoliubov parameters.
States of Low Energy (SLE). SLE are a class of Hadamard states that are explicitly constructable on any spatially homogeneous background. The original construction [31] was for Friedmann-Lemaître cosmologies; further properties and the extension to Bianchi I spacetimes can be found in [21, 27]. The term SLE stems from the fact that in Fourier space the temporarily averaged Hamiltonian is minimized modewise. The averaging is done with a $f\in C_{c}^{\infty}(\mathbb{R})$ ‘window’ function and the minimization produces a set of Bogoliubov parameters that depend on $f$ and the momenta. This construction can be applied to the Hamiltonian governing the dynamics of the Gowdy scalar in the reduced phase space formulation, see (4.23) below. Specifically, taking the wave equation (2.23) as basic the formulas in [21, 27] apply with the specialization to $d=1$ and $\omega_{n}(\tau)=|n|e^{\tau}$. The SLE solution can be expressed solely in terms of the commutator function, see Thm.II.2 of [27]. Using the expression (3.1) for the commutator function of the Gowdy system SLE solution for $G_{n}^{\rm SLE}(\tau)=|T_{n}^{\rm SLE}(\tau)|^{2}$ can be evaluated. The result is of the following form
$G_{n}(\tau)=\frac{\pi}{2}c_{1,n}J_{0}(|n|e^{\tau})^{2}+\frac{\pi}{2}c_{2,n}J_{0}(|n|e^{\tau})Y_{0}(|n|e^{\tau})+\frac{\pi}{2}c_{3,n}Y_{0}(|n|e^{\tau})^{2}\,$,
$c_{1,n}=\frac{1}{2}|\lambda_{n}+\mu_{n}|^{2}\,,\quad c_{3,n}=\frac{1}{2}|\lambda_{n}-\mu_{n}|^{2}\,,\quad c_{2,n}=-i(\lambda_{n}\mu_{n}^{*}-\lambda_{n}^{*}\mu_{n})\,$,
$4c_{1,n}c_{3,n}-c_{2,n}^{2}=(|\lambda_{n}|^{2}-|\mu_{n}|^{2})^{2}=1\,,\quad n\neq 0\,.$ (4.9)
For SLE the $n\neq 0$ coefficients are
$c_{1,n}^{f}$ $\!\!=\!\!$ $\frac{\pi n^{2}}{8{\cal E}_{n}^{\rm SLE}}\int\!d\tau f(\tau)^{2}e^{2\tau}\big[Y_{0}(|n|e^{\tau})^{2}+Y_{1}(|n|e^{\tau})^{2}\big]\,$,
$c_{2,n}^{f}$ $\!\!=\!\!$ $-\frac{\pi n^{2}}{4{\cal E}_{n}^{\rm SLE}}\int\!d\tau f(\tau)^{2}e^{2\tau}\big[J_{0}(|n|e^{\tau})Y_{0}(|n|e^{\tau})+J_{1}(|n|e^{\tau})Y_{1}(|n|e^{\tau})\big]\,$,
$c_{3,n}^{f}$ $\!\!=\!\!$ $\frac{\pi n^{2}}{8{\cal E}_{n}^{\rm SLE}}\int\!d\tau f(\tau)^{2}e^{2\tau}\big[J_{0}(|n|e^{\tau})^{2}+J_{1}(|n|e^{\tau})^{2}\big]\,.$ (4.10)
Here ${\cal E}_{n}^{\rm SLE}$ is the energy of the time averaged Hamiltonian in the SLE. By inverting the $c_{1,n},c_{2,n},c_{3,n}$ versus $\lambda_{n},\mu_{n}$ relations in (LABEL:GDGowdysol) one obtains expressions for the $f$-dependent Bogoliubov parameters $\lambda_{n}^{f},\mu_{n}^{f}$ that define the SLE solution $T_{n}^{\rm SLE}(\tau)$ in (2.3). The induced $n$-dependence is rather complicated and it is not obvious that $|T_{n}^{\rm SLE}(\tau)|^{2}$ will satisfy (4.4). Actually, it does, as one can show
$c_{1,n}^{f}\asymp 1/2,\quad c_{3,n}^{f}\asymp 1/2,\quad c_{2,n}^{f}\asymp 0,$ (4.11)
where $\asymp$’ means “equality in an asymptotic expansion in powers of 1/|n|1/|n|”. For $\lambda_{n}^{f},\mu_{n}^{f}$ this means $\mu_{n}^{f}\asymp 0$, $|\lambda_{n}^{f}|\asymp 1$, while the phase of $\lambda_{n}^{f}$ would require separate examination. This phase drops out in the asymptotics of $G_{n}^{\rm SLE}(\tau)=|T_{n}^{\rm SLE}(\tau)|^{2}$, which therefore has the same asymptotics as (4.6), i.e. (4.4).
In summary, both the Bunch-Davies state and the SLE define Hadamard states for the Gowdy scalar by specifying a set of Bogoliubov parameters in (2.3). Via (2.26) the Bogoliubov parameters define initial data, which enter the initial value parameterization of the two-point function (LABEL:twopmat2), (LABEL:twopmat3). The Result 3.1 delineates conditions under which these states are also time consistent.
4.2 Hadamard and time consistent states
Both the Bunch–Davies vacuum (BD) and State of Low Energy (SLE) are also time consistent. Technically, this is because their Bogoliubov parameters $(\lambda_{n},\mu_{n})$ are independent of the reference time $\tau_{0}$, albeit for distinct reasons.
The Bunch-Davies state corresponds in the parameterization (2.3) simply to $\lambda_{n}^{\rm BD}=1$, $\mu_{n}^{\rm BD}=0$, $n\in\mathbb{Z}$. Then (3.21), (2.31) gives for $n\neq 0$
$\tilde{\lambda}_{n}^{\rm BD}(\tau_{0})=1-\frac{i}{\pi}\ln\gamma_{n}(0)\,,\makebox[28.45274pt]{ }\;\;\tilde{\mu}_{n}^{\rm BD}(\tau_{0})=-\frac{i}{\pi}\ln\gamma_{n}(0)\,$,
$\tilde{z}_{n}^{\rm BD}(\tau_{0})=\frac{\sqrt{\pi}}{2}-\frac{i}{\sqrt{\pi}}\ln\gamma_{n}(\tau_{0})\,,\quad\tilde{w}_{n}^{\rm BD}(\tau_{0})=-\frac{i}{\sqrt{\pi}}\,.$ (4.12)
For $n=0$ one has from (3.27) $\tilde{\lambda}_{0}=1,\tilde{\mu}_{0}=0$ and (2.31) then gives
$\tilde{z}_{0}(\tau_{0})=\frac{\sqrt{\pi}}{2}-i\frac{\tau_{0}}{\sqrt{\pi}}\,,\quad\tilde{w}_{0}(\tau_{0})=-\frac{i}{\sqrt{\pi}}\,.$ (4.13)
The SLE are likewise time consistent. Indeed, according to the result in Section 3.2 this should hold because the parameters $c_{1,n}^{f},c_{2,n}^{f},c_{3,n}^{f}$ defined in (4.1) are independent of $\tau_{0}$. Returning to (LABEL:GDGowdysol), (4.1) and considering the $\tau\rightarrow$-$\infty$ behavior of the solution $G_{n}^{\rm SLE}(\tau)$ one finds for $n\neq 0$
$G_{n}^{\rm SLE}(\tau)$ $\!\!=\!\!$ $\tilde{c}_{3,n}\tau^{2}+\tilde{c}_{2,n}\tau+\tilde{c}_{1,n}+O(\tau e^{2\tau})\,$,
$\tilde{c}_{3,n}$ $\!\!=\!\!$ $\frac{2}{\pi}c_{3,n}^{f}\,$,
$\tilde{c}_{2,n}$ $\!\!=\!\!$ $c_{2,n}^{f}+c_{3,n}^{f}\frac{4}{\pi}\ln(e^{\gamma_{E}}|n|/2)\,$,
$\tilde{c}_{1,n}$ $\!\!=\!\!$ $\frac{\pi}{2}c_{1,n}^{f}+c_{2,n}^{f}\ln(e^{\gamma_{E}}|n|/2)+c_{3,n}^{f}\frac{2}{\pi}\ln^{2}(e^{\gamma_{E}}|n|/2)\,.$ (4.14)
For $n=0$ the SLE parameters in (4.1) are not really defined and we omit them for now. For the induced Bogoliubov parameters one has
$\tilde{\lambda}_{n}^{f}$ $\!\!=\!\!$ $\lambda_{n}^{f}-\frac{i}{\pi}(\lambda_{n}^{f}\!-\!\mu_{n}^{f})\ln\gamma_{n}(0)\,$,
$\tilde{\mu}_{n}^{f}$ $\!\!=\!\!$ $\mu_{n}^{f}-\frac{i}{\pi}(\lambda_{n}^{f}\!-\!\mu_{n}^{f})\ln\gamma_{n}(0)\,$ (4.15)
and the velocity dominated initial values are (up to a shared phase that may be needed to have the correct large $n$ asymptotic expansion) one has
$\tilde{z}_{n}^{f}(\tau_{0})$ $\!\!=\!\!$ $\frac{\sqrt{\pi}}{2}(\lambda_{n}^{f}+\mu_{n}^{f})-\frac{i}{\sqrt{\pi}}\ln\gamma_{n}(\tau_{0})(\lambda_{n}^{f}-\mu_{n}^{f})\,$,
$\tilde{w}_{n}^{f}(\tau_{0})$ $\!\!=\!\!$ $-\frac{i}{\sqrt{\pi}}(\lambda_{n}^{f}-\mu_{n}^{f})\,.$ (4.16)
There are of course also Hadamard states which are not time consistent simply because the Bogoliubov parameters depend on the initial value time $\tau_{0}$. For example
$\lambda_{n}(\tau_{0})=\big(1+e^{-\frac{n^{2}}{2\tau_{0}}}\big)^{1/2}\,,\quad\mu_{n}(\tau_{0})=e^{-\frac{n^{2}}{2\tau_{0}}}\,$, (4.17)
differs from the BD parameters $\lambda_{n}^{\rm BD}=1,\mu_{n}^{\rm BD}=0$ by $\tau_{0}$-dependent terms that do not have an expansion in powers of $1/n$. It is therefore a Hadamard state but not time consistent. The interplay between time consistent states and Hadamard states is schematically summarized in Figure 4.
time‐consistentHadamardBDSLE
Figure 4: Intersection of time‐consistent and Hadamard states.
Using the notion of time consistent states as introduced in Subsection 3.2 the AVD Results 3.2, 3.3 can be paraphrased as follows: all time consistent states give rise to (matrix-) two-point functions that are asymptotically velocity dominated in the sense of (LABEL:avdresult). These states include many Hadamard states, in particular the Bunch-Davies vacuum and the States of Low Energy.
4.3 Quantum $\sigma$‑field
So far, our analysis has focused on the Gowdy scalar $\phi$, canonically quantized as in (2.22). The gravitational origin of the classical system is encoded in the constraints (2.4). As noted in (2.1) they can be solved for the gradient of $\sigma$. In proper time gauge and $(t=e^{\tau},\zeta)$ coordinates the gradient in (LABEL:evolpt) simplifies to
$F_{0}^{\rm pt}(e^{\tau},\phi)$ $\!\!=\!\!$ $\frac{1}{2}[(\partial_{\tau}\phi)^{2}+e^{2\tau}(\partial_{\zeta}\phi)^{2}]=:T_{00}=T_{11}\,.$
$F_{1}^{\rm pt}(e^{\tau},\phi)$ $\!\!=\!\!$ $\partial_{\tau}\phi\partial_{\zeta}\phi=:T_{01}=T_{10}\,.$ (4.18)
Here, we introduced the "would-be energy–momentum tensor" $T_{\mu\nu}$ of $\phi$ in the time dependent background. For on-shell $\phi$ it obeys the partial conservation equation $\partial_{\tau}T_{01}=\partial_{\zeta}T_{11}$. This ensures that the gradient formulas can consistently be integrated to give [12]
$\sigma(\tau,\zeta)=\sigma(\tau_{0},\zeta_{0})+\int_{\tau_{0}}^{\tau}\!d\tau^{\prime}\,T_{00}(\tau^{\prime},\zeta)+\int_{\zeta_{0}}^{\zeta}\!d\zeta^{\prime}\,T_{01}(\tau_{0},\zeta^{\prime})\,.$ (4.19)
Since $T_{00}=T_{11}$ is spatially periodic the last relation implies that $C_{0}$ below is conserved, $\partial_{\tau}C_{0}=0$. On the other hand, one sees from (4.19) that in order for $\sigma(\tau_{0},\zeta)=\sigma(\tau_{0},\zeta+2\pi)$ to hold one needs $C_{0}=0$ at $\tau=\tau_{0}$. Together,
$C_{0}\stackrel{{\scriptstyle\displaystyle{!}}}{{=}}0\,,\quad C_{0}:=\int_{0}^{2\pi}\frac{d\zeta}{2\pi}\,T_{01}\,.$ (4.20)
This is the remnant of the original constraints that still needs to be incorporated into the solution for $\phi$. In the quantum theory, one must define the field $\sigma$—that is, the components $T_{00}$ and $T_{01}$—as composite operators satisfying the partial conservation law and the constraint (4.20). An adiabatic renormalization [28] of $T_{00}$ has been considered early on in [11].
For later use we prepare the mode expansions
$T_{00}$ $\!\!=\!\!$ $\sum_{n\in\mathbb{Z}}e^{in\zeta}(T_{00})_{n}(\tau)\,,\quad T_{01}=\sum_{n\in\mathbb{Z}}e^{in\zeta}(T_{01})_{n}(\tau)\,$,
$(T_{00})_{n}(\tau)$ $\!\!=\!\!$ $\frac{1}{2}\sum_{m}\big[\dot{\phi}_{n-m}\dot{\phi}_{m}-(n-m)m\,e^{2\tau}\phi_{n-m}\phi_{m}\big]$
$(T_{01})_{n}(\tau)$ $\!\!=\!\!$ $i\sum_{m}m\,\dot{\phi}_{n-m}\phi_{m}\,.$ (4.21)
Here, the Fourier modes $\phi_{m}=\phi_{m}(\tau)$ are functions of $\tau$ only and we often write $\dot{\phi}_{m}=\partial_{\tau}\phi_{m}$. The consistency condition needed is
$\partial_{\tau}(T_{01})_{n}=in(T_{00})_{n}\,,\quad n\in\mathbb{Z}\,$, (4.22)
which indeed holds on account of $\ddot{\phi}_{n}=-n^{2}\phi_{n}$.
These relations remain valid upon canonical quantization, provided that the ordering of the Fock space operators is respected and the operator sums are treated as formal. For example, the time dependent Hamilton operator is
$\mathbb{H}(\tau)$ $\!\!=\!\!$ $\frac{1}{2}\int_{0}^{2\pi}\!\frac{d\zeta}{2\pi}T_{00}(\tau,\zeta)=\frac{1}{2}(T_{00})_{0}(\tau)=$ (4.23)
$\!\!=\!\!$ $\frac{1}{2}\sum_{n\in\mathbb{Z}}\big\{\varepsilon_{n}(\tau)(a_{n}a_{n}^{*}+a_{-n}^{*}a_{-n})+d_{n}(\tau)a_{n}a_{-n}+d_{n}(\tau)^{*}a_{-n}^{*}a_{n}^{*}\big\}\,$,
with $\varepsilon_{n}(\tau)=|\partial_{\tau}T_{n}|^{2}+n^{2}e^{2\tau}|T_{n}(\tau)|^{2}$, $d_{n}(\tau)=(\partial_{\tau}T_{n})^{2}+n^{2}e^{2\tau}T_{n}(\tau)^{2}$. The zero-mode part of the $\sigma$ operator is by (4.19) then formally given by integrating $\partial_{\tau}\int\!\frac{d\zeta}{2\pi}\sigma(\tau,\zeta)=2\mathbb{H}(\tau)$. Of course, these formal sums cannot act on the Fock space and some regularization and/or renormalization is needed.
Hadamard states are tailored towards renormalization of quantum field theories in curved spacetimes [25, 26]. In the standard setting with a spatial slices diffeomorphic to $\mathbb{R}^{d}$ one can render the two‐point function finite by subtracting its singular “Hadamard parametrix” in position space—i.e. the part singular on the lightcone—from the full two‐point function. For spatially homogeneous backgrounds subtraction of a truncated Hadamard parametrix is consistent with the older adiabatic renormalization, see e.g. [29]. The precise relation has been elucidated in [24], see also Appendix A of [27].
In the polarized Gowdy model the spatial slices are circles, so the light‐cone singularity acquires infinitely many $2\pi$–periodic images, see (4.5). Therefore, it is more convenient to work in Fourier space and parameterize the two‐point function using its initial data $Z_{n}(\tau_{0})$. Since Hadamard states are singled out by the universal large $n$ asymptotics (4.4) of $|T_{n}(\tau)|^{2}$, it follows that the initial–value matrix $Z_{n}(\tau_{0})$ itself admits a universal large $n$ expansion for every Hadamard state. Explicitly,
$Z_{n}(\tau_{0})\asymp\left.\begin{pmatrix}\frac{2}{x}-\frac{1}{4x^{3}}&-\frac{1}{x}+\frac{3}{8}\frac{1}{x^{3}}\\[5.69054pt] -\frac{1}{x}+\frac{3}{8}\frac{1}{x^{3}}&\frac{x}{2}+\frac{9}{16}\frac{1}{x}-\frac{105}{256}\frac{1}{x^{3}}\end{pmatrix}\right|_{x=|n|e^{\tau_{0}}}+O(n^{-5})\,.$ (4.24)
We denote by $Z_{n}^{\rm Hr}(\tau_{0})$ the matrix obtained by truncating the large $n$ expansion of $Z_{n}(\tau_{0})$ up to and including terms of order $\mathcal{O}(n^{-r})$; the superscript “Hr” stands for “Hadamard expansion truncated at order rr.” Next, we consider point-split definitions of the vacuum expectation values of $T_{00}$ and $T_{01}$ from (4.3), aimed at:
$T_{00}^{W}(\tau,\tau^{\prime},\zeta\!-\!\zeta^{\prime}):=\frac{1}{4}\big[\partial_{\tau}\partial_{\tau^{\prime}}+e^{\tau+\tau^{\prime}}\partial_{\zeta}\partial_{\zeta^{\prime}}\big]W^{s}(\tau,\tau^{\prime},\zeta\!-\!\zeta^{\prime})$
$\quad=\frac{1}{4}W_{0}^{s}(\tau,\tau^{\prime})+\frac{1}{2}\sum_{n\geq 1}\cos n(\zeta\!-\!\zeta^{\prime})\big[\partial_{\tau}\partial_{\tau^{\prime}}W_{n}^{s}(\tau,\tau^{\prime})+e^{\tau+\tau^{\prime}}n^{2}W_{n}^{s}(\tau,\tau^{\prime})\big]\,$,
$T_{01}^{W}(\tau,\tau^{\prime},\zeta\!-\!\zeta^{\prime}):=\frac{1}{2}\partial_{\tau}\partial_{\zeta^{\prime}}W^{s}(\tau,\tau^{\prime},\zeta\!-\!\zeta^{\prime})$
$\quad=\sum_{n\geq 1}n\sin n(\zeta\!-\!\zeta^{\prime})\partial_{\tau}W_{n}^{s}(\tau,\tau^{\prime})\,$, (4.25)
where $W_{n}^{s}(\tau,\tau^{\prime}):=T_{n}(\tau)T_{n}(\tau^{\prime})^{*}+T_{n}(\tau)^{*}T_{n}(\tau^{\prime})$, $n\in\mathbb{Z}$. They obey the quasi-conservation equation
$\partial_{\zeta}T_{00}^{W}(\tau,\tau^{\prime},\zeta\!-\!\zeta^{\prime})=\frac{1}{2}\big(\!\!-\partial_{\tau^{\prime}}+e^{-\tau+\tau^{\prime}}\partial_{\tau}\big)\,T_{01}^{W}(\tau,\tau^{\prime},\zeta\!-\!\zeta^{\prime})\,$, (4.26)
which we seek to preserve after renormalization.
To proceed, we recall the initial data realization of the symmetric two point function
$W_{n}^{s}(\tau,\tau^{\prime})=\big(\!\!-\partial_{\tau_{0}}\Delta_{n}(\tau,\tau_{0}),\Delta_{n}(\tau,\tau_{0})\big)Z_{n}(\tau_{0})\begin{pmatrix}-\partial_{\tau_{0}}\Delta_{n}(\tau^{\prime},\tau_{0})\\[5.69054pt] \Delta_{n}(\tau^{\prime},\tau_{0})\end{pmatrix}\,$. (4.27)
Used in the Fourier kernels $\partial_{\tau}\partial_{\tau^{\prime}}W_{n}^{s}(\tau,\tau^{\prime})+e^{\tau+\tau^{\prime}}n^{2}W_{n}^{s}(\tau,\tau^{\prime})$ and $\partial_{\tau}W_{n}^{s}(\tau,\tau^{\prime})$, one obtains realizations of $T^{W}_{00}$ and $T^{W}_{01}$ in terms of the commutator function and the initial value matrix $Z_{n}(\tau_{0})$. Importantly, subtractions that are carried out on the level of the initial value matrix will not affect the time dependence and in particular preserve the validity of (4.26). Our proposal for a “truncated Hadamard renormalization” in this context is to replace $Z_{n}(\tau_{0})$ in (4.27) with
$Z_{n}^{\rm subHr}(\tau_{0}):=Z_{n}(\tau_{0})-Z_{n}^{\rm Hr}(\tau_{0})\,$, (4.28)
for a suitable order $r\in\mathbb{N}$. From (LABEL:PPtwop3) one recalls that pointwise in $\tau,\tau_{0}$ the leading large $n$ behavior of $\Delta_{n}(\tau,\tau_{0})$ is $O(1/n)$, that of $\partial_{\tau_{0}}\Delta_{n}(\tau,\tau_{0})$ is $O(1)$, and that of $\partial_{\tau_{0}}\partial_{\tau}\Delta_{n}(\tau,\tau_{0})$ is $O(n)$. For the three relevant Fourier kernels $W_{n}^{s}(\tau,\tau^{\prime})$, $n\partial_{\tau}W_{n}^{s}(\tau,\tau^{\prime})$, and $\partial_{\tau}\partial_{\tau^{\prime}}W_{n}^{s}(\tau,\tau^{\prime})+e^{\tau+\tau^{\prime}}n^{2}W_{n}^{s}(\tau,\tau^{\prime})$, we need a subtraction that produces a leading decay of $O(1/n^{2})$ in order to ensure absolute convergence of the sums. Inspection of the matrix multiplications in (4.27) and its derivatives shows that in order to obtain coincidence limits represented by absolutely convergent series the following minimal subtractions are needed:
$\phi^{2}:\;\;r=1\,,\quad T_{01}:\;\;r=2\,,\quad T_{00}:\;\;r=2\,.$ (4.29)
We denote the subtracted Fourier kernels by ${\rm kernel}]^{\rm subHr}$. After the subtraction they decay at least like $1/n^{2}$, pointwise in $\tau,\tau^{\prime}$, including $\tau^{\prime}=\tau$. The $\zeta^{\prime}\rightarrow\zeta$ limit is therefore well-defined and by common abuse of notation we interpret the limit as the matrix element $\langle 0|:\!{\rm operator}\!:_{\rm Hr}|0\rangle$ of the composite operator aimed at (even if the existence of the operator has not yet fully been justified, see [25]). In this notation one has
$\langle 0|:\!\phi(\tau,\zeta)^{2}\!:_{H1}|0\rangle$ $\!\!=\!\!$ $W_{0}^{s}(\tau,\tau)+2\sum_{n\geq 1}[W_{n}^{s}]^{\rm subH1}(\tau,\tau)\,.$
$\langle 0|:\!T_{00}(\tau,\zeta)\!:_{H2}|0\rangle$ $\!\!=\!\!$ $\frac{1}{4}\partial_{\tau}\partial_{\tau^{\prime}}W_{0}^{s}(\tau,\tau^{\prime})\big|_{\tau^{\prime}=\tau}+\sum_{n\geq 1}[\partial_{\tau}\partial_{\tau^{\prime}}W_{n}^{s}+e^{\tau+\tau^{\prime}}n^{2}W_{n}^{s}]^{\rm subH2}(\tau,\tau)\,.$
$\langle 0|:\!T_{01}(\tau,\zeta)\!:_{H2}|0\rangle$ $\!\!=\!\!$ $0\,.$ (4.30)
The first expression defines the counterpart of the “power spectrum”, the second is independent of $\zeta$ on account of translation invariance, and the third is consistent with the extension of (4.26) to the coincidence limit. Alternatively, the matrix element of the $n=0$ version of (4.22) should at least be constant, $\partial_{\tau}\langle 0|\int_{0}^{2\pi}\!\frac{d\zeta}{2\pi}:\!T_{01}(\tau,\zeta)\!:_{H2}|0\rangle=0$.
The vanishing of this constant extents to higher order subtractions
$\langle 0|:\!T_{01}(\tau,\zeta)\!:_{Hr}|0\rangle=0\,,\quad r\geq 2\,.$ (4.31)
This is because, once $n\partial_{\tau}W_{n}^{s}]^{\rm subHr}(\tau,\tau)$ is absolutely summable, its Fourier-sine series will vanish in the $\zeta\rightarrow\zeta^{\prime}$ coincidence limit. Normally, the Fourier subtractions (4.28) cannot be pushed to arbitrarily high orders, as the coefficients in (4.24) are rapidly increasing beyond $r=10$, say, reflecting the merely asymptotic nature of the expansion. The case (4.31) is exceptional in that the large coefficients are rendered irrelevant by the $\zeta^{\prime}\rightarrow\zeta$ limit of the convergent sine-Fourier series. For (4.31) we can therefore conclude that the all order subtraction, representing the actual Hadamard subtraction, also vanishes, $\langle 0|\!:\!T_{01}(\tau,\zeta)\!:_{\rm H}\!|0\rangle=0$. Recall that the Hadamard normal product would be defined (after stripping off test functions) by the coincidence limit of
$\partial_{\tau}\phi(\tau,\zeta)\partial_{\zeta^{\prime}}\phi(\tau^{\prime},\zeta^{\prime})\!:_{H}\,=\partial_{\tau}\phi(\tau,\zeta)\partial_{\zeta^{\prime}}\phi(\tau^{\prime},\zeta^{\prime})-T_{01}^{W^{\rm subH}}(\tau,\tau^{\prime},\zeta-\zeta^{\prime})\,$, (4.32)
where $T_{01}^{W^{\rm subH}}(\tau,\tau^{\prime},\zeta-\zeta^{\prime})$ derives from $T_{01}^{W}(\tau,\tau^{\prime},\zeta\!-\!\zeta^{\prime})=\frac{1}{2}\partial_{\tau}\partial_{\zeta^{\prime}}W^{s}(\tau,\tau^{\prime},\zeta\!-\!\zeta^{\prime})$ by subtracting the full position space Hadamard parametrix from $W^{s}$; see e.g. [25]. In the coincidence limit the subtraction is $\langle 0|\!:\!T_{01}(\tau,\zeta)\!:_{\rm H}\!|0\rangle$, so its vanishing leads to the “non-renormalization” result
$\lim_{\tau^{\prime}\rightarrow\tau,\zeta^{\prime}\rightarrow\zeta}\big\{:\!\partial_{\tau}\phi(\tau,\zeta)\partial_{\zeta^{\prime}}\phi(\tau^{\prime},\zeta^{\prime})+\partial_{\zeta^{\prime}}\phi(\tau^{\prime},\zeta^{\prime})\partial_{\tau}\phi(\tau,\zeta)\!:_{H}\big\}$
$\quad=\lim_{\tau^{\prime}\rightarrow\tau,\zeta^{\prime}\rightarrow\zeta}\big\{\partial_{\tau}\phi(\tau,\zeta)\partial_{\zeta^{\prime}}\phi(\tau^{\prime},\zeta^{\prime})+\partial_{\zeta^{\prime}}\phi(\tau^{\prime},\zeta^{\prime})\partial_{\tau}\phi(\tau,\zeta)\big\}\,.$ (4.33)
As a heuristic check one may consider the naive quantum version of the Fourier modes $(T_{01})_{n}$ from (4.3), written as $(T_{01})_{n}=\frac{1}{2}\sum_{m}m[\dot{\phi}_{n-m}\phi_{m}+\phi_{m}\dot{\phi}_{n-m}]$, with $\phi_{n}(\tau)=T_{n}(\tau)a_{n}+T_{n}(\tau)^{*}a_{-n}^{*}$, $\dot{\phi}_{n}(\tau)=\dot{T_{n}}(\tau)a_{n}+\dot{T_{n}}(\tau)^{*}a_{-n}^{*}$. Clearly, for $n\neq 0$ there are no ordering ambiguities and the termwise interpreted expectation value (being linear in $a_{n-m},a_{-(n-m)}^{*}$ and $a_{m},a_{-m}^{*},n\neq 0)$ in the Fock vacuum simply vanishes. For $n=0$ it is the antisymmetry in the $\sum_{m}m$ sum (corresponding to the $\sin m(\zeta-\zeta^{\prime})$ structure in the point-split version) that removes the symmetric terms containing $a_{m}a_{-m}$, $a_{-m}^{*}a_{m}^{*}$, as well as terms arising by reordering $a_{m}^{*}a_{m}$. The upshot is that the quantum version of the conserved charge $C_{0}$ can reasonably be interpreted as
$C_{0}=\,:\!C_{0}\!:_{H}\;=-i\lim_{N\rightarrow\infty}\sum_{n=-N}^{N}n\,a_{n}^{*}a_{n}\,.$ (4.34)
Its vacuum expectation value vanishes and so does therefore that of $T_{01}(\tau,\zeta)\!\!:_{H}=T_{01}(\tau,\zeta)=\sum_{n}e^{in\zeta}(T_{01})_{n}$. Further, one may verify that
$iC_{0},\phi(\tau,\zeta)]=\partial_{\zeta}\phi(\tau,\zeta)\,,\quad[iC_{0},\mathbb{H}(\tau)]=0\,$, (4.35)
where $\mathbb{H}(\tau)$ is the above Fock space Hamiltonian. The operator $C_{0}$ can be shown to be selfadjoint on a dense (finite occupation number) domain of the Fock space and by Stone’s theorem therefore generates a unitary group $\mathbb{R}\ni\zeta\mapsto e^{i\zeta C_{0}}$. It acts on the Fourier modes in the expected way
$e^{i\zeta C_{0}}\phi_{n}(\tau)e^{-i\zeta C_{0}}=\sum_{n\geq 0}\frac{\zeta^{n}}{n!}{\rm ad}^{n}(iC_{0})\phi_{n}(\tau)=e^{in\zeta}\phi_{n}(\tau)\,$, (4.36)
and similarly for $\pi_{n}(\tau)$. Here ${\rm ad}(iC_{0})A=[iC_{0},A]$, ${\rm ad}^{n}(iC_{0})A=[iC_{0},{\rm ad}^{n-1}(iC_{0})A]$, $n\geq 1$. In particular, for $\zeta=2\pi$ the operator $e^{i2\pi C_{0}}$ acts like the identity on the field algebra generated by (the $2\pi$-periodic) $\phi(\tau,\zeta),\pi(\tau,\zeta)$. Finally, $e^{i\zeta C_{0}}|0\rangle=|0\rangle$, expresses the translation invariance of the Fock vacuum, and accounts for the dependence of the two-point functions considered on $\zeta\!-\!\zeta^{\prime}$ only.
5.  Conclusions
Gowdy cosmologies have played an important role in shaping our understanding of the Big Bang singularity, specifically through proofs of Asymptotic Velocity Domination (AVD) in the polarized [3] and unpolarized case [4]. Here we investigated for the polarized system a quantum version of the AVD property, formulated in terms of a matrix two-point function ${\cal W}^{s}$ of the basic Gowdy scalar. Since the latter governs via (LABEL:diractwopt2) also the two-point functions of the integrands of Dirac observables the main question is whether ${\cal W}^{s}$ exhibits some form of AVD. This was answered in the affirmative through the results (LABEL:grad20), (LABEL:avdresult). We take this as a ‘proof of principle’ for the viability of a quantum AVD scenario. Hopefully it will turn out to admit as significant generalizations as the seminal classical result in [3].
The immediate extension is of course to the non-polarized Gowdy system. A formal expansion of classical solutions around those of the velocity dominated system has been obtained in [32]. The construction of classical Dirac observables is feasible by taking advantage of an underlying Lax pair [33]. The required spatial periodicity (for the $T^{3}$ topology) however presents considerable complications compared to other two-Killing vector reductions. Through a Riemannian sigma-model formulation the quantum theory is amenable to an all-order perturbative (weak Newton coupling) analysis [8, 9, 10]. The quantum AVD property, on the other hand, is likely to be related to an expansion in inverse powers of Newton’s constant (strong coupling) and requires a different methodology.
Acknowledgments. We would like to thank R. Banerjee and R. Penna for discussions and valuable comments on the manuscript. We also appreciate W. Musk’s and C. Ives’ participation in other aspects of this project. This work is supported in part by the PittPACC initiative.
Appendix A The Hamiltonian action and its VD counterpart
It is a general property of two-Killing vector reductions that entering with (2.1) into the Einstein-Hilbert action will produce a valid reduced action principle, i.e. one whose variations produce field equation coinciding with the ones obtained by directly specializing the Einstein field equations to metrics of the form (2.1). We therefore take the reduced Hamiltonian action of the polarized $T^{3}$ system as a starting point, see [13, 15]. The velocity dominated system is normally only defined through its field equations, see [3, 5]. In preparation for the quantum theory we develop here also an off-shell formulation in terms of an Hamiltonian action principle and its symmetries.
A.1 Gowdy cosmologies: the Hamiltonian action and its symmetries
In some fiducial foliation the Hamiltonian action reads
$S^{{\textsc{h}}}$ $\!\!=\!\!$ $\int_{t_{i}}^{t_{f}}\!\!dx^{0}\int_{0}^{2\pi}\!\!dx^{1}\big\{\pi^{\rho}e_{0}(\rho)+\pi^{\sigma}e_{0}(\sigma)+\pi^{\phi}e_{0}(\phi)-n{\cal H}_{0}\big\}\,$,
${\cal H}_{0}$ $\!\!=\!\!$ $-\lambda_{{\textsc{n}}}\pi^{\sigma}\pi^{\rho}-\frac{1}{\lambda_{{\textsc{n}}}}(\partial_{1}\rho\partial_{1}\sigma-2\partial_{1}^{2}\rho)+\frac{\lambda_{{\textsc{n}}}}{2\rho}(\pi^{\phi})^{2}+\frac{\rho}{2\lambda_{{\textsc{n}}}}(\partial_{1}\phi)^{2}\,$,
${\cal H}_{1}$ $\!\!=\!\!$ $\pi^{\rho}\partial_{1}\rho+\pi^{\sigma}\partial_{1}\sigma-2\partial_{1}\pi^{\sigma}+\pi^{\phi}\partial_{1}\phi\,$, (A.1)
where $e_{0}=\partial_{0}-{\cal L}_{s}$, with ${\cal L}_{s}$ the Lie derivative along the one-dimensional shift $s$. Upon integration by parts one can render the $s$ dependence explicit to find $s{\cal H}_{1}$, with ${\cal H}_{1}$ as given. For definiteness we identify $S^{1}$ with $\mathbb{R}/(2\pi\mathbb{Z})$ and assume all fields in (A.1) to be $2\pi$-periodic in $x^{1}$. Further, $\lambda_{{\textsc{n}}}>0$ is the dimensionless reduced Newton constant and $\sigma:=\tilde{\sigma}+\frac{1}{2}\ln\rho$. The phase space is equipped with a Poisson structure, $\{\,\cdot\,$$\cdot\,\}$ determined by the basic brackets, $\{\rho(x^{1}),\pi^{\rho}(y^{1})\}=\delta(x^{1}\!-\!y^{1})$, $\{\sigma(x^{1}),\pi^{\sigma}(y^{1})\}=\delta(x^{1}\!-\!y^{1})$, $\{\phi(x^{1}),\pi^{\phi}(y^{1})\}=\delta(x^{1}\!-\!y^{1})$, where the equal $x^{0}$ arguments are suppressed and all $\delta$ distributions are spatial +1+1 densities.
The constraints obey a closed Poisson algebra
$\{{\cal H}_{0}(x),{\cal H}_{1}(y)\}$ $\!\!=\!\!$ $\partial_{1}{\cal H}_{0}\delta(x-y)+2{\cal H}_{0}(x)\delta^{\prime}(x-y)\,$,
$\{{\cal H}_{1}(x),{\cal H}_{1}(y)\}$ $\!\!=\!\!$ $\partial_{1}{\cal H}_{1}\delta(x-y)+2{\cal H}_{1}(x)\delta^{\prime}(x-y)\,$,
$\{{\cal H}_{0}(x),{\cal H}_{0}(y)\}$ $\!\!=\!\!$ $\delta^{\prime}(x-y)[{\cal H}_{1}(x)+{\cal H}_{1}(y)]\,.$ (A.2)
The first two relations just express that the density weights of ${\cal H}_{0},{\cal H}_{1}$ are +2+2. By specialization of the general (model independent) gravitational constraint algebra one would expect the last relation to contain a factor explicitly dependent on the spatial metric. Specifically, with the present density conventions one would for the 1+11\!+\!1 dimensional constraint algebra expect an additional $\gamma(\gamma_{11})^{-1}$, $\gamma:=-\det\gamma_{\mu\nu}$, term on the right hand side of the last relation in (A.1). Using (2.2) this evaluates to $\gamma(\gamma_{11})^{-1}=n^{2}e^{\tilde{\sigma}}$, which is a spatial scalar. A short computation shows that the redefinition ${\cal H}_{0}^{{\textsc{ADM}}}:=ne^{\tilde{\sigma}/2}{\cal H}_{0}$, ${\cal H}_{1}^{{\textsc{ADM}}}:={\cal H}_{1}$, maps (A.1) into the expected ADM type constraint algebra.
On general grounds ${\cal H}_{0}(\epsilon)+{\cal H}_{1}(\epsilon^{1})$ is the generator of Hamiltonian gauge variations of any (not explicitly time dependent) functional $F$ on phase space built from the canonical variables, here $\phi,\rho,\sigma$ and their canonical momenta. The descriptors $(\epsilon,\epsilon^{1})$ play the role of $(n,s)$ and can both be viewed as spatial 1-1 densities. The variations $\delta_{\epsilon}^{H}F=\{F,{\cal H}_{0}(\epsilon)+{\cal H}_{1}(\epsilon^{1})\}$ have to be augmented by gauge variations of lapse and shift in order to obtain an invariance of the Hamiltonian action (A.1). The result is:
$\delta^{{\textsc{h}}}_{\epsilon}S^{{\textsc{h}}}=0\,,\quad\mbox{if}\quad\epsilon|_{t_{i}}=0=\epsilon|_{t_{f}}\,$, (A.3)
where the restriction on the temporal gauge descriptors has a well understood origin and carries over from full Einstein gravity [16]. For convenient reference we display the complete set of Hamiltonian gauge variations:
$\delta^{{\textsc{h}}}_{\epsilon}n$ $=$ $\partial_{0}\epsilon-(s\partial_{1}\epsilon-\epsilon\partial_{1}s)+\epsilon^{1}\partial_{1}n-n\partial_{1}\epsilon^{1}\,$,
$\delta^{{\textsc{h}}}_{\epsilon}s$ $=$ $\partial_{0}\epsilon^{1}+(\epsilon^{1}\partial_{1}s-s\partial_{1}\epsilon^{1})+\epsilon\partial_{1}n-n\partial_{1}\epsilon\,$, (A.4a)
$\delta^{{\textsc{h}}}_{\epsilon}\sigma$ $=$ $-\lambda_{{\textsc{n}}}\epsilon\pi^{\rho}+\epsilon^{1}\partial_{1}\sigma+2\partial_{1}\epsilon^{1}\,$,
$\delta^{{\textsc{h}}}_{\epsilon}\rho$ $=$ $-\lambda_{{\textsc{n}}}\epsilon\pi^{\sigma}+\epsilon^{1}\partial_{1}\rho\,$,
$\delta^{{\textsc{h}}}_{\epsilon}\phi$ $=$ $\lambda_{{\textsc{n}}}\frac{\epsilon}{\rho}\pi^{\phi}+\epsilon^{1}\partial_{1}\phi\,$, (A.4b)
$\delta^{{\textsc{h}}}_{\epsilon}\pi^{\sigma}$ $=$ $\partial_{1}\big(\!-\frac{\epsilon}{\lambda_{{\textsc{n}}}}\partial_{1}\rho+\epsilon^{1}\pi^{\sigma}\big)\,$,
$\delta^{{\textsc{h}}}_{\epsilon}\pi^{\rho}$ $=$ $\partial_{1}\big(\!-\frac{\epsilon}{\lambda_{{\textsc{n}}}}\partial_{1}\sigma+\epsilon^{1}\pi^{\rho}\big)-\frac{2}{\lambda_{{\textsc{n}}}}\partial_{1}^{2}\epsilon-\frac{\epsilon}{2\lambda_{{\textsc{n}}}}(\partial_{1}\phi)^{2}+\frac{\epsilon\lambda_{{\textsc{n}}}}{2\rho^{2}}(\pi^{\phi})^{2}\,$,
$\delta^{{\textsc{h}}}_{\epsilon}\pi^{\phi}$ $=$ $\partial_{1}\big(\frac{\epsilon\rho}{\lambda_{{\textsc{n}}}}\partial_{1}\phi+\epsilon^{1}\pi^{\phi}\big)\,.$ (A.4c)
Unless stated otherwise we take the descriptors $\epsilon,\epsilon^{1}$ to be spatially periodic. The field equations of the system are readily obtained by varying $S^{{\textsc{h}}}$. In addition to the constraints $\delta S^{{\textsc{h}}}/\delta n=-{\cal H}_{0}$, $\delta S^{{\textsc{h}}}/\delta s=-{\cal H}_{1}$, there are the velocity-momentum relations $e_{0}(\rho)+\lambda_{{\textsc{n}}}n\pi^{\sigma}=0$, $e_{0}(\sigma)+\lambda_{{\textsc{n}}}n\pi^{\rho}=0$, $\rho e_{0}(\phi)-\lambda_{{\textsc{n}}}n\pi^{\phi}=0$, and the evolution equations
$0=\frac{\delta S^{{\textsc{h}}}}{\delta\rho}$ $\!\!=\!\!$ $-e_{0}(\pi^{\rho})-\frac{1}{\lambda_{{\textsc{n}}}}\partial_{1}(n\partial_{1}\sigma)-\frac{2}{\lambda_{{\textsc{n}}}}\partial_{1}^{2}n+\frac{\lambda_{{\textsc{n}}}n}{2\rho^{2}}(\pi^{\phi})^{2}-\frac{n}{2\lambda_{{\textsc{n}}}}(\partial_{1}\phi)^{2}\,$,
$0=\frac{\delta S^{{\textsc{h}}}}{\delta\sigma}$ $\!\!=\!\!$ $-e_{0}(\pi^{\sigma})-\frac{1}{\lambda_{{\textsc{n}}}}\partial_{1}(n\partial_{1}\rho)\,$,
$0=\frac{\delta S^{{\textsc{h}}}}{\delta\phi}$ $\!\!=\!\!$ $-e_{0}(\pi^{\phi})+\frac{1}{\lambda_{{\textsc{n}}}}\partial_{1}(\rho n\partial_{1}\phi)\,.$ (A.5)
These are such that the constraints are preserved under time evolution
$e_{0}({\cal H}_{0})-\frac{1}{n}\partial_{1}(n^{2}{\cal H}_{1})=0\,,\quad e_{0}({\cal H}_{1})-\frac{1}{n}\partial_{1}(n^{2}{\cal H}_{0})=0\,.$ (A.6)
Generally, let $F$ be a (not explicitly time dependent) functional $F$ on phase space that is also a spatial tensor. Then $\{F,{\cal H}_{0}(n)\}=e_{0}(F)$ encodes the evolution equations.
A.2 VD Gowdy cosmologies: the Hamiltonian action and its symmetries
The Hamiltonian action of the velocity dominated system is
$S^{{\textsc{hvd}}}$ $\!\!=\!\!$ $\int_{t_{i}}^{t_{f}}\!\!dx^{0}\int_{0}^{2\pi}\!\!dx^{1}\big\{\pi^{\varrho}e_{0}(\varrho)+\pi^{\varsigma}e_{0}(\varsigma)+\pi^{\varphi}e_{0}(\varphi)-n{\cal H}_{0}^{{\textsc{vd}}}\big\}\,$,
${\cal H}_{0}^{{\textsc{vd}}}$ $\!\!=\!\!$ $-\lambda_{{\textsc{n}}}\pi^{\varsigma}\pi^{\varrho}+\frac{\lambda_{{\textsc{n}}}}{2\varrho}(\pi^{\varphi})^{2}\,$,
${\cal H}_{1}^{{\textsc{vd}}}$ $\!\!=\!\!$ $\pi^{\varrho}\partial_{1}\varrho+\pi^{\varsigma}\partial_{1}\varsigma-2\partial_{1}\pi^{\varsigma}+\pi^{\varphi}\partial_{1}\varphi\,.$ (A.7)
Here we write $\varrho,\varsigma,\varphi$ for the counterparts of $\rho,\sigma,\phi$ in the VD system. Strictly speaking, different symbols should be used also for the VD versions of $n$, $s$, but for readability’s sake we keep the original ones. The phase space is again equipped with the natural Poisson structure. The constraint algebra simplifies to
$\{{\cal H}_{0}^{{\textsc{vd}}}(x),{\cal H}_{1}^{{\textsc{vd}}}(y)\}$ $\!\!=\!\!$ $\partial_{1}{\cal H}_{0}^{{\textsc{vd}}}\delta(x-y)+2{\cal H}_{0}^{{\textsc{vd}}}(x)\delta^{\prime}(x-y)\,$,
$\{{\cal H}_{1}^{{\textsc{vd}}}(x),{\cal H}_{1}^{{\textsc{vd}}}(y)\}$ $\!\!=\!\!$ $\partial_{1}{\cal H}_{1}^{{\textsc{vd}}}\delta(x-y)+2{\cal H}_{1}^{{\textsc{vd}}}(x)\delta^{\prime}(x-y)\,$,
$\{{\cal H}_{0}^{{\textsc{vd}}}(x),{\cal H}_{0}^{{\textsc{vd}}}(y)\}$ $\!\!=\!\!$ $0\,.$ (A.8)
The Poisson commutativity of the Hamiltonian constraint is characteristic for Carroll type gravity systems, see e.g. [22, 23], and accounts for the dynamical decoupling of spatial points. Nevertheless, the action (A.2) is still invariant under diffeomorphism type gauge symmetries parameterized by two independent functions of two variables. Explicitly,
$\delta^{{\textsc{hvd}}}_{\epsilon}S^{{\textsc{hvd}}}=0\,,\quad\mbox{if}\quad\epsilon|_{t_{i}}=0=\epsilon|_{t_{f}}\,$, (A.9)
holds with the modified gauge variations
$\delta^{{\textsc{hvd}}}_{\epsilon}n$ $=$ $\partial_{0}\epsilon-(s\partial_{1}\epsilon-\epsilon\partial_{1}s)+\epsilon^{1}\partial_{1}n-n\partial_{1}\epsilon^{1}\,$,
$\delta^{{\textsc{hvd}}}_{\epsilon}s$ $=$ $\partial_{0}\epsilon^{1}+(\epsilon^{1}\partial_{1}s-s\partial_{1}\epsilon^{1})$ (A.10a)
$\delta^{{\textsc{hvd}}}_{\epsilon}\varsigma$ $=$ $-\lambda_{{\textsc{n}}}\epsilon\pi^{\varrho}+\epsilon^{1}\partial_{1}\varsigma+2\partial_{1}\epsilon^{1}\,$,
$\delta^{{\textsc{hvd}}}_{\epsilon}\varrho$ $=$ $-\lambda_{{\textsc{n}}}\epsilon\pi^{\varsigma}+\epsilon^{1}\partial_{1}\varrho\,$,
$\delta^{{\textsc{hvd}}}_{\epsilon}\varphi$ $=$ $\lambda_{{\textsc{n}}}\frac{\epsilon}{\varrho}\pi^{\varphi}+\epsilon^{1}\partial_{1}\varphi\,$, (A.10b)
$\delta^{{\textsc{hvd}}}_{\epsilon}\pi^{\varsigma}$ $=$ $\partial_{1}(\epsilon^{1}\pi^{\varsigma})\,$,
$\delta^{{\textsc{hvd}}}_{\epsilon}\pi^{\varrho}$ $=$ $\partial_{1}(\epsilon^{1}\pi^{\varrho})+\frac{\epsilon\lambda_{{\textsc{n}}}}{2\varrho^{2}}(\pi^{\varphi})^{2}\,$,
$\delta^{{\textsc{hvd}}}_{\epsilon}\pi^{\varphi}$ $=$ $\partial_{1}(\epsilon^{1}\pi^{\varphi})\,.$ (A.10c)
So far we treated the VD system as a Carroll-type gravitational theory in its own right. The asymptotic velocity domination property posits its dynamical relevance for the full Gowdy system, initially classically and on-shell. There is, however, also a kinematic relation in that the VD system arises as a (classical) scaling limit from the original one. For Carroll systems this is usually done via a (coordinate dependent) “speed of light to zero” limit. In the present context, we adopt a merely foliation dependent scaling defined on phase space as follows:
[…]
In the line element (2.1) this enhances for $\ell\gg 1$ spacelike distances compared to timelike ones. Neighboring worldlines are harder to communicate with and the light cone structure appears anti-Newtonian. The scaling of the reduced Newton constant $\lambda_{{\textsc{n}}}$ is adjusted such that the rescaled action (A.1) has a well defined limit as $\ell\rightarrow\infty$. Indeed, the limit is precisely $S^{{\textsc{hvd}}}$ from (A.2), upon renaming of the fields. Likewise, the constraint algebra (A.2) is a contraction of (A.1), and the gauge variations (A.10) are limiting versions of (A.4).
Appendix B Classical Dirac observables
Two Killing vector reductions of Einstein gravity often admit an infinite set of nonlocal conserved charges whose existence is related to that of a Lax pair. On general grounds such conserved charges are expected to give rise to (normally elusive) Dirac observables, Poisson commuting with the constraints. For the polarized Gowdy system the conserved charges become linear functionals of the basic Gowdy scalar. A systematic characterization of Dirac observables for the polarized Gowdy system has been given by Torre [15], using an extended phase space formulation. Due to the use of a polar decomposition in the last step it falls however slightly short of providing fully explicit expressions for them. Here we provide such a construction on the original rather than an extended phase space in a way that makes contact to the velocity dominated limit and in principle generalizes to the non-polarized case.
A Dirac observable $\cal O$} is by definition a functional of the phase space variables that weakly Poisson commutes with the constraints, see e.g. [16]. In the situation at hand, we aim at the strong vanishing
$\{{\cal H}_{0},{\cal O}\}=0=\{{\cal H}_{1},{\cal O}\}\,.$ (B.1)
In addition to the local gauge invariance (A.4) the action $S^{{\textsc{h}}}$ also has a trivial global invariance under constant shifts $\phi(x)\mapsto\phi(x)+\alpha$, $\alpha\in\mathbb{R}$. It is generated by the Noether charge $Q=(2\pi)^{-1}\int_{0}^{2\pi}\!dx^{1}\pi^{\phi}(x)$. Since the constraints ${\cal H}_{0},{\cal H}_{1}$ Poisson commute with $Q$ the iterated images $\{Q,\ldots\{Q,{\cal O}\}$...\} are Dirac observables as well. To avoid such rather trivial chains it is natural to require that the Dirac observables Poisson commute with $Q$ as well
$\{Q,{\cal O}\}=0\,.$ (B.2)
The goal in the following is to construct an infinite set of Dirac observables which satisfy (B.1), (B.2) strongly, off-shell, and without gauge fixing. See e.g. [16] for the significance of the latter. Since ${\cal H}_{1}$ generates infinitesimal spatial translations spatial integrals of spatial +1+1 densities will satisfy the second condition in (B.1). A solution of the first condition, on the other hand, necessitates to the construction of (nonlocal) conserved charges, which is normally elusive as it requires a constructive approach to the solution of the initial value problem. As mentioned, this is often feasible for the two Killing vector reductions of Einstein gravity, but for Gowdy cosmologies has not been accomplished beyond the polarized sector. In the following we describe a construction of off-shell Dirac observables for the polarized Gowdy system that generalizes to the unpolarized sector.
Gauge variation of the generating current. Consider the basic currents in the Hamiltonian formalism
$\jmath_{0}^{\textsc{h}}:=\pi^{\phi}\,,\quad\jmath_{1}^{\textsc{h}}:=\frac{\rho n}{\lambda_{{\textsc{n}}}}\partial_{1}\phi\,$, (B.3)
whose conservation is one of the Hamiltonian evolution equations $e_{0}(\pi^{\phi})=\partial_{1}(\frac{\rho n}{\lambda_{{\textsc{n}}}}\partial_{1}\phi)$. Using the Hamiltonian gauge variations (A.4) one can check that
$\delta_{\epsilon}^{\textsc{h}}\jmath_{0}^{\textsc{h}}=\partial_{1}\Big(\frac{\epsilon}{n}\jmath_{1}^{\textsc{h}}+\epsilon^{1}\jmath_{0}^{\textsc{h}}\Big)\,$, (B.4)
holds without using any equations of motion.
Next, we aim at a one-parameter generalization of the form [^6] The origin of these expressions lies in the $\partial/\partial\theta$ derivatives of a Lax pair for the 2-Killing vector reductions, see [33]. Use of the Lax pair is overkill for the polarized case but (B) turns out to generalize to the non-polarized Gowdy system.
$\jmath_{0}^{\textsc{h}}(\theta)$ $:=$ $\frac{\lambda_{{\textsc{n}}}^{2}}{[\lambda_{{\textsc{n}}}^{2}(\theta+\tilde{\rho})^{2}-\rho^{2}]^{3/2}}[\lambda_{{\textsc{n}}}(\theta+\tilde{\rho})\pi^{\phi}-\frac{\rho^{2}}{\lambda_{{\textsc{n}}}}\partial_{1}\phi]\,$,
$\jmath_{1}^{\textsc{h}}(\theta)$ $:=$ $\frac{\lambda_{{\textsc{n}}}^{2}}{[\lambda_{{\textsc{n}}}^{2}(\theta+\tilde{\rho})^{2}-\rho^{2}]^{3/2}}[(\theta+\tilde{\rho})\rho n\partial_{1}\phi-\rho n\pi^{\phi}]\,$, (B.5)
for complex $\theta\in\mathbb{C}$, ${\rm Im}\theta\neq 0$. Here we define
$\tilde{\rho}(x^{0},x^{1}):=-\int_{y^{1}}^{x^{1}}\!dx\,\pi^{\sigma}(x^{0},x)+\tilde{\rho}(x^{0},y^{1})\,$, (B.6)
for some reference point $y^{1}$. For its Hamiltonian gauge variation one can self-consistently take
$\delta_{\epsilon}^{\textsc{h}}\tilde{\rho}=-\epsilon^{1}\pi^{\sigma}+\frac{\epsilon}{\lambda_{{\textsc{n}}}}\partial_{1}\rho\,.$ (B.7)
This comes about as follows: since $\tilde{\rho}$ is not spatially periodic (see below) the natural class of gauge-type spatial diffeomorphisms for its are those of the real line vanishing outside a compact region. Evaluating the gauge variation of the integral in (B.7) one finds that that contribution from the $y^{1}$ endpoint vanishes if $y^{1}\rightarrow$-$\infty$. One can therefore consistently take $\tilde{\rho}(x^{0},-\infty)=0$ and obtains (B.7). Once (B.7) is in place we treat $\tilde{\rho}$ as spatially quasi-periodic via (B.9) and for the fundamental domain $0,2\pi]$ use gauge descriptors of the same type as in (A.4).
Using only (A.4) and (B.7) one finds
$\delta_{\epsilon}^{\textsc{h}}\jmath_{0}^{\textsc{h}}(\theta)=\partial_{1}\Big(\frac{\epsilon}{n}\jmath_{1}^{\textsc{h}}(\theta)+\epsilon^{1}\jmath_{0}^{\textsc{h}}(\theta)\Big)\,$, (B.8)
holds, again without using any equations of motion. For noncompact spatial sections and with suitable fall-off conditions the relation (B.8) would directly identify the spatial integral of $\jmath_{0}^{\textsc{h}}(\theta)$ as a one-parameter family of Dirac observables for the polarized Gowdy system.
Periodic extension. For the $T^{3}$ Gowdy cosmologies of prime interest in the context of AVD, the identity (B.8) does not directly provide Dirac observables by integration: while all the basic fields and hence $\jmath_{0}^{\textsc{h}},\jmath_{1}^{\textsc{h}}$ are spatially periodic, the off-shell currents $\jmath_{0}^{\textsc{h}}(\theta),\jmath_{1}^{\textsc{h}}(\theta)$ are not. This is because $\tilde{\rho}$ is not spatially periodic (neither on- nor off-shell). From (B.6) one has
$\tilde{\rho}(x^{0},x^{1}\!+\!2\pi)-\tilde{\rho}(x^{0},x^{1})=-\!\int_{0}^{2\pi}\!dx\,\pi^{\sigma}(x^{0},x)=:-\pi^{\sigma}_{0}\,$, (B.9)
using the spatial periodicity of $\pi^{\sigma}$. Importantly, $\pi_{0}^{\sigma}$ itself is gauge invariant,
$\delta^{H}_{\epsilon}\pi^{\sigma}_{0}=0\,$, (B.10)
as can be seen from (A.4). Here and from now on we take the descriptors $\epsilon,\epsilon^{1}$ to be spatially periodic. An alternative way of looking at (B.10) is by noting that $\{\sigma_{0},\pi_{0}^{\sigma}\}=1$, where $\sigma_{0}=(2\pi)^{-1}\int_{0}^{2\pi}\!dx\,\sigma(x)$ is the zero mode of $\sigma$. This zero mode simply does not occur in ${\cal H}_{0},{\cal H}_{1}$, which accounts for (B.10). Taking into account the nonperiodicity of (B) one merely has
$\big\{\int_{0}^{2\pi}\!\!dx\jmath_{0}^{\textsc{h}}(x;\theta),{\cal H}_{0}(\epsilon)+{\cal H}_{1}(\epsilon^{1})\big\}$
$=\Big(\frac{\epsilon}{n}\Big)(0)\big[\jmath_{1}^{\textsc{h}}(2\pi;\theta)-\jmath_{1}^{\textsc{h}}(0;\theta)\big]+\epsilon^{1}(0)\big[\jmath_{0}^{\textsc{h}}(2\pi;\theta)-\jmath_{0}^{\textsc{h}}(0;\theta)\big]\,.$ (B.11)
Here we suppressed the shared time argument and assumed as before that the descriptors $\epsilon,\epsilon^{1}$ are spatially periodic. The basic idea now is to replace $\jmath_{0}^{\textsc{h}}(x,\theta)$ on the left hand side with its $2\pi$-periodic extension, i.e. with $\sum_{n}\jmath_{0}^{\textsc{h}}(x\!+\!2\pi n,\theta)$. Assuming that this sum converges, it will give rise on the right hand side to a telescopic sum which vanishes. In principle, the plain periodic extension can be modified by multiplying the $n$-th term by a function $s_{n}$ that is itself gauge invariant and obeys $s_{n}|_{x^{1}+2\pi}=s_{n+1}$ as a function of $x^{1}$. Since the only source of non-periodicity in our context is the combination $\theta+\tilde{\rho}$ it is natural to take $s_{n}$ a function of $\theta+\tilde{\rho}$, in which case $\delta^{\textsc{h}}_{\epsilon}s_{n}=0$ is a stringent requirement.
Proceeding along these lines, we consider for ${\rm Im}\theta\neq 0$ the $s_{n}$-modified periodic extension of (B). It leads to sums of the form
$c_{0}(\tilde{\rho}+\theta,\rho)$ $:=$ $\sum_{n\in\mathbb{Z}}\frac{s_{n}(\theta\!+\!\tilde{\rho})\lambda_{{\textsc{n}}}^{3}(\theta+\tilde{\rho}-n\pi_{0}^{\sigma})}{[\lambda_{{\textsc{n}}}^{2}(\theta+\tilde{\rho}-n\pi_{0}^{\sigma})^{2}-\rho^{2}]^{3/2}}\,$,
$c_{1}(\tilde{\rho}+\theta,\rho)$ $:=$ $\sum_{n\in\mathbb{Z}}\frac{\lambda_{{\textsc{n}}}^{3}s_{n}(\theta\!+\!\tilde{\rho})}{[\lambda_{{\textsc{n}}}^{2}(\theta+\tilde{\rho}-n\pi_{0}^{\sigma})^{2}-\rho^{2}]^{3/2}}\,.$ (B.12)
For reasons that will become clear shortly we take for $s_{n}$ the generalized sign function $s_{n}(z-in\pi_{0}^{\sigma})=\exp\{3i{\rm Arg}(z-in\pi_{0}^{\sigma})\}(\exp\{2i{\rm Arg}(z-in\pi_{0}^{\sigma})\})^{-3/2}$, ${\rm Im}z\neq 0$. The sums (B) then converge absolutely and for ${\rm Im}\theta\neq 0$ define spatially periodic functions of $x^{0},x^{1}$. The summed expansions (LABEL:csums2) below also show them to be smooth in $\rho,\tilde{\rho}$ for ${\rm Im}(\theta)\neq 0$. Inserting (B) into (B) gives
$\jmath_{0}^{\textsc{p}}(\theta):=\sum_{n\in\mathbb{Z}}s_{n}(\theta+\tilde{\rho})\jmath^{\textsc{h}}_{0}(\theta-n\pi_{0}^{\sigma})=c_{0}(\tilde{\rho}+\theta,\rho)\pi^{\phi}-c_{1}(\tilde{\rho}+\theta,\rho)\frac{\rho^{2}}{\lambda_{{\textsc{n}}}^{2}}\partial_{1}\phi\,$,
$\jmath_{1}^{{\textsc{p}}}(\theta):=\sum_{n\in\mathbb{Z}}s_{n}(\theta+\tilde{\rho})\jmath^{\textsc{h}}_{1}(\theta-n\pi_{0}^{\sigma})=c_{0}(\tilde{\rho}+\theta,\rho)\frac{\rho}{\lambda_{{\textsc{n}}}}n\partial_{1}\phi-c_{1}(\tilde{\rho}+\theta,\rho)\frac{\rho}{\lambda_{{\textsc{n}}}}n\pi^{\phi}\,.$ (B.13)
The condition for the summed versions to obey equation (B.8) amounts to a simple pair of differential equations for $c_{0},c_{1}$,
$\partial_{\tilde{\rho}}(\rho^{2}c_{1})=-\lambda_{{\textsc{n}}}^{2}\rho\partial_{\rho}c_{0}\,,\quad\rho\partial_{\tilde{\rho}}c_{0}=-\partial_{\rho}(\rho^{2}c_{1})\,.$ (B.14)
These imply that $c_{0}$ and $\rho^{2}c_{1}$ are spatially periodic solutions of $\lambda_{{\textsc{n}}}^{-2}\partial_{\tilde{\rho}}^{2}f=\partial_{\rho}^{2}f\pm\rho^{-1}\partial_{\rho}f$, respectively.
The spatial periodicity can be rendered explicit by expanding in powers of $\rho$ using
$\!\!\!\!\!\!\!\!\!\!\frac{\lambda_{{\textsc{n}}}^{3}(\theta\!+\!\tilde{\rho}\!-\!n\pi_{0}^{\sigma})}{[\lambda_{{\textsc{n}}}^{2}(\theta\!+\!\tilde{\rho}\!-\!n\pi_{0}^{\sigma})^{2}-\rho^{2}]^{3/2}}=\sum_{l\geq 0}{-3/2\choose l}(-\rho^{2})^{l}s_{n}(\theta\!+\!\tilde{\rho})\lambda_{{\textsc{n}}}^{2}[\lambda_{{\textsc{n}}}(\theta\!+\!\tilde{\rho}\!-\!n\pi_{0}^{\sigma})]^{-2l-2},$
$\!\!\!\!\!\!\!\!\!\!\frac{\lambda_{{\textsc{n}}}^{3}}{[\lambda_{{\textsc{n}}}^{2}(\theta\!+\!\tilde{\rho}\!-\!n\pi_{0}^{\sigma})^{2}-\rho^{2}]^{3/2}}=\sum_{l\geq 0}{-3/2\choose l}(-\rho^{2})^{l}s_{n}(\theta\!+\!\tilde{\rho})\lambda_{{\textsc{n}}}^{3}[\lambda_{{\textsc{n}}}(\theta\!+\!\tilde{\rho}\!-\!n\pi_{0}^{\sigma})]^{-2l-3},$ (B.15)
The point of the $s_{n}$ modulated sums in (B) is simply to cancel the corresponding $s_{n}$’s in the expansions (LABEL:csums2). The resulting sums over $n$ can then readily be performed. For $2\leq l\in\mathbb{N}$ one has
$\sum_{n\in\mathbb{Z}}\frac{1}{(a\!+\!bn)^{l}}=\frac{(-)^{l}}{(l\!-\!1)!}\Big(\frac{\partial}{\partial a}\Big)^{l-2}\sum_{n\in\mathbb{Z}}\frac{1}{(a\!+\!bn)^{2}}=\frac{\pi^{2}}{b^{2}}\frac{(-)^{l}}{(l\!-\!1)!}\Big(\frac{\partial}{\partial a}\Big)^{l-2}\frac{1}{\sin^{2}(\pi a/b)}\,.$ (B.16)
[…]
$\!\!\!\!\!\!\!\!\!\!c_{0}(\theta\!+\!\tilde{\rho},\rho)$ $\!\!=\!\!$ $\Big(\frac{\pi}{\pi_{0}^{\sigma}}\Big)^{2}\sum_{l\geq 0}{-3/2\choose l}\frac{(-\rho^{2}/\lambda_{{\textsc{n}}}^{2})^{l}}{(2l+1)!}\Big(\frac{\partial}{\partial\theta}\Big)^{2l}\!\frac{1}{\sin^{2}\big(\frac{\pi}{\pi_{0}^{\sigma}}(\theta\!+\!\tilde{\rho})\big)}$,
$\!\!\!\!\!\!\!\!\!\!c_{1}(\theta\!+\!\tilde{\rho},\rho)$ $\!\!=\!\!$ $-\Big(\frac{\pi}{\pi_{0}^{\sigma}}\Big)^{2}\sum_{l\geq 0}{-3/2\choose l}\frac{(-\rho^{2}/\lambda_{{\textsc{n}}}^{2})^{l}}{(2l+2)!}\Big(\frac{\partial}{\partial\theta}\Big)^{2l+1}\!\frac{1}{\sin^{2}\big(\frac{\pi}{\pi_{0}^{\sigma}}(\theta\!+\!\tilde{\rho})\big)}$. (B.17)
As a check on the innocuous nature of the $s_{n}$ modulation in (B) one can verify that the expressions (B) indeed solve the differential equations (B.14). This yields an expansion of $\jmath_{0}^{{\textsc{p}}}(\theta)$ of the form $\jmath_{0}^{{\textsc{p}}}(\theta)=\sum_{j\geq 0}\jmath_{0}^{{\textsc{p}}}(\theta)_{2j}$, where the $j$-th term has scaling weight $\ell^{-2j}$ under (A.11). Explicitly,
$\jmath_{0}^{{\textsc{p}}}(\theta)_{0}$ $:=$ $\Big(\frac{\pi}{\pi_{0}^{\sigma}}\Big)^{2}\frac{\pi^{\phi}}{\sin^{2}\big(\frac{\pi}{\pi_{0}^{\sigma}}(\theta+\tilde{\rho})\big)}\,$, (B.18)
$\jmath_{0}^{{\textsc{p}}}(\theta)_{2j}$ $:=$ $\Big(\frac{\pi}{\pi_{0}^{\sigma}}\Big)^{2}{-3/2\choose j}\frac{(-1)^{j}}{(2j\!+\!1)!}\Big(\frac{\rho^{2}}{\lambda_{{\textsc{n}}}^{2}}\Big)^{j}$
$\times\Big(\frac{\partial}{\partial\theta}\Big)^{2j}\Big[\frac{\pi^{\phi}}{\sin^{2}\big(\frac{\pi}{\pi_{0}^{\sigma}}(\theta+\tilde{\rho})\big)}+\frac{1}{(2j\!+\!2)}\frac{\partial}{\partial\theta}\frac{\partial_{1}\phi}{\sin^{2}\big(\frac{\pi}{\pi_{0}^{\sigma}}(\theta+\tilde{\rho})\big)}\Big$]\,.
Finally, we note a third rewriting of the $c_{0},c_{1}$. Using
$\frac{1}{\sin^{2}(z)}=-4\sum_{n\geq 1}ne^{\pm 2inz}\,,\quad\pm{\rm Im}z>0\,$,
$\sum_{l\geq 0}{-3/2\choose l}\frac{x^{2l}}{(2l+1)!}=J_{0}(x)\,,\quad\sum_{l\geq 0}{-3/2\choose l}\frac{x^{2l}}{(2l+2)!}=\frac{J_{1}(x)}{x}\,$, (B.19)
to manipulate (B) one finds [^7] Inserting for the left hand side the original sums (B) produces Bessel function identities which appear to be new.
$c_{0}(\theta+\tilde{\rho},\rho)$ $\!\!=\!\!$ $-\Big(\frac{2\pi}{\pi_{0}^{\sigma}}\Big)^{2}\sum_{n\geq 1}ne^{\pm 2i\pi n(\theta+\tilde{\rho})/\pi_{0}^{\sigma}}J_{0}\Big(\frac{2n\rho\pi}{\lambda_{{\textsc{n}}}\pi_{0}^{\sigma}}\Big)\,$,
$c_{1}(\theta+\tilde{\rho},\rho)$ $\!\!=\!\!$ $\pm i\Big(\frac{2\pi}{\pi_{0}^{\sigma}}\Big)^{2}\sum_{n\geq 1}ne^{\pm 2i\pi n(\theta+\tilde{\rho})/\pi_{0}^{\sigma}}\;\frac{\lambda_{{\textsc{n}}}}{\rho}J_{1}\Big(\frac{2n\rho\pi}{\lambda_{{\textsc{n}}}\pi_{0}^{\sigma}}\Big)\,$, (B.20)
for $\pm{\rm Im}\theta>0$. The occurrence of Bessel functions is explained by the comment after (B.14), where regularity at $\rho=0$ excludes the Bessel $Y$-type solutions. Conversely, it follows from (B) that the Dirac observables (B.21) are a complete set in the sector regular for $\rho/\lambda_{{\textsc{n}}}\rightarrow 0$. Let us stress that (B) holds off-shell, conceptually the Bessel functions entering here are unrelated to those in (2.3).
Result B.1.
Set
${\cal O}(\theta):=\int_{0}^{2\pi}\!\frac{dx^{1}}{2\pi}\big\{c_{0}(\tilde{\rho}+\theta,\rho)\pi^{\phi}-c_{1}(\tilde{\rho}+\theta,\rho)\frac{\rho^{2}}{\lambda_{{\textsc{n}}}^{2}}\partial_{1}\phi\big\}\,,\quad{\rm Im}\theta\neq 0\,$, (B.21)
with $c_{0}$, $c_{1}$ given alternatively by (B), (B), (B). Then ${\cal O}(\theta)$ is a one-parameter family of strong, off-shell Dirac observables for the polarized $T^{3}$ Gowdy cosmologies. The integrand $\jmath_{0}^{{\textsc{p}}}(\theta)$ has an ‘anti-Newtonian’ expansion in even powers of $\rho/\lambda_{{\textsc{n}}}$ given in (B.18). The leading term gives rise to one-parameter family of strong, off-shell Dirac observables in the VD system via
${\cal O}^{{\textsc{vd}}}(\theta)=\int_{0}^{2\pi}\!\!\frac{dx^{1}}{2\pi}\,\jmath_{0}^{{\textsc{p}}}(\theta)_{0}\big|_{\pi^{\phi}\mapsto\pi^{\varphi},\pi^{\sigma}\mapsto\pi^{\varsigma}}\,.$ (B.22)
Recall that $\rho>0$ is a temporal function whose $\rho=0$ level set can be identified with the Big Bang. The Dirac observables (B.21) therefore remain regular at the Big Bang and (B.18) provides a systematic expansion away from it, notably in inverse powers of the reduced Newton constant $\lambda_{{\textsc{n}}}$. Upon substitution of the dynamical variables of the VD system the $\rho=0$ term defines Dirac observables in the VD system, as indicated in (B.22). The invariance $\delta_{\epsilon}^{{\textsc{hvd}}}{\cal O}^{{\textsc{vd}}}(\theta)=0$ can be verified by direct computation from (A.10).
On-shell specialization. Subject to the evolution equations the left hand side of the defining relation $\delta_{\epsilon}^{\textsc{h}}\jmath_{0}^{\textsc{p}}(\theta)=\partial_{1}[(\epsilon/n)\jmath_{1}^{\textsc{p}}(\theta)+\epsilon^{1}\jmath_{0}^{\textsc{p}}(\theta)]$ must for $\epsilon=n,\epsilon^{1}=0$ produce the infinitesimal time evolution $e_{0}(\jmath_{0}^{{\textsc{p}}}(\theta))$. The Dirac observables thereby give rise to a one-parameter family of on-shell conserved currents
$e_{0}(\jmath^{{\textsc{p}}}_{0}(\theta))=\partial_{1}\jmath_{1}^{{\textsc{p}}}(\theta)\,.$ (B.23)
Using the evolution equations (A.1) and (B.14) this can indeed be verified directly. The on-shell version of $\tilde{\rho}$ obeys $e_{0}(\rho)=\lambda_{{\textsc{n}}}n\partial_{1}\tilde{\rho}$, $\lambda_{{\textsc{n}}}e_{0}(\tilde{\rho})=n\partial_{1}\rho$, and the periodicity constant is $\pi_{0}^{\sigma}=\lambda_{{\textsc{n}}}^{-1}\int_{0}^{2\pi}dxe_{0}(\rho)/n$. This holds without gauge fixing and/or the subsequent coordinate specifications (2.7), (2.13). In particular, the scaling transformations (A.11) remain intact and can be used to introduce an $\ell$-grading as before. After the scaling weights have been identified we proceed to fix the proper time gauge and impose the coordinate choices (2.7), (2.13). The latter spoil the scaling as $t$ and $\zeta$ would have to be assigned weights $\ell^{2}$ and $\ell$, respectively, but the limit $\ell\rightarrow\infty$ is not supposed to act on the coordinates. We therefore also set $\lambda_{{\textsc{n}}}=1$ and identify $\tilde{\rho}$ with $\zeta$, consistent with the specialization to $\pi_{0}^{\sigma}=2\pi$ and the comment after (2.7). We denote the so-specialized coefficient functions by $c_{0}(\theta+\zeta,e^{\tau})$, $c_{1}(\theta+\zeta,e^{\tau})$ (with $\lambda_{{\textsc{n}}}=1,\pi_{0}^{\sigma}=2\pi$ understood). They can be expressed alternatively as the specializations of (B),(B) or (B). For the Dirac observables (B.21) this yields an on-shell specialization of the form
${\cal O}^{\rm on}(\theta)$ $\!\!=\!\!$ $\int_{0}^{2\pi}\!\frac{d\zeta}{2\pi}\big\{c_{0}(\theta\!+\!\zeta,e^{\tau})\partial_{\tau}\phi^{\rm on}-c_{1}(\theta\!+\!\zeta,e^{\tau})e^{2\tau}\partial_{\zeta}\phi^{\rm on}\big\}\,$,
$\phi^{\rm on}$ $\!\!=\!\!$ $\sum_{n\in\mathbb{Z}}\big\{T_{n}(\tau)e^{in\zeta}a_{n}+T_{n}(\tau)^{*}e^{-in\zeta}a_{n}^{*}\big\}\,$, (B.24)
with $T_{n}$ a Wronskian normalized solution of $\partial_{\tau}^{2}+e^{2\tau}n^{2}]T_{n}(\tau)=0$. By construction $\partial_{\tau}{\cal O}^{\rm on}(\theta)=0$, which can also be verified directly from (B.14). The value of ${\cal O}^{\rm on}(\theta)$ can therefore be found from the $\tau\rightarrow$-$\infty$ limit of the integrand. Since $c_{0}$, $c_{1}$ are regular at $\rho=0$ and AVD implies $\partial_{\tau}\phi\rightarrow v$ for $\tau\rightarrow$-$\infty$ (see Section 2.2) one infers ${\cal O}^{\rm on}(\theta)={\cal O}_{{\textsc{vd}}}^{\rm on}(\theta)$. In fact, a fairly explicit expression for the on-shell value can be found as follows: inserting $\phi^{\rm on}$ into the first line of (B) one finds
${\cal O}^{\rm on}(\theta)$ $\!\!=\!\!$ $\sum_{n\neq 0}\big\{{\cal O}_{n}(\theta)a_{n}+{\cal O}_{n}(\theta)^{*}a_{n}^{*}\big\}\,$,
${\cal O}_{n}(\theta)$ $\!\!=\!\!$ $\frac{i}{\sqrt{\pi}}(\lambda_{n}\!-\!\mu_{n})\theta(\mp n)|n|e^{\pm i|n|\theta}\,,\quad\pm{\rm Im}\theta>0\,.$ (B.25)
We outline the derivation. In a first step, the coefficients ${\cal O}_{n}(\theta)$ come out as a linear combination of $\partial_{\tau}T_{n}(\tau)$ and $ine^{2\tau}T_{n}(\tau)$ with coefficients that are Fourier coefficients of the $c_{0},c_{1}$. The latter are conveniently evaluated using (B) and come out as
$\int_{0}^{2\pi}\!\frac{d\zeta}{2\pi}e^{in\zeta}c_{0}(\zeta+\theta,e^{\tau})=-J_{0}(e^{\tau}|n|)e^{\pm i|n|\theta}\theta(-|n|)\,$,
$\int_{0}^{2\pi}\!\frac{d\zeta}{2\pi}e^{in\zeta}c_{1}(\zeta+\theta,e^{\tau})=\pm i|n|J_{1}(e^{\tau}|n|)e^{\pm i|n|\theta}\theta(-|n|)\,$, (B.26)
where $\theta(\mp n)$ is the step function, imposing $\mp n\in\mathbb{N}$. Expressing $\partial_{\tau}T_{n}(\tau)$ and $ine^{2\tau}T_{n}(\tau)$ in terms of Bessel $J_{1},Y_{1}$ and $J_{0},Y_{0}$ functions via (2.3) many terms cancel and use of the Wronskian identity $Y_{1}(x)J_{0}(x)-Y_{0}(x)J_{1}(x)=-2/(\pi x)$ yields the above ${\cal O}_{n}(\theta)$. Importantly, for ${\rm Im}\theta\neq 0$ the ${\cal O}_{n}(\theta)$ decay faster than any power in $|n|$. Indeed, for Hadamard states we saw in Section 4.2 that $\lambda_{n}\rightarrow 1$, $\mu_{n}\rightarrow 0$, for $|n|\rightarrow\infty$; the rapid decay would however persist for any $\lambda_{n},\mu_{n}$ polynomially bounded in $|n|$.
An analogous computation can be done for the VD observables (B.22). Here $\varphi^{\rm on}=\sum_{n}\{\mathfrak{t}_{n}(\tau)e^{in\zeta}a_{n}+\mathfrak{t}_{n}(\tau)^{*}e^{-n\zeta}a_{n}^{*}\}$, $\partial_{\tau}^{2}\mathfrak{t}_{n}=0$, needs to be inserted into the much simpler (specialized) on-shell form of ${\cal O}_{0}(\theta)|_{\pi^{\phi}\mapsto\pi^{\varphi},\pi^{\sigma}\mapsto\pi^{\varsigma}}$. The integrals encountered are for $\pm{\rm Im}\theta>0$ readily evaluated by contour deformation, while $\partial_{\tau}\mathfrak{t}_{n}=(-i/\sqrt{\pi})(\tilde{\lambda}_{n}\!-\!\tilde{\mu}_{n})$ is already $\tau$ independent. The result is
${\cal O}_{{\textsc{vd}}}^{\rm on}(\theta)$ $\!\!=\!\!$ $\sum_{n\neq 0}\big\{\widetilde{{\cal O}}_{n}(\theta)a_{n}+\widetilde{{\cal O}}_{n}(\theta)^{*}a_{n}^{*}\big\}\,$,
$\widetilde{{\cal O}}_{n}(\theta)$ $\!\!=\!\!$ $\frac{i}{\sqrt{\pi}}(\tilde{\lambda}_{n}\!-\!\tilde{\mu}_{n})\theta(\mp n)|n|e^{\pm i|n|\theta}\,,\quad\pm{\rm Im}\theta>0\,.$ (B.27)
Since by (3.21) $\tilde{\lambda}_{n}\!-\!\tilde{\mu}_{n}=\lambda_{n}\!-\!\mu_{n}$ one confirms
${\cal O}^{\rm on}(\theta)={\cal O}_{{\textsc{vd}}}^{\rm on}(\theta)\,$, (B.28)
which is another manifestation of AVD.